### v1

# Studying propagating turbulent structures in the near wake of a sphere using Hilbert proper orthogonal decomposition

Abstract
Turbulent flows, despite their apparent randomness, exhibit coherent structures that underpin their dynamics. Proper orthogonal decomposition (POD) has been widely used to extract these structures from experimental data. While periodic features like vortex shedding can be identified using POD mode pairs when periodicity dominates the flow, detecting such structures in complex flows is more challenging. The Hilbert proper orthogonal decomposition (HPOD) addresses this by applying POD to the analytic signal of the turbulent fluctuations, yielding complex modes with a $90^{\circ}$ phase shift between the real and imaginary components. These modes capture propagating structures effectively but introduce filtering artefacts from the Hilbert transform that is used to derive the analytic signal. The current work investigates the relationship between the modes of the POD and those of the HPOD on the velocity fluctuations in the wake of a sphere. By comparing their outputs, POD mode pairs that correspond to the same propagating structures revealed by HPOD are identified. Furthermore, this study explored whether computing the analytic signal of the POD modes can replicate the HPOD modes, offering a more computationally efficient method for determining the pairs of POD modes that represent propagating structures. The results show that the pairs of POD modes identified by the HPOD can be more efficiently determined using the Hilbert transform directly on the POD modes. This method enhances the interpretive power of POD, enabling more detailed analysis of turbulent dynamics without introducing the filtering from the Hilbert transform.
1 Introduction
While turbulence often appears random, many turbulent flows contain underlying structures, such as eddies and vortices, the scale and nature of which vary depending on the conditions of the flow. Identifying these structures allows for a more detailed description of the flow, as well as the isolation of the dominant structures in the turbulence based on their contribution to the quantity or phenomenon of interest, such as turbulent kinetic energy (TKE) or mixing. Ranking these structures based on their significance to a given application or study allows the construction of parsimonious reduced-order models, which can be used to reduce the computational intensity of analyses or to reduce noise from experimental data.
The complex flow phenomena present in the flow over bluff bodies, even at low Reynolds numbers, develop as the Reynolds number increases, and are a subject of ongoing study. In this respect, the flow over spheres serves as a canonical model for flows over axisymmetric bluff bodies, and is of particular interest for developing a broad understanding of these flow phenomena. The phenomena observed in the wake of a sphere include transition from axisymmetry to planar symmetry to asymmetry, the formation and shedding of vortices, the development of shear layer instabilities, and flow separation. As the flow over a sphere varies significantly with Reynolds number, experimental and numerical investigations have covered a wide range of Reynolds number regimes.
In the context of experimental studies of the flow over spheres, one of the main challenges is fixing the sphere in a wind or water tunnel test section with minimal disturbance of the flow, since more often than not, the mounting/support structure of the sphere can have a significant influence on the flow. The supports chosen for an experiment should simultaneously provide rigid and stable positioning of the sphere, and minimise the disturbance to the flow, while also facilitating the placement of probes without obstructing optical measurement access. Some of the supports used in these studies include rigid downstream supports [1], honeycomb structures with rigid supports upstream of the sphere [2, 3], and wires placed azimuthally [4, 5]. The drag on the spheres has been measured using hot-wire probes [1, 6], with flow visualisations performed using aluminium dust [7], smoke [8, 5], and dyes [9, 4, 10] to gain a qualitative understanding of the recirculation region, flow separation, and vortex shedding in the wake. Various forms of particle image velocimetry (PIV), including planar [5], stereographic [3] and tomographic [11], have been used to quantitatively measure the flow around spheres.
While mounting the sphere is not a concern in numerical studies, simulations are subject to their own set of challenges. The geometry of the problem and the range of scales involved require careful consideration of the mesh and solver to resolve the flow at high Reynolds numbers. Because of this complexity, large eddy simulation (LES[12, 13, 14, 15] and detached eddy simulation (DES[14, 16] methods have been favoured over the more computationally expensive direct numerical simulation (DNS[12, 15, 17, 18] to resolve the main features of the flow at high Reynolds numbers. In conjunction with the advancement of computational capabilities, these models have extended the feasibility of numerical investigations of the flow over spheres to large Reynolds numbers.
Experimental and numerical studies have examined the development of the flow over spheres over a wide range Reynolds number regimes. At low Reynolds number, $Re_{D}<20$, the flow in the wake is laminar and axisymmetric, and there is no recirculation [7, 13]. A small vortex ring forms near the rear stagnation point around $Re_{D}\approx 25$ [7], and elongates in the streamwise direction with increasing Reynolds number. This vortex ring begins to oscillate around $Re_{D}\approx 130$ with a long period, which becomes more prominent at higher Reynolds numbers. The recirculation region formed behind the sphere, which exhibits reverse flow in the mean, extends further downstream of the sphere as Reynolds number increases [9].
The axisymmetry in the wake at low Reynolds number is lost due to the formation of a double-threaded wake around $Re_{D}\approx 210$ [9, 13], which consists of two streamwise vortices of opposing sign that can be seen as dye threads in the wake of a free-falling sphere [19]. Single-frequency vortex shedding begins around $Re_{D}\approx 300$ [7, 4], and hairpin-shaped vortices form at $Re_{D}\approx 400$ [10]. The generation of an unsteady random side force, which has been observed in both experimental [8] and numerical [20] studies results in the loss of planar symmetry in the wake in the range of $400<Re_{D}<500$ [7].
From $Re_{D}=800$, the small-scale Kelvin-Helmholtz instability in the shear layer at the edge of the recirculation region manifests as axisymmetric vortex shedding [16, 21]. This is the higher frequency of two instability modes identified in the wake of the sphere by [22], with the lower frequency mode arising from large-scale vortex shedding. Small-scale vortical structures are shed from the main vortical structure around $Re_{D}\approx 1000$ [20], and the recirculation region behind the sphere begins to shrink with increasing Reynolds number. Periodic fluctuations are present in the wake for much higher Reynolds numbers, $Re_{D}=O(10^{6})$ [10].
Since its first application to turbulent flows by [23], proper orthogonal decomposition (POD) has been widely used for identifying and analysing structures in turbulent flows [24, 25]. As POD does not require temporal resolution of the flow and can thus be applied to non-time-resolved data, it is a robust and simple method for identifying the structures within a turbulent flow. In the case of the flow over a cylinder, the turbulent fluctuations due to the vortices shed from the sphere can be identified in the leading POD modes [26, 27]. The energy of these modes is significantly larger than the higher-order modes, and their phase angle is clearly related to that of the pressure fluctuations resulting from the shed vortices [28]. Similarly, periodic structures in screeching, impinging and free jets [29, 30, 31], are represented by the leading POD modes. Identifying pairs of POD modes which represent periodic structures allows for the a posteriori phase-averaging of experimental data which is collected asynchronously or at low frequency relative to these structures. However, the complexity of the flow over a sphere in this regime makes identification of periodic structures in the turbulent fluctuations challenging, as POD modes representing different phases of a structure are not linked by the decomposition.
The analytic signal of the turbulent fluctuations, which is defined using the Hilbert transform, can be decomposed into complex modes, which include phase information, using POD [32]. While the Hilbert transform is performed in time for time-resolved data, it can also performed in an analogous spatial direction when the data is not time-resolved [33, 34, 35]. This results in modes emphasising structures which are periodic in the chosen direction, with the streamwise direction being the most sensible choice to identify structures which propagate downstream from the sphere. Although a useful method for identifying propagating structures, Hilbert POD (HPOD) introduces artefacts to the modes as a result of the Hilbert transform.
The current work applies both POD and HPOD, with the Hilbert transform performed in the streamwise direction, to experimentally-measured velocity fluctuations in the wake of a sphere at $Re_{D}=7780$. The modes resulting from each of these decompositions are compared, and the correlation between the POD modes and the HPOD modes is used to determine pairs of POD modes representing propagating structures. The data is then phase-averaged using the elucidated POD mode pairs to observe the periodic behaviour of the propagating structures. By pairing POD modes using the HPOD modes, the dynamics of the turbulent structures within the wake can be determined while avoiding the filtering that is introduced by the Hilbert transform. Additionally, the same mode pairing can be achieved using the streamwise Hilbert transform of the individual POD modes. This method significantly reduces the computational cost of identifying paired POD modes, as it does not require POD to be performed on the complex-valued analytic signal. It is shown that the leading modes of POD and HPOD are similar in shape, and have comparable phase angles throughout the experimental data. However, the HPOD modes accentuate the propagating parts of the structures and introduce artefacts which are not present in the POD modes.
2 Theoretical Considerations
2.1 Proper Orthogonal Decomposition
POD decomposes statistically independent realisations of the random field $\boldsymbol{f}(\boldsymbol{x},t)$ into orthonormal spatial modes $\psi_{k}(\boldsymbol{x})$ with temporal coefficients $a_{k}(t)$, which are ranked by their energy $\lambda_{k}$, from largest to smallest, such that
$\boldsymbol{f}(\boldsymbol{x},t)=\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x}),$ (1)
and is analogous to singular value decomposition (SVD) and principle component analysis (PCA). In the application of POD to turbulence, $\boldsymbol{f}(\boldsymbol{x},t)$ represents the turbulent fluctuations, and $\lambda_{k}$ represents the contribution of the $k^{th}$ mode to the total TKE of $\boldsymbol{f}(\boldsymbol{x},t)$. A common use for POD is the definition of a reduced-order representation of $\boldsymbol{f}(\boldsymbol{x},t)$ by reconstructing the original data using $\hat{K}<K$ modes
$\hat{\boldsymbol{f}}(\boldsymbol{x},t)=\sum_{k=1}^{\hat{K}}a_{k}(t)\psi_{k}(\boldsymbol{x}),$ (2)
which represents a portion of the total kinetic energy equal to
$TKE=100\times\frac{\sum_{k=1}^{\hat{K}}\lambda_{k}}{\sum_{k=1}^{K}\lambda_{k}}\ \%.$ (3)
In order to perform POD, the two-point correlation of $\boldsymbol{f}(\boldsymbol{x},t)$ in space
[…]
is calculated, where $T$ is the temporal domain. The spatial modes and the corresponding energy are given by the eigendecomposition of $C(\boldsymbol{x},\boldsymbol{x}^{\prime})$
[…]
where $S$ is the spatial domain. The temporal coefficients are the projection of $\boldsymbol{f}(\boldsymbol{x},t)$ onto $\psi(\boldsymbol{x})$
[…]
Alternatively, POD can be performed using the method of snapshots, which first determines the temporal coefficients, and then the spatial modes. This method is more computationally efficient when the number of spatial coordinates included in $\boldsymbol{f}(\boldsymbol{x},t)$ is larger than the number of realisations of the field. The two-point correlation in time
[…]
is used instead of that in space, given in equation (4). The eigendecomposition of $C(t,t^{\prime})$
[…]
yields the eigenvalues and temporal coefficients. The spatial modes are then determined by the projection of $\boldsymbol{f}(\boldsymbol{x},t)$ onto $a(t)$
[…]
The temporal coefficients of the $i^{th}$ and $j^{th}$ POD modes can be expressed in terms of their phase angle $\phi_{i,j}$ as
$a_{i}(t)=\sqrt{2\lambda_{i}}\cos{\big{(}\phi_{i,j}(t)\big{)}},$ (10)
and
$a_{j}(t)=\sqrt{2\lambda_{j}}\sin{\big{(}\phi_{i,j}(t)\big{)}},$ (11)
respectively. Thus, the phase angle $\phi_{i,j}(t)$ is given by
$\phi_{i,j}(t)=\tan^{-1}\left(\frac{a_{j}(t)/\sqrt{2\lambda_{j}}}{a_{i}(t)/\sqrt{2\lambda_{i}}}\right).$ (12)
The phase average of the data using the $i^{th}$ and $j^{th}$ POD modes at the phase angle $\phi$ is given by
$\widetilde{\boldsymbol{f}}(\boldsymbol{x},\phi)=\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})\,|\,\phi_{i,j}(t)=\phi$ (13)
where $\phi_{i,j}(t)$ is defined by equation (12). While any two POD modes can be used for phase-averaging $\boldsymbol{f}(\boldsymbol{x},t)$, not all selections of modes result in physically meaningful phase averages. In the case of strongly dominant periodic fluctuations, the temporal coefficients of the corresponding modes form a circle [26]. Large periodic structures can also be identified by comparing the phase angle of the modes to an independent measurement of the phase of the flow, such as pressure signals [28]. However, the less prominent the periodic structures, the more difficult it is to relate the POD modes which represent them.
2.2 Hilbert Proper Orthogonal Decomposition
HPOD embeds the phase information in the modes by replacing $\boldsymbol{f}(\boldsymbol{x},t)$ with its analytic signal [32]
$\boldsymbol{f}^{a}(\boldsymbol{x},t)=\boldsymbol{f}(\boldsymbol{x},t)+i\mathcal{H}[\boldsymbol{f}(\boldsymbol{x},t)],$ (14)
where $\mathcal{H}$ denotes the Hilbert transform of $\boldsymbol{f}(\boldsymbol{x},t)$ in time
[…]
The Hilbert transform is equivalent to a $\mp 90^{\circ}$ phase shift [36] of the fundamental frequencies of $\boldsymbol{f}(\boldsymbol{x},t)$ in Fourier space
$\mathcal{H}\big{[}\boldsymbol{f}(\boldsymbol{x},t)\big{]}=\mathcal{F}^{-1}\Big{[}-i\operatorname{sgn}(f)\mathcal{F}\big{[}\boldsymbol{f}(\boldsymbol{x},t)\big{]}\Big{]},$ (16)
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ are the Fourier transform and inverse Fourier transform, respectively, and
[…]
is the sign of the instantaneous frequency $f$ of each term in the Fourier series given by $\mathcal{F}\big{[}\boldsymbol{f}(\boldsymbol{x},t)\big{]}$.
For time-resolved data, the Hilbert transform can be performed in time to determine the analytic signal $\boldsymbol{f}^{a}(\boldsymbol{x},t)$. HPOD is performed on time-resolved data by replacing $\boldsymbol{f}(\boldsymbol{x},t)$ with its analytic signal in equations (4) to (6). However, where the coherent structures propagate downstream, the Hilbert transform can be performed in the streamwise direction, as propagation in this direction serves as an analogy for a change in time, to apply this method to data that is not time-resolved [33, 34, 35]
[…]
In this case, the analytic signal of $\boldsymbol{f}(\boldsymbol{x},t)$ is defined using $\mathcal{H}_{x}$
$\boldsymbol{f}^{a}_{x}(\boldsymbol{x},t)=\boldsymbol{f}(\boldsymbol{x},t)+i\mathcal{H}_{x}\big{[}\boldsymbol{f}(\boldsymbol{x},t)\big{]},$ (19)
and substituted for $\boldsymbol{f}(\boldsymbol{x},t)$ in equations (7) to (9). The resulting eigenvalues $\lambda^{a}_{k}$ are the contribution of the $k^{th}$ mode to the total TKE of the analytic signal, and are real-valued. The spatial modes $\psi^{a}_{k}(\boldsymbol{x})$ are complex, with the imaginary part of $\psi^{a}_{k}(\boldsymbol{x})$ representing a phase shift of its real part by $\mp 90^{\circ}$.
The instantaneous amplitude and phase angle of the $k^{th}$ mode are defined by the real $\Re$ and imaginary $\Im$ parts of $a^{a}_{k}(t)$ as
$|a^{a}_{k}(t)|=\sqrt{\Re{\big{(}a^{a}_{k}(t)\big{)}}^{2}+\Im{\big{(}a^{a}_{k}(t)\big{)}}^{2}},$ (20)
and
$\phi^{a}_{k}(t)=\tan^{-1}\left(\mp\frac{\Im\big{(}a^{a}_{k}(t)\big{)}}{\Re\big{(}a^{a}_{k}(t)\big{)}}\right),$ (21)
respectively. The instantaneous amplitude denotes the strength of the mode at each instant in time, and the instantaneous phase angle denotes the mode shape, which oscillates between the real and imaginary parts of $\psi_{k}^{a}(\boldsymbol{x})$ at $\phi_{k}^{a}(t)$ equal to $0^{\circ}$ and $\mp 90^{\circ}$, respectively.
The analytic signal can be reconstructed using $\hat{K}$ modes by
$\hat{\boldsymbol{f}^{a}}(\boldsymbol{x},t)=\sum_{k=1}^{\hat{K}}a^{a}_{k}(t)\psi^{a}_{k}(\boldsymbol{x}),$ (22)
the real part of which is the reduced-order model of $\boldsymbol{f}(\boldsymbol{x},t)$ including the modes which contribute the most TKE to the analytic signal. This reconstruction favours the propagating components of $\boldsymbol{f}(\boldsymbol{x},t)$, but introduces artefacts due to the Hilbert transform.
2.3 Hilbert Transform of POD Modes
Although the POD modes do not contain temporal information, the streamwise Hilbert transform can be performed directly on each of the modes. Thus, the analytic signal of the POD mode $\psi_{k}(\boldsymbol{x})$ can be defined as
$(\psi_{k})^{a}(\boldsymbol{x})=\psi_{k}(\boldsymbol{x})+i\mathcal{H}_{x}\big{[}\psi_{k}(\boldsymbol{x})\big{]}.$ (23)
Substituting the analytic signal of the modes for the POD modes in equation (1) yields
$\sum_{k=1}^{K}a_{k}(t)\Big{(}(\psi_{k})^{a}(\boldsymbol{x})\Big{)}=\sum_{k=1}^{K}a_{k}(t)\Big{(}\psi_{k}(\boldsymbol{x})+i\mathcal{H}_{x}\big{[}\psi_{k}(\boldsymbol{x})\big{]}\Big{)},$ (24)
which can be expanded to
$\sum_{k=1}^{K}a_{k}(t)\Big{(}(\psi_{k})^{a}(\boldsymbol{x})\Big{)}=\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})+a_{k}(t)i\mathcal{H}_{x}\big{[}\psi_{k}(\boldsymbol{x})\big{]}.$ (25)
As the Hilbert transform is a linear operator, this can be rewritten as
[…]
which is the analytic signal of $\boldsymbol{f}(\boldsymbol{x},t)$.
Therefore, applying the transform to each of the POD modes of $\boldsymbol{f}(\boldsymbol{x},t)$ individually and using the resulting analytic signals in place of the original modes allows for the reconstruction of $\boldsymbol{f}^{a}(\boldsymbol{x},t)$. However, the analytic signal of the POD modes are not equal to the HPOD modes due to the filtering of the Hilbert transform being introduced after the decomposition of $\boldsymbol{f}(\boldsymbol{x},t)$. While this method is not equivalent to HPOD, where the propagation of structures is embedded in the modes, it is a computationally efficient alternative for matching pairs of POD modes which represent the same propagating structure.
3 Methodology
3.1 Experimental Facility
Experiments were performed in a vertical water tunnel, the test section of which is 1.5 m long with a cross-section of 0.25 m $\times$ 0.25 m, shown in figure 1[37, 38, 39]. The settling chamber is connected to the top of the test section by a 16:1 contraction, which includes a honeycomb and screens to reduce the turbulence level and scale in the tunnel test section. Water is pumped from the plenum chamber below the test section up to the settling chamber such that the flow through the test section is downwards. The walls of the test section are made of 15 mm thick clear acrylic. A removable panel is located on one wall for access to the test section, and is set flush to the internal side of the wall to minimise asymmetry in the tunnel. Seeding and filtration of the tunnel is performed via an auxiliary circuit, which is isolated during experiments. A sting of diameter 9.3 mm is mounted to a crossbeam of thin aerofoil cross-sections above the test section, and tapers in diameter over the last 100 mm down to 4.65 mm in diameter and ends in a threaded rod which is used to mount a sphere in the test section. A close-up view of the sphere mounting structure is shown in figure 1b.
A sphere with a diameter of $D=40$ mm was manufactured using a Phrozen Sonic Mighty 8k 3-D printer, which has a 35 µm layer height and a planar resolution of 22 µm $\times$ 22 µm, as shown in figure 1c. The sphere has a shell thickness of 2.5 mm and an external boss at the top for mounting to a threaded rod on the end of the sting. The boss begins parallel to the sting and smoothly transitions to the sphere surface. The sphere is supported internally by a conical structure, which features slits to prevent uncured resin from being trapped in the cavities during printing.
Figure 1: (a) Vertical water tunnel facility, and (b) close-up of the mounting structure and sphere, and (c) 3-D printed sphere design [39].
3.2 Experimental Method
Measurements of the flow in the wake of the sphere were taken with a freestream velocity of $U_{\infty}=200$ m/s, resulting in a Reynolds number of $Re_{D}=7780$. Under these conditions, the background turbulence is 2%, and the mean convective time of the flow
$t_{c}=\frac{D}{U_{\infty}},$ (27)
is 0.2 seconds [39]. The flow was seeded with 11 µm glass spheres, and the single-exposed particle image pairs were recorded using a PCO Panda 26 DS camera with a 5120 $\times$ 5120 px array of 2.5 µm $\times$ 2.5 µm pixels. A Zeiss lens with a fixed 50 mm focal length was used to capture a 190 mm $\times$ 190 mm field of view, resulting in a spatial resolution is 37 µm/px. The field of view began just behind the trailing edge of the sphere, to avoid reflections off the surface, and extended to 5.2 diameters downstream of the sphere. The field of view was aligned with the centreline of the sphere, and extended to 2.3 diameters either side of the centreline in the transverse direction. The f-stop was set to 2.8, resulting in a 5.0 mm depth of field. The diffraction limited minimum image diameter and expected particle image diameter for an 11 µm particle at these setting are 3.9 µm and 1.6 px, respectively.
The particle images were illuminated using 7 ns pulses from a pair of 400mJ Spectra Physics Nd:YAG lasers. The beams of these lasers were aligned co-linearly before being shaped into a sheet aligned to the centre of the tunnel and the corresponding meridian of the sphere. The thickness of the laser sheet throughout the field of view was 300 µm. Pairs of single-exposed images where captured with a separation of $\Delta t=1$ ms at an acquisition rate of 1 Hz ($\Delta T=1$ s), using a BeagleBone Black pulse generator [40] to control the timing and synchronisation of the camera and laser pulses.
3.3 Data Processing
The mean background images of the first and second exposures were subtracted from the corresponding single-exposed particle image pairs. The instantaneous streamwise and transverse velocities, $u(x,y,t)$ and $v(x,y,t)$, were determined from the particle image pairs using two-component-two-dimensional (2C-2D) multigrid/multipass cross-correlation digital (MCCD) PIV [41, 42]. A final interrogation window of 64 px $\times$ 32 px on a grid with spacing of 32 px $\times$ 16 px, in $x$ and $y$, respectively, was used. This aspect ratio was chosen to capture the higher streamwise velocity component and the higher velocity gradients in the transverse direction, while still being sufficiently large for the interrogation window to contain approximately 10 particles common to both images of each pair. The quality of the cross-correlation was increased through the application of an 8-point Hart filter. A maximum displacement limit of 20 px and normalised local median filter [43], with a threshold of 2 standard errors, was applied to the raw displacement vectors. The validation criteria resulted in a rejection of 2.3% of vectors, which were replaced by an interpolation of a 2nd-order polynomial fit of their 13 nearest neighbours [42]. A total of $N_{T}=12,600$ two-component velocity vector fields, each containing 50,000+ vectors, were determined. The instantaneous velocity fluctuations $u^{\prime}(x,y,t)$ and $v^{\prime}(x,y,t)$ were calculated by subtracting the mean velocity field, $\overline{u}(x,y)$ and $\overline{v}(x,y)$, from the instantaneous velocity field, $u(x,y,t)$ and $v(x,y,t)$.
The 95% confidence interval of the 2C-2D MCCDPIV algorithm used is 0.06 pixels [42], which corresponds to a standard error of $\sigma_{\varepsilon_{u}}=0.03$ pixels. The standard error of the instantaneous velocities is thus $1.1\times 10^{-3}$ m/s. The uncertainty of the mean velocity arising from the random nature of the turbulent fluctuations can be estimated by including the mean streamwise Reynolds stress [44, 45]
$\sigma_{\overline{u}}=\sqrt{\frac{\overline{u^{\prime}u^{\prime}}+\sigma^{2}_{\varepsilon_{u}}}{N_{T}}},$ (28)
where $\sigma_{\varepsilon_{u}}$ is the standard error of the instantaneous streamwise velocity measurements. As the maximum of both $\overline{u^{\prime}u^{\prime}}$ and $\overline{v^{\prime}v^{\prime}}$ are of the order of $0.05U_{\infty}^{2}$ [39], a conservative estimate of the uncertainty in the mean velocity components is $4.0\times 10^{-4}$ m/s. The uncertainty in the fluctuations is thus
$\sigma_{u^{\prime}}=\sqrt{\sigma^{2}_{\overline{u}}+\sigma^{2}_{\varepsilon_{u}}},$ (29)
which is equal to $1.2\times 10^{-3}$ m/s. The key experimental and processing parameters are summarised in table 1.
Parameter Symbol Units Relative Units
Freestream velocity $U_{\infty}$ 200 mm/s --
Sphere diameter $D$ 40 mm --
Reynolds number $Re_{D}$ 7780 --
Convective time $t_{c}$ 0.2 s --
Time between laser pulses $\Delta t$ 1 ms $0.005t_{c}$
Time between velocity fields $\Delta T$ 1 s $5t_{c}$
Depth of field -- 5.0 mm $0.125D$
Laser sheet thickness -- 0.3 mm $0.0075D$
Diffraction limited particle diameter -- 3.9 µm
Spatial resolution $SR$ 37 µm --
Grid spacing -- 32 px $\times$ 16 px $0.03D\times 0.015D$
Interrogation window -- 64 px $\times$ 32 px $0.06D\times 0.06D$
Field of view -- 190 mm $\times$ 190 mm $4.7D\times 4.7D$
Number of 2C-2D velocity fields $N_{T}$ 12,600 --
Velocity vectors per field $N_{S}$  50,000+ --
Velocity vector validation rate -- 97.7%
Instantaneous velocity fluctuation uncertainty $\sigma_{u^{\prime}}$ $1.2\times 10^{-3}$ m/s $0.006U_{\infty}$
Table 1: Summary of experimental and processing parameters.
4 Analysis and Results
4.1 POD and HPOD Modes
POD was performed by arranging the turbulent fluctuations into the matrix
$\boldsymbol{X}=\begin{bmatrix}u^{\prime}(\boldsymbol{x}_{1},t_{1})&\cdots&u^{\prime}(\boldsymbol{x}_{N_{S}},t_{1})&v^{\prime}(\boldsymbol{x}_{1},t_{1})&\cdots&v^{\prime}(\boldsymbol{x}_{N_{S}},t_{1})\\ \vdots&u^{\prime}(\boldsymbol{x}_{n},t_{m})&\vdots&\vdots&v^{\prime}(\boldsymbol{x}_{n},t_{m})&\vdots\\ u^{\prime}(\boldsymbol{x}_{1},t_{N_{T}})&\cdots&u^{\prime}(\boldsymbol{x}_{N_{S}},t_{N_{T}})&v^{\prime}(\boldsymbol{x}_{1},t_{N_{T}})&\cdots&v^{\prime}(\boldsymbol{x}_{N_{S}},t_{N_{T}})\\ \end{bmatrix}$ (30)
where $N_{S}$ is the number of velocity vectors per velocity field and $\boldsymbol{x}_{n}={(x,y)}_{n}$ is the spatial coordinates of the $n^{th}$ vector, which has fluctuating components denoted by $u^{\prime}$ and $v^{\prime}$ in the streamwise and transverse directions, respectively. As $N_{S}>N_{T}$, the method of snapshots was used to perform the POD, with the two-point correlation of $\boldsymbol{X}$ in time given by the matrix form of equation (7)
$\boldsymbol{C}=\boldsymbol{X}\boldsymbol{X}^{T},$ (31)
the eigendecomposition of which
$\boldsymbol{C}\boldsymbol{\Phi}=\lambda\boldsymbol{\Phi},$ (32)
yielded the eigenvectors $\boldsymbol{\Phi}$ and eigenvalues $\lambda$. The temporal coefficients were determined from $\boldsymbol{\Phi}$
$a_{k}(t_{m})=\boldsymbol{\Phi}_{m,k},$ (33)
and the spatial modes were given by the projection of $\boldsymbol{X}$ onto $\boldsymbol{\Phi}$
$\psi_{k}(\boldsymbol{x}_{n})=(\boldsymbol{\Phi}^{T}\boldsymbol{X})_{k,n}.$ (34)
The analytic signal of $\boldsymbol{X}$
$\boldsymbol{X}^{a}=\boldsymbol{X}+i\mathcal{H}_{x}[\boldsymbol{X}],$ (35)
was determined using equations (18) and (19), and HPOD was performed by substituting $\boldsymbol{X}^{a}$ for $\boldsymbol{X}$ in equations (31) to (34). The resulting eigenvalues, temporal coefficients and spatial modes are denoted by $\lambda^{a}_{k}$, $a_{k}^{a}(t)$ and $\psi_{k}^{a}(\boldsymbol{x})$, respectively.
The contribution of the leading HPOD modes to the TKE of $\boldsymbol{X}^{a}$ is significantly larger than the contribution of the leading POD modes to the TKE of $\boldsymbol{X}$. The first, second and third modes of each decomposition contribute 7.70%, 4.53% and 3.70%, and 4.8%, 3.39% and 2.55%, respectively. However, the difference in the contribution of the higher order modes to the TKE is less significant, as shown in figure 2a. The result is that the cumulative contribution to TKE of a given number of the HPOD modes is larger than that of the equivalent number of POD modes, as shown in figure 2b and table 2. However, this is not a one-to-one comparison, as the total TKE of the analytic signal $\boldsymbol{X}^{a}$ is larger than that of the data $\boldsymbol{X}$, at 571,871 m2/s2 and 300,542 m2/s2, respectively. This is due to the imaginary component of $\boldsymbol{X}^{a}$, which is not present in $\boldsymbol{X}$.
Figure 2: (a) Individual and (b) cumulative TKE contribution of the POD and HPOD modes.
TKE [%] POD Modes HPOD Modes Ratio
25 12 7 0.58
50 65 38 0.58
75 311 180 0.58
90 1027 593 0.58
95 1973 1172 0.59
99 5904 4148 0.70
Table 2: Number of POD modes and HPOD modes required to meet 25% to 99% of the TKE of the data $\boldsymbol{X}$ and its analytic signal $\boldsymbol{X}^{a}$, respectively, and the ratio of HPOD modes to POD modes required to meet each proportion of TKE.
Figure 3: (i) Streamwise and (ii) transverse components of the first ten POD modes.
[…]
The sixth and eighth POD modes, shown in figures 3f and 3h, respectively, are similar in structure to the first and second POD modes but with a shorter wavelength than those modes. The seventh POD mode superficially resembles the fourth mode in its transverse component, shown in figure 3g(ii). However, the streamwise component, shown in figure 3g(i), does not exhibit the streamwise periodicity of the other leading modes. While this mode may represent an asymmetric aspect of the structure represented by the forth and fifth modes, it is less clearly related to these modes than the third mode appears to be to the first and second modes. The ninth and tenth POD modes respectively exhibit V-shaped structures in their streamwise components, shown in figures 3i(i) and 3j(i), respectively, similar to the fourth and fifth modes for $x/D>2.5$, which increase in wavelength further downstream. These modes differ in shape from the lower-order modes upstream of $x/D=2$, where the structures have an aspect ratio closer to unity than those in the higher-order modes. The transverse components of the ninth and tenth modes, shown in figures 3i(ii) and 3j(ii), respectively, resemble those of the fourth and fifth modes with a short wavelength. The wavelength of these modes also increase in the downstream direction.
Figure 4: (i) Streamwise and (ii) transverse components of the first four HPOD modes.
The first HPOD mode, the real and imaginary parts of which are shown in figure 4a and 4b, respectively, is qualitatively similar in structure to the first two POD modes, shown in figure 3a and 3b, respectively. The streamwise components of the first HPOD mode, shown in figures 4a(i) and 4b(i) for the real and imaginary parts, respectively, consist of alternating positive and negative regions in the streamwise direction and is antisymmetric about the streamwise axis. The transverse components, shown in figures 4a(ii) and 4b(ii) for the real and imaginary parts, respectively, consist of symmetric regions of alternating sign. This structure is very similar to that of the first two POD modes, with the exception of the ends of the measurement domain. This difference is due to the filtering of the Hilbert transform, which places more emphasis on the components of the structure which propagate in the streamwise direction. The effect of this filtering is seen most clearly comparing figures 4b(i) and 3b(i) around $x/D=5$. The second HPOD mode features the same V-shaped structure in the streamwise components, the real and imaginary parts of which are shown in figures 4c(i) and 4d(i), respectively, present in the fourth and fifth POD modes, shown in figures 3d(i) and 3e(i), respectively. The transverse components of the real and imaginary parts of the second HPOD mode are antisymmetric with regions of alternating sign in the streamwise direction, as shown in figures 4c(ii) and 4d(ii), respectively. While similar to transverse components of the fourth and fifth POD modes, shown in figures 3d(ii) and 3e(ii), respectively, the HPOD modes are significantly more symmetrical/anti-symmetrical than the POD modes, due to the filtering effect of the Hilbert transform.
The third HPOD mode, the real and imaginary components of which are shown in figures 4e and 4f, have a similar structure to the sixth and eight POD modes, shown in figures 3f and 3h, respectively. Namely, antisymmetric streamwise velocity regions and symmetric transverse velocity regions of alternating sign in the streamwise direction, with a shorter wavelength than the first HPOD mode and the first two POD modes. The fourth HPOD mode, the real and imaginary components of which are shown in figures 4g and 4h, have a similar structure to the ninth and tenth POD modes, shown in figures 3i and 3j, respectively. The streamwise components of these modes feature symmetrical ovular regions of alternating sign in the streamwise direction in the near wake becoming V-shaped structures around $x/D=2$, while the transverse components feature antisymmetric regions of alternating sign in the streamwise direction.
4.2 Pairing POD Modes using HPOD Modes
As $\boldsymbol{f}(\boldsymbol{x},t)$ is given by the real part of $\boldsymbol{f}^{a}(\boldsymbol{x},t)$, it can be expressed in terms of the HPOD modes as
$\boldsymbol{f}(\boldsymbol{x},t)=\Re\big{(}\boldsymbol{f}^{a}(\boldsymbol{x},t)\big{)}=\sum_{k=1}^{K}\Re\big{(}a_{k}^{a}(t)\psi^{a}_{k}(\boldsymbol{x})\big{)},$ (36)
which can be expanded to
$\boldsymbol{f}(\boldsymbol{x},t)=\Re\big{(}\boldsymbol{f}^{a}(\boldsymbol{x},t)\big{)}=\sum_{k=1}^{K}\Re\big{(}a_{k}^{a}(t)\big{)}\Re\big{(}\psi^{a}_{k}(\boldsymbol{x})\big{)}-\Im\big{(}a_{k}^{a}(t)\big{)}\Im\big{(}\psi_{k}^{a}(\boldsymbol{x})\big{)}.$ (37)
As the real and imaginary parts of $a_{k}^{a}(t)$ can be expressed as $|a^{a}_{k}(t)|\cos(\phi^{a}_{k}(t))$ and $-|a^{a}_{k}(t)|\sin(\phi^{a}_{k}(t))$, respectively, the shape of the $k^{th}$ HPOD mode at the phase angle $\phi_{k}^{a}(t)$ is given by
$\frac{\psi_{k}^{a}(\boldsymbol{x},\phi_{k}^{a}(t))}{|a_{k}^{a}(t)|}=\cos\big{(}\phi_{k}^{a}(t)\big{)}\Re\big{(}\psi_{k}^{a}(\boldsymbol{x})\big{)}+\sin\big{(}\phi_{k}^{a}(t)\big{)}\Im\big{(}\psi_{k}^{a}(\boldsymbol{x})\big{)}.$ (38)
The mode shape of the $k^{th}$ HPOD mode matches the $j^{th}$ POD mode when
$\psi_{j}(\boldsymbol{x})=\begin{bmatrix}\cos(\Delta\phi^{a}_{k,j})&\sin(\Delta\phi^{a}_{k,j})\\ \end{bmatrix}\begin{bmatrix}\Re\big{(}\psi^{a}_{k}(\boldsymbol{x})\big{)}\\ -\Im\big{(}\psi^{a}_{k}(\boldsymbol{x})\big{)}\\ \end{bmatrix},$ (39)
where $\Delta\phi_{k,j}^{a}$ is the phase offset of the $k^{th}$ HPOD mode which results in the best match of the $j^{th}$ POD mode. The negative of the imaginary part is used to account for the negative introduced by the Hilbert transform for positive instantaneous frequencies, given in equation (16), which results in the phase of the HPOD modes being reversed relative to the motion of the structure.
The phase angles at which the $k^{th}$ HPOD mode matches each of the POD modes can be determined simultaneously by
$\begin{bmatrix}\cos(\Delta\phi^{a}_{k,1})&\sin(\Delta\phi^{a}_{k,1})\\ \vdots&\vdots\\ \cos(\Delta\phi^{a}_{k,j})&\sin(\Delta\phi^{a}_{k,j})\\ \vdots&\vdots\\ \cos(\Delta\phi^{a}_{k,K})&\sin(\Delta\phi^{a}_{k,K})\\ \end{bmatrix}^{T}={\left(\begin{bmatrix}\Re\big{(}\psi^{a}_{k}(\boldsymbol{x})\big{)}\\ -\Im\big{(}\psi^{a}_{k}(\boldsymbol{x})\big{)}\\ \end{bmatrix}\begin{bmatrix}\Re\big{(}\psi^{a}_{k}(\boldsymbol{x})\big{)}\\ -\Im\big{(}\psi^{a}_{k}(\boldsymbol{x})\big{)}\\ \end{bmatrix}^{T}\right)}^{-1}\begin{bmatrix}\Re\big{(}\psi^{a}_{k}(\boldsymbol{x})\big{)}\\ -\Im\big{(}\psi^{a}_{k}(\boldsymbol{x})\big{)}\\ \end{bmatrix}\begin{bmatrix}\psi_{1}(\boldsymbol{x})\\ \vdots\\ \psi_{j}(\boldsymbol{x})\\ \vdots\\ \psi_{K}(\boldsymbol{x})\\ \end{bmatrix}^{T},$ (40)
and
$\Delta\phi^{a}_{k,j}=\tan^{-1}\left(\frac{\sin(\Delta\phi^{a}_{k,j})}{\cos(\Delta\phi^{a}_{k,j})}\right).$ (41)
The correlation coefficient of the $k^{th}$ HPOD mode at $\Delta\phi^{a}_{k,j}$ and the $j^{th}$ POD mode is
$\mathcal{R}_{k,j}=\frac{\operatorname{cov}\big{(}\psi_{k}^{a}(\boldsymbol{x},\Delta\phi^{a}_{k,j}),\psi_{j}(\boldsymbol{x})\big{)}}{\sqrt{\operatorname{var}\big{(}\psi_{k}^{a}(\boldsymbol{x},\Delta\phi^{a}_{k,j})\big{)}\operatorname{var}\big{(}\psi_{j}(\boldsymbol{x},)\big{)}}},$ (42)
and the corresponding $R^{2}$ value is
$R^{2}_{k,j}=1-\frac{\sum\big{(}\psi_{k}^{a}(\boldsymbol{x},\Delta\phi^{a}_{k,j})-\psi_{j}(\boldsymbol{x})\big{)}^{2}}{\sum\big{(}\psi_{j}(\boldsymbol{x})\big{)}^{2}}.$ (43)
The HPOD modes which best match each of the first ten POD modes, and the corresponding phase offset, correlation coefficient, and $R^{2}$ value, are shown in table 3.
POD Mode, $j$ HPOD Mode, $k$ $\Delta\phi_{k,j}^{a}$ $\mathcal{R}_{k,j}$ $R^{2}_{k,j}$
1 1 $1.71\pi$ 0.89 0.78
2 1 $1.18\pi$ 0.97 0.93
3 3 $0.18\pi$ 0.62 0.26
4 2 $1.70\pi$ 0.84 0.65
5 2 $1.21\pi$ 0.96 0.91
6 3 $1.52\pi$ 0.82 0.65
7 4 $1.70\pi$ 0.63 0.29
8 3 $0.88\pi$ 0.76 0.50
9 4 $1.10\pi$ 0.85 0.70
10 4 $0.50\pi$ 0.80 0.49
Table 3: HPOD modes and phase shifts for best match of the first ten POD modes, and the corresponding $\mathcal{R}_{k,j}$ values.
The first and second POD modes are the best matches to the first HPOD mode, with correlation coefficients of 0.89 and 0.97, and $R^{2}$ values of 0.78 and 0.93, respectively. The corresponding phase angles are $1.71\pi$ and $1.18\pi$, which are approximately $90^{\circ}$ apart. This is consistent with the separation of the real and imaginary parts of the HPOD modes, suggesting that the relationship between these POD modes is similar to that of the real and imaginary parts of the first HPOD mode. Similarly, the fourth and fifth POD modes are the best matches to the second HPOD mode, with correlation coefficients of 0.84 and 0.96, and $R^{2}$ values of 0.65 and 0.91, respectively. The corresponding phase angles are $1.7\pi$ and $1.21\pi$, which are also separated by approximately $90^{\circ}$.
The sixth and eighth POD modes are the best match to the third HPOD mode, with correlation coefficients of 0.82 and 0.76, and $R^{2}$ values of 0.65 and 0.50, respectively. The separation of the corresponding phase angles, $1.52\pi$ and $0.88\pi$, respectively, is greater than $90^{\circ}$. This is also the case for the ninth and tenth POD modes, which are the best matches to the fourth HPOD mode at phase angles of $1.10\pi$ and $0.50\pi$, respectively. The corresponding correlation coefficients and $R^{2}$ values of these POD modes are 0.85 and 0.80, and 0.70 and 0.49, respectively.
The phase average of the paired POD modes were compared to the corresponding HPOD modes at phase angles from 0 to $2\pi$ in increments of $0.01\pi$, accounting for the phase angle shift in the first POD mode in each pair, $\Delta\phi_{k,i}$. The $R^{2}$ values of each phase average
$R_{k,(i,j)}^{2}=1-\frac{\sum{\Big{(}\psi^{a}_{k}(\boldsymbol{x},\phi+\Delta\phi^{a}_{k,i})-\widetilde{\psi_{i,j}}(\boldsymbol{x},\phi)\Big{)}}^{2}}{\sum{\big{(}\widetilde{\psi_{i,j}}(\boldsymbol{x},\phi)\big{)}}^{2}},$ (44)
was calculated for each case. The joint probability density functions of the phase angle of the HPOD modes and the phase angle of the corresponding pairs of POD modes, adjusted for the phase angle shift between the HPOD mode and the first POD mode of each pair.
HPOD Mode $k$ POD Modes $i,j$ $R_{k,(i,j)}^{2}$
1 1,2 0.81
2 4,5 0.76
3 6,8 0.57
4 9,10 0.61
Table 4: $R_{k,(i,j)}^{2}$ values of the phase-averaged HPOD modes and the corresponding POD modes.
Figure 5: Phase-averaged (a) streamwise and (c) transverse components of the first and second POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the first HPOD mode, adjusted for the phase angle offset $\Delta\phi_{1,1}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
The streamwise component of the phase-averaged first and second POD modes, shown on figure 5a, is good agreement with that of the phase average of the first HPOD mode at the adjusted phase angle, shown in figure 5b, with the main differences in the structures being a result of the filtering from the Hilbert transform. Namely, the intensity of the structures in the POD mode is higher directly behind the sphere than the HPOD modes, while the intensity in the structures further downstream is higher in the HPOD modes. This difference is also seen in the transverse comments, shown in figures 5c and 5d for the POD modes and HPOD modes, respectively.
Additionally, the structures in the near wake of the sphere in the POD modes, shown in figure 5a(iv), are filtered out of the HPOD mode, shown in figure 5b(iv), as they do not propagate downstream. Despite these differences, it is clear that the first and second POD modes and the first HPOD mode represent the same propagating structure consisting of planar antisymmetric streamwise fluctuations and alternating positive and negative transverse fluctuations. This structure is consistent with flapping of the wake, and the similarity of the third POD mode, shown in figure 3c, to the first and second suggests that the flapping also has an out-of-plane component.
Figure 6: Phase-averaged (a) streamwise and (c) transverse components of the fourth and fifth POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the second HPOD mode, adjusted for the phase angle offset $\Delta\phi_{2,4}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
While the streamwise component of the phase-averaged fourth and fifth POD modes, shown in figure 6a, is similar to that of the second HPOD mode, shown in figure 6b, the latter demonstrates significantly stronger symmetry, due to the inherent filtering of propagating features by HPOD. This is most clearly seen when comparing figures 6a(ii) and 6a(vi) to figures 6b(ii) and 6b(vi), respectively, where the V-shaped structures in the HPOD case remain symmetric downstream, while they become asymmetric in the case of the POD modes. The transverse component of the second HPOD mode, shown in figure 6d, shows a stronger antisymmetry than the transverse components of the phase average of the fourth and fifth POD modes, shown in figure 6c.
The variation between the phase average of the fourth and fifth POD modes and the second HPOD mode highlights the filtering effect of HPOD, which prioritises the structures that are periodic in the direction of the Hilbert transform. While the fourth and fifth POD modes are highly asymmetric, particularly after $x/D>3$, the second HPOD mode is symmetric in the streamwise component and antisymmetric in the transverse direction. The asymmetry in the POD modes is likely due to the three-dimensionality in the flow being more significant in the structure represented by the fourth and fifth POD modes than that represented by the first and second POD modes. In this case, the use of HPOD is beneficial for isolating the propagating structures in the flow in the presence of asymmetry in the POD modes. The alternating positive and negative streamwise fluctuations in these modes suggests that they represent a pulsating motion in the wake of the sphere.
Figure 7: Phase-averaged (a) streamwise and (c) transverse components of the sixth and eighth POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the third HPOD mode, adjusted for the phase angle offset $\Delta\phi_{3,6}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
The streamwise component of the phase average of the sixth and eight POD modes, shown in figure 7a, is qualitatively similar to that of the third HPOD mode, shown in figure 7b. However, the alternating positive and negative structures are better defined in the HPOD mode than the phase average of the POD modes. This is most evident when comparing figures 7a(i) and 7a(ii) to figures 7b(i) and 7b(ii), in which the positive region in the HPOD modes around $x/D=3$ and $y/D>0$ is not present in the phase-averaged POD modes. This is most likely due to the structure moving in the out-of-plane direction as it propagates, which is filtered out by the Hilbert transform. Additionally, the intensity of the structures is generally lower in the HPOD modes than those in the POD modes due to the filtering of the non-propagating components.
The transverse component of the phase average of the sixth and eighth POD modes, shown in figure 7c, shows a transverse shifting of the structures of alternating sign, particularly in figures 7c(i)-(iii). This shifting is not present in the transverse component of the third HPOD mode, shown in figure 7d. The shifting of the structures in the transverse velocity of the POD modes is consistent with the the three-dimensional motion of the wake, and coupled with the behaviour of the streamwise components of the POD modes. The structure represented by these modes appears to be a flapping in the wake with a shorter wavelength than that represented by the first and second POD modes and the first HPOD mode.
Figure 8: Phase-averaged (a) streamwise and (c) transverse components of the ninth and tenth POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the fourth HPOD mode, adjusted for the phase angle offset $\Delta\phi_{4,9}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
The streamwise component of the phase-averaged ninth and tenth POD modes, shown in figure 8a, is qualitatively similar to the phase average of the fourth HPOD mode, shown in figure 8b, for $x/D>2.5$. In this domain, the streamwise components of both the POD modes and HPOD mode exhibit V-shaped structures similar to those in the fourth and fifth POD modes and the second HPOD mode. For $x/D<2.5$, there is more detail in the structure of the POD modes, which is most noticeable in figures 8a(iii) and 8a(iv) compared to figures 8b(iii) and 8b(iv). As the V-shaped structures in the streamwise components are more prominent in the phase-averaged ninth and tenth POD modes than the fourth HPOD mode, it can be inferred that they do not propagate in the streamwise direction.
The transverse components of phase average of the ninth and tenth POD modes and the phase average of the fourth HPOD modes, shown in figures 8c and 8d, respectively, are also qualitatively similar. As is the case for the streamwise components, the transverse component of the phase-averaged POD modes contains features not present in the HPOD mode in the very near wake, due to the filtering effect of the Hilbert transform. This can be seen clearly in figures 8c(ii) and 8c(iii) compared to figures 8d(ii) and 8d(iii).
Figure 9: Joint probability density function of the phase angles of (a) the first HPOD mode and the first and second POD modes, (b) the second HPOD mode and the fourth and fifth POD modes, (c) the third HPOD mode and the sixth and eighth POD modes, and (d) the fourth HPOD mode and the ninth and tenth POD modes. Phase angles are relative to the phase angle of each mode in the first instantaneous velocity field.
The joint probability density functions of the HPOD modes and the corresponding POD mode pairs, shown in figure 9, are all the greatest around $\phi^{a}_{k}=\phi_{i,j}$. The phase angles of both have been adjusted by the phase angle of the first instantaneous velocity field to account for the phase shift between the POD modes and the HPOD modes and to ensure all phase angles are in the domain $\phi\in[0,2\pi]$.
4.3 Pairing POD Modes using the Hilbert Transform
While HPOD modes can be used to determine matched pairs of POD modes, this process requires performing both decompositions. As the Hilbert transform can be applied in the streamwise direction, it can be applied directly to the POD modes despite their lack of temporal information. The result is a shift of the fundamental wavenumbers of each POD mode by $\mp 90^{\circ}$, which is analogous to a propagation of the mode downstream. Thus, the streamwise Hilbert transform of a POD mode is analogous to a POD mode which represents a $\mp 90^{\circ}$ phase shift of the same streamwise periodic structure.
In order to determine matched pairs of POD modes using the Hilbert transform, the Hilbert transform of each POD mode $\psi_{k}(\boldsymbol{x})$ in the streamwise direction $\mathcal{H}_{x}[\psi_{k}(\boldsymbol{x})]$ is used in place of the imaginary part of the HPOD modes to define the analytic signal of $\psi_{k}(\boldsymbol{x})$
${(\psi_{k})}^{a}(\boldsymbol{x})=\psi_{k}(\boldsymbol{x})+i\mathcal{H}_{x}\big{[}\psi_{k}(\boldsymbol{x})\big{]}.$ (45)
The shape of the analytic signal of the $k^{th}$ POD mode at the phase angle $\phi_{k}$ is given by
$\big{(}\psi_{k}\big{)}^{a}(\boldsymbol{x},\phi_{k})=\cos(\phi_{k})\psi_{k}(\boldsymbol{x})+\sin(\phi_{k})\mathcal{H}_{x}[\psi_{k}(\boldsymbol{x})],$ (46)
and the phase angle at which the $k^{th}$ POD mode best matches the $j^{th}$ POD mode is given by
$\begin{bmatrix}\cos(\Delta\phi_{k,1})&\sin(\Delta\phi_{k,1})\\ \vdots&\vdots\\ \cos(\Delta\phi_{k,j})&\sin(\Delta\phi_{k,j})\\ \vdots&\vdots\\ \cos(\Delta\phi_{k,K})&\sin(\Delta\phi_{k,K})\\ \end{bmatrix}^{T}=\left(\begin{bmatrix}\psi_{k}(\boldsymbol{x})\\ \mathcal{H}_{x}[\psi_{k}(\boldsymbol{x})]\\ \end{bmatrix}\begin{bmatrix}\psi_{k}(\boldsymbol{x})\\ \mathcal{H}_{x}[\psi_{k}(\boldsymbol{x})]\\ \end{bmatrix}^{T}\right)^{-1}\begin{bmatrix}\psi_{k}(\boldsymbol{x})\\ \mathcal{H}_{x}[\psi_{k}(\boldsymbol{x})]\end{bmatrix}\begin{bmatrix}\psi_{1}(\boldsymbol{x})\\ \vdots\\ \psi_{j}(\boldsymbol{x})\\ \vdots\\ \psi_{K}(\boldsymbol{x})\\ \end{bmatrix}^{T},$ (47)
and
$\Delta\phi_{k,j}=\tan^{-1}\frac{\sin(\Delta\phi_{k,j})}{\cos(\Delta\phi_{k,j})}.$ (48)
The corresponding correlation coefficient is
$\mathcal{R}_{k^{a},j}=\frac{\operatorname{cov}\big{(}(\psi_{k})^{a}(\boldsymbol{x},\Delta\phi_{k,j}),\psi_{j}(\boldsymbol{x})\big{)}}{\sqrt{\operatorname{var}\big{(}(\psi_{k})^{a}(\boldsymbol{x},\Delta\phi_{k,j})\big{)}\operatorname{var}\big{(}\psi_{j}(\boldsymbol{x})\big{)}}}.$ (49)
The $R^{2}$ value of the phase average of the $k^{th}$ and $j^{th}$ POD modes compared to the phase average of the analytic signal of the $k^{th}$ POD mode is given by
$R^{2}_{k^{a},\widetilde{(k,j)}}=1-\frac{\sum\big{(}(\psi_{k})^{a}(\boldsymbol{x},\phi)-\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)\big{)}^{2}}{\sum\big{(}\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)\big{)}^{2}},$ (50)
where $(\psi_{k})^{a}(\boldsymbol{x},\phi)$ is the analytic signal of the $k^{th}$ POD mode and $\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)$ is the phase average of the $k^{th}$ and $j^{th}$ POD modes at phase angle $\phi$, which was determined using phase angles $\phi=[0,2\pi]$ in increments of $0.01\pi$. The POD mode which best matches the analytic signal of the each of the first ten POD modes, the corresponding correlation coefficient, and the $R^{2}$ of the phase averages are given in table 5.
POD Mode $k$ Paired POD Mode $j$ $\Delta\phi_{k,j}$ $\mathcal{R}_{k^{a},j}$ $R^{2}_{k^{a},\widetilde{(k,j)}}$
1 2 $-0.5\pi$ 0.88 0.89
2 1 $0.5\pi$ 0.76 0.76
3 2 $0.5\pi$ 0.70 0.74
4 5 $-0.5\pi$ 0.79 0.81
5 4 $0.5\pi$ 0.72 0.72
6 8 $-0.5\pi$ 0.80 0.82
7 9 $-0.5\pi$ 0.60 0.67
8 6 $0.5\pi$ 0.76 0.77
9 10 $-0.5\pi$ 0.64 0.65
10 9 $0.5\pi$ 0.63 0.65
Table 5: Paired POD modes using the analytic signal of the $k^{th}$ POD mode, and the corresponding phase angle, correlation coefficient and $R^{2}$ of the phase averages.
For each of the first ten POD modes, the paired POD mode corresponds to the analytic signal at a phase angle of $\mp 0.5\pi$. Thus, the $j^{th}$ POD mode best matches the Hilbert transform of the corresponding $k^{th}$ POD mode. A pair of POD modes representing a structure propagating in the streamwise direction therefore requires that the POD modes are a mutual best match with phase angle shifts of $0.5\pi$ with opposite signs. Thus, the POD mode pairs identified using the Hilbert transform are the first and second POD modes, the fourth and fifth POD modes, the sixth and eighth POD modes, and the ninth and tenth POD modes. These are the same pairings as those identified using the HPOD modes.
Figure 10: Phase-averaged (a) streamwise and (c) transverse components of the first and second POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the first POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
The streamwise component of the analytic signal of the first POD mode, shown in figure 10b, is a close match to the phase average of the first two POD modes, shown in figure 10a. Both show the same long regions of alternating sign propagating downstream from the sphere, with the main variation between them being for $x/D<2$, where the non-propagating features are located, as shown in figure 5a(iv). This is an artefact of applying the Hilbert transform to the POD mode, due to these structures being present in the second POD mode but not the first POD mode, on which the Hilbert transform was performed. This also explains the variation in the $R^{2}$ of the analytic signal of the first POD mode and the second POD mode and that of the analytic signal of the second POD mode and the first POD mode, being 0.89 and 0.76, respectively. This difference is due to the non-propagating components of these modes being absent in the other mode, but propagated by the Hilbert transform.
The variation between the transverse components of the phase-averaged first and second POD modes, shown in figure 10c, and the analytic signal of the first POD mode, shown in figure 10d, is significantly less than in the streamwise components. Both exhibit the same symmetric regions of alternating sign, which increase in length and decrease in intensity as they travel downstream. This is consistent with the vertical fluctuations from the flow coming around the sphere decaying once they are sufficiently far into the wake.
Figure 11: Phase-averaged (a) streamwise and (c) transverse components of the fourth and fifth POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the fourth POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
The streamwise component of the analytic signal of the fourth POD mode, shown in figure 11b, retains more of the asymmetry of the phase-averaged fourth and fifth POD modes, shown in figure 11a, than the second HPOD mode, shown in figure 6b. This is also true for the transverse component of the analytic signal of the fourth POD mode, shown in figure 11d. The transverse component of the analytic signal of the fourth POD mode, shown in figure 11d, also retains the asymmetry seen in the phase average of the fourth and fifth POD modes, shown in figure 11c. Although the filtering of the Hilbert transform is less prominent in the analytic signal of the fourth POD mode than in the second HPOD mode, shown in figures 6b and 6d, it is still present, most noticeably figures 11b(iii) and 11d(iv).
Figure 12: Phase-averaged (a) streamwise and (c) transverse components of the sixth and eighth POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the sixth POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
While there is some filtering in the analytic signal of the sixth POD mode, the asymmetry seen in the phase average of the sixth and eighth POD modes is preserved where the HPOD modes were planar symmetric. This can be seen in the positive region introduced to the streamwise component of the third HPOD mode at $x/D=2.5$, shown in figure 7b(i), which is not present in the phase average of the sixth and eight POD modes, shown in figure 12a(i), or the analytic signal of the sixth POD mode, shown in figure 12b(i). It is also shown in the asymmetric region at $x/D=3.5$ in the transverse component of the sixth and eighth POD modes, shown in figure 12c(i), which is also off-centre in the analytic signal of the sixth POD mode, shown in figure 12d(i), but located at the centreline in the third HPOD mode, shown in figure 7d(i).
Additionally, the intensity of the structures further downstream, such as in figure 12a(i), is clearly preserved in the analytic signal of the sixth POD mode, as shown in figure 12b(i), where the intensity of the same region is significantly reduced in the third HPOD mode, shown in figure 7b(i). This intensity in the phase average of the sixth and eighth POD modes around $x/D=4$, shown in figure 12a(v), is propagated downstream in the analytic signal of the sixth mode, as shown in figure 12b(v), whereas the intensity is lower in the same region in the third HPOD mode, as shown in figure 7b(v). This shows that the filtering which occurs from the Hilbert transform also affects the analytic signal of the POD modes.
Figure 13: Phase-averaged (a) streamwise and (c) transverse components of the ninth and tenth POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the ninth POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
While there is a significant difference between the streamwise components of the analytic signal of the ninth POD mode, shown in figure 13b, and that of the phase average of the ninth and tenth POD modes, shown in figure 13a, for $x/D<2$, they are more similar further downstream of the sphere. The structures in the phase average of the ninth and tenth modes are smaller scale for $x/D<2$ than they are for $x/D>2$, which is consistent with smaller scale fluctuations in the near wake breaking down as they travel downstream while the larger scales remain. As a result, these scales are not propagated by the Hilbert transform and are thus filtered in the analytic signal. The differences between the transverse components of the analytic signal of the ninth POD mode, shown in figure 13d, and the phase average of the ninth and tenth POD modes, shown in figure 13c, are significantly less prominent. However, there is still variation in the structures upstream of $x/D=1.5$, as shown in figure 13c(ii) and 13d(ii), and figures 13c(iii) and 13d(iii).
4.4 Propagating Structures in the Wake of a Sphere
The contribution of the pairs of POD modes identified to the total TKE of $\boldsymbol{X}$ is given in table 6.
HPOD Mode TKE [%] POD Modes TKE [%]
1 7.70 1, 2 8.24
2 4.53 4, 5 4.75
3 3.70 6, 8 3.20
4 2.70 9,10 2.60
Table 6: Turbulent kinetic energy contributions of the HPOD modes to $\boldsymbol{X}^{a}$ and of the POD mode pairs to $\boldsymbol{X}$.
In order to determine the nature and effects of the propagating structures identified from the POD modes, the paired POD modes were used to phase average the instantaneous velocity fluctuations
$\begin{bmatrix}\widetilde{u^{\prime}}(\boldsymbol{x},\phi)\\ \widetilde{v^{\prime}}(\boldsymbol{x},\phi)\\ \end{bmatrix}=\frac{1}{N_{\phi}}\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})\,|\,\phi_{i,j}=\phi,$ (51)
where $i$ and $j$ are paired POD modes, and $N_{\phi}$ is the number of instantaneous velocity fields included for the phase angle $\phi$. The mean velocity was added to the phase-averaged velocity fluctuations to determine the phase-averaged velocity
$\begin{bmatrix}\widetilde{u}(\boldsymbol{x},\phi)\\ \widetilde{v}(\boldsymbol{x},\phi)\\ \end{bmatrix}=\begin{bmatrix}\widetilde{u^{\prime}}(\boldsymbol{x},\phi)\\ \widetilde{v^{\prime}}(\boldsymbol{x},\phi)\\ \end{bmatrix}+\begin{bmatrix}\overline{u}(\boldsymbol{x})\\ \overline{v}(\boldsymbol{x})\\ \end{bmatrix},$ (52)
from which the phase-averaged vorticity is defined as
$\widetilde{\omega}(\boldsymbol{x},\phi)=\frac{\partial\widetilde{v}(\boldsymbol{x},\phi)}{\partial x}-\frac{\partial\widetilde{u}(\boldsymbol{x},\phi)}{\partial y}.$ (53)
The phase-averaged planar Reynolds stress is given by the phase average of the product of the streamwise and transverse components of the instantaneous velocity fluctuations. The TKE of the phase-averaged velocity is given by
$\widetilde{\operatorname{TKE}}(\boldsymbol{x},\phi)=\big{(}\widetilde{u^{\prime}}(\boldsymbol{x},\phi)\big{)}^{2}+\big{(}\widetilde{v^{\prime}}(\boldsymbol{x},\phi)\big{)}^{2}.$ (54)
Figure 14: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the first and second POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
Figure 15: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the first HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
The motion represented by the first and second POD modes, shown in figure 14, and the first HPOD mode, shown in figure 15, is a flapping of the wake. The phase-averaged velocity from using the first and second POD modes and the first HPOD mode, shown in figure 14a and 15a, respectively, are qualitatively indistinguishable, as is the phase-averaged vorticity, shown in figure 14b and 15b, respectively. The structures in the planar Reynolds stresses of the phase-average of POD modes, shown in figure 14c, are longer and more intense for $x/D<3$ than those in the first HPOD mode, shown in figure 15c, and disappear after $x/D=3$. The structures in the planar Reynolds stress of the HPOD mode are shorter and more consistent in size across the measured domain. The intensity of the structures is lower than those in the POD modes, but decays more slowly, with structure visible past $x/D=3$. The phase-averaged TKE of the POD modes, shown in figure 14d, is concentrated in periodically shed packets which initially increase in intensity until $x/D=2$, and then decay by $x/D=3$. The TKE of the HPOD mode, shown in figure 15d, reaches exhibits the same periodic structures as the POD modes, however, the intensity is increased. This is consistent with the TKE of the analytic signal of the turbulent fluctuations being significantly higher than that of the actual turbulent fluctuations.
Figure 16: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the fourth and fifth POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
Figure 17: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the second HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
The structure represented by the fourth and fifth POD modes and the second HPOD mode is a pulsating in the streamwise fluctuations, as can be seen from the deformation of the low velocity region behind the sphere, shown in figure 16a and 17a, respectively. The streamwise pulsation can also be seen in the phase-averaged out-of-plane vorticity, with the high-vorticity regions either side of the sphere extend further toward the centreline close to the sphere in figures 16b(ii) and 17b(ii). This extension toward the centreline occurs around $x/D=2$ in figures 16b(v) and 17b(v). While the structures in the planar Reynolds stress are highly variable in shape and size in the phase-averaged POD modes, shown in figure 16c, they are consistent in size and shape in the HPOD mode, shown in figure 17c. The size and shape of the phase-averaged TKE of the POD modes, shown in figure 16d, is more consistent than that of the planar Reynolds stresses. These structures are similar to, but less planar symmetric, than the structures in the phase-averaged TKE of the HPOD mode, which also exhibit a higher intensity with a more consistent pattern. The intensity of the TKE increases from directly behind the sphere until $x/D=2$, as seen in figure 17d(ii), before decreasing further downstream, as seen in figure 17d(iv).
Figure 18: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the sixth and eighth POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
Figure 19: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the third HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
The sixth and eighth POD modes represent a flapping motion in the wake, similar to that of the first and second POD modes but with a shorter wavelengths. Thus, the phase-averaged velocity, shown in figure 18a, and phase-averaged vorticity, shown in figure 18b, are similar to those of the first and second POD modes. The smaller structures are evident in the phase-averaged planar Reynolds stress, shown in figure 18c, particularly near the sphere. Further downstream of the sphere, the structures increase in streamwise extent, most notable in figure 18c(iv). The smaller structures in the phase-averaged TKE, shown in figure 18d, are more definitively separated across the centreline of the sphere than in the case of the first and second POD modes, where the larger structures extended to the centreline. The structures in the phase-averaged planar Reynolds stresses of the third HPOD mode, shown in figure 19c, are more consistent in size, shape and arrangement than those of the POD modes. There is also a region with low planar Reynolds stress around $x/D=2$ for all phase angles shown in figure 19, which reflects the smaller structures in the streamwise turbulent fluctuations in this regions, shown in figure 7c. The structures in the phase-averaged TKE are also more clearly defined for the HPOD modes, as shown in figure 19d.
Figure 20: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the ninth and tenth POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
Figure 21: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the fourth HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
The structure represented by the ninth and tenth POD modes, and the fourth HPOD mode, is similar to that represented by the fourth and fifth POD modes, and the second HPOD mode, namely periodic streamwise fluctuations in the wake. The phase-averaged velocity, shown in figures 20a and 21a for the POD modes and HPOD mode, respectively, and the phase-averaged vorticity, shown in figures 20b and 21b for the POD modes and HPOD mode, respectively, are qualitatively similar to those of the lower order modes. The structures in the planar Reynolds stress phase-averaged using the POD modes, shown in figure 20c, are small and round in the near wake of the sphere and larger and more ovular further downstream, with the latter being angles away from the centreline in accordance with the V-shaped structures in the POD modes. While the arrangement of these structures is variable, particularly close to the sphere, in the phase-average using the POD modes, the planar Reynolds stress of the HPOD mode, shown in figure 21c, clearly shows the round structures for $x/D<2$ and more V-shaped structures for $x/D>2$. Both the POD modes and HPOD modes show a noticeable reduction in the planar Reynolds stress around $x/D=2$, which delineates the rounder structures from the V-shaped ones. However, the HPOD modes display this more prominently than the POD modes, which also exhibit additional structures in the very near wake which are filtered out in the HPOD modes. The change between round structures near the sphere to more V-shaped structures is reflected in the phase-averaged TKE, shown in figure 20d and 21d, for the POD modes and HPOD mode, respectively. This also shows that the structures close to the sphere are more intense and decay in intensity toward $x/D=2$, with the V-shaped structures increasing in intensity around $x/D=3$ before decaying in intensity downstream. As in the case of the planar Reynolds stress, the structures in the TKE are more clearly defined in the HPOD mode due to the filtering of the Hilbert transform.
The structures elucidated by the POD mode pairs identified from the HPOD modes differ from those in the corresponding HPOD modes. The primary difference between these methods results from the filtering inherent in the Hilbert transform, as the non-propagating structures are given less importance in the HPOD, while propagating structures are amplified. The phase-averaging of the resulting modes thus results in more consistently shaped structures making it easier to identify and interpret the propagating structures. However, the filtering out of non-propagating structures may result in a poorer representation of the flow structure where the flow is three-dimensional. While applying the Hilbert transform directly to the POD modes is more computational efficient than performing the HPOD for pairing POD modes, it also forcibly propagates structures in the flow which may not propagate in reality. Thus, both methods of pairing POD modes introduce non-physical artefacts into the analysis of the flow and this should be considered when interpreting results.
5 Concluding Remarks
The velocity in the wake of a sphere at $Re_{D}=7780$ was measured using 2C-2D MCCDPIV, and POD and HPOD were applied to the turbulent fluctuations. The resulting POD and HPOD modes were compared to determine if the HPOD modes, which accentuate propagating structures, could be used to match POD modes which represented the same structure at different phase angles. The Hilbert transform of the POD modes was then used as a substitute for the HPOD modes in order to determine paired POD modes without performing the computationally-intensive HPOD. The POD mode pairs identified using these methods were then used to phase average the turbulent fluctuations to examine the propagating structures in the wake of the sphere.
The leading HPOD modes are shown to represent the same structures as the leading POD modes, with a greater emphasis on the propagating structures in the former. While each HPOD mode consists of a real and imaginary part representing a phase-shift of 90 in the mode, this comes at the cost of filtering from the Hilbert transform. As a result, it can be difficult to differentiate between what is part of the flow structures and what is an artefact of this filtering, particularly in the case of two-dimensional measurements of a three-dimensional flow. Performing the Hilbert transform on the POD modes to create an analytic signal of the POD modes is shown to identify the same pairs of POD modes identified using the HPOD modes. This method allows the same structures identified using the HPOD modes to be identified, without introducing the filtering resulting from the HPOD. The phase average of the turbulent fluctuations using the identified POD mode pairs represented flapping and pulsating motions in wake of the sphere.
As the velocity fields in the current work are not time-resolved, the Hilbert transform was performed in the streamwise direction, and thus the structures identified are those which propagate in that direction. As a result, the temporal information of the structures is not given by either decomposition. However, the paired POD modes gave insight into the dynamics of the structures in the turbulent fluctuations when the complexity of the flow made identifying related POD modes challenging.
Acknowledgements
Shaun Davey gratefully acknowledges the support of the Australian Commonwealth Government through a Research Training Program (RTP) Scholarship. This research was undertaken with the support of resources provided via a Monash HPC-NCI Merit Allocation from the National Computational Infrastructure (NCI Australia), an NCRIS-enabled capability supported by the Australian Government. The authors would like to acknowledge the support of the Australian Research Council (ARC) in funding this research through Linkage Infrastructure, Equipment and Facilities (LIEF) grants LE100100222 and LE180100166.
Declaration of Interests
The authors report no conflict of interest.

### v2

@@
 Proper orthogonal decomposition (POD) has been widely used to extract these structures from experimental data.
-While periodic features like vortex shedding can be identified using POD mode pairs when periodicity dominates the flow, detecting such structures in complex flows is more challenging.
-The Hilbert proper orthogonal decomposition (HPOD) addresses this by applying POD to the analytic signal of the turbulent fluctuations, yielding complex modes with a $90^{\circ}$ phase shift between the real and imaginary components.
-These modes capture propagating structures effectively but introduce filtering artefacts from the Hilbert transform that is used to derive the analytic signal.
+Periodic features such as vortex shedding can appear as POD mode pairs in strongly periodic flows, but detecting propagating structures in more complex flows is challenging.
+Hilbert proper orthogonal decomposition (HPOD) addresses this by applying POD to the analytic signal of the turbulent fluctuations, which yields complex modes with a $\pi/2$ phase shift between the real and imaginary components.
+These modes capture propagating structures effectively but introduce spectral leakage from the Hilbert transform used to derive the analytic signal.
 The current work investigates the relationship between the modes of the POD and those of the HPOD on the velocity fluctuations in the wake of a sphere.
 By comparing their outputs, POD mode pairs that correspond to the same propagating structures revealed by HPOD are identified.
-Furthermore, this study explored whether computing the analytic signal of the POD modes can replicate the HPOD modes, offering a more computationally efficient method for determining the pairs of POD modes that represent propagating structures.
-The results show that the pairs of POD modes identified by the HPOD can be more efficiently determined using the Hilbert transform directly on the POD modes.
-This method enhances the interpretive power of POD, enabling more detailed analysis of turbulent dynamics without introducing the filtering from the Hilbert transform.
+Furthermore, this study explores whether computing the analytic signal of the POD modes can replicate the HPOD modes, offering a more computationally efficient method for determining the pairs of POD modes that represent propagating structures.
+The results show that the pairs of POD modes identified by the HPOD can be determined more efficiently using the Hilbert transform directly on the POD modes.
+This method enhances the interpretive power of POD, enabling more detailed analysis of turbulent dynamics without the need to compute the analytic signal of the entire turbulent fluctuation data.
+Keywords: Wakes, Turbulent flows, Low-dimensional models
 1 Introduction
-While turbulence often appears random, many turbulent flows contain underlying structures, such as eddies and vortices, the scale and nature of which vary depending on the conditions of the flow.
-Identifying these structures allows for a more detailed description of the flow, as well as the isolation of the dominant structures in the turbulence based on their contribution to the quantity or phenomenon of interest, such as turbulent kinetic energy (TKE) or mixing.
+While turbulence often appears random, many turbulent flows contain underlying structures, such as eddies and vortices.
+The scale and nature of these structures vary depending on the flow conditions.
+Identifying these structures allows for a more detailed description of the flow, as well as the isolation of the dominant structures within the turbulence based on their contribution to the quantity or phenomenon of interest, such as turbulent kinetic energy (TKE) or mixing.
 Ranking these structures based on their significance to a given application or study allows the construction of parsimonious reduced-order models, which can be used to reduce the computational intensity of analyses or to reduce noise from experimental data.
-The complex flow phenomena present in the flow over bluff bodies, even at low Reynolds numbers, develop as the Reynolds number increases, and are a subject of ongoing study.
-In this respect, the flow over spheres serves as a canonical model for flows over axisymmetric bluff bodies, and is of particular interest for developing a broad understanding of these flow phenomena.
-The phenomena observed in the wake of a sphere include transition from axisymmetry to planar symmetry to asymmetry, the formation and shedding of vortices, the development of shear layer instabilities, and flow separation.
-As the flow over a sphere varies significantly with Reynolds number, experimental and numerical investigations have covered a wide range of Reynolds number regimes.
-In the context of experimental studies of the flow over spheres, one of the main challenges is fixing the sphere in a wind or water tunnel test section with minimal disturbance of the flow, since more often than not, the mounting/support structure of the sphere can have a significant influence on the flow.
-The supports chosen for an experiment should simultaneously provide rigid and stable positioning of the sphere, and minimise the disturbance to the flow, while also facilitating the placement of probes without obstructing optical measurement access.
-Some of the supports used in these studies include rigid downstream supports [1], honeycomb structures with rigid supports upstream of the sphere [2, 3], and wires placed azimuthally [4, 5].
-The drag on the spheres has been measured using hot-wire probes [1, 6], with flow visualisations performed using aluminium dust [7], smoke [8, 5], and dyes [9, 4, 10] to gain a qualitative understanding of the recirculation region, flow separation, and vortex shedding in the wake.
-Various forms of particle image velocimetry (PIV), including planar [5], stereographic [3] and tomographic [11], have been used to quantitatively measure the flow around spheres.
-While mounting the sphere is not a concern in numerical studies, simulations are subject to their own set of challenges.
-The geometry of the problem and the range of scales involved require careful consideration of the mesh and solver to resolve the flow at high Reynolds numbers.
-Because of this complexity, large eddy simulation (LES[12, 13, 14, 15] and detached eddy simulation (DES[14, 16] methods have been favoured over the more computationally expensive direct numerical simulation (DNS[12, 15, 17, 18] to resolve the main features of the flow at high Reynolds numbers.
-In conjunction with the advancement of computational capabilities, these models have extended the feasibility of numerical investigations of the flow over spheres to large Reynolds numbers.
-Experimental and numerical studies have examined the development of the flow over spheres over a wide range Reynolds number regimes.
-At low Reynolds number, $Re_{D}<20$, the flow in the wake is laminar and axisymmetric, and there is no recirculation [7, 13].
+The complex flow phenomena present in the flow over bluff bodies, even at low Reynolds numbers, develop with increasing Reynolds number, and remain a subject of ongoing study.
+In this respect, the flow over spheres serves as a canonical model for flows over axisymmetric bluff bodies, and is of particular interest for advancing the understanding of these flow phenomena.
+The phenomena observed in the wake of a sphere include transition from axisymmetry to planar symmetry, from planar symmetry to asymmetry, the formation and shedding of vortices, the development of shear layer instabilities, and flow separation.
+As the flow over a sphere varies significantly with Reynolds number, experimental and numerical investigations have explored a wide range of Reynolds number regimes.
+In the context of experimental studies of the flow over spheres, one of the main challenges is fixing the sphere in a wind or water tunnel test section with minimal disturbance of the flow, since the mounting or support structure of the sphere frequently has a significant influence on the flow.
+The supports chosen for an experiment should simultaneously provide rigid and stable positioning of the sphere, and minimise the disturbance to the flow, while facilitating probe placement and maintaining unobstructed optical access for measurements.
+Some of the supports used in these studies include rigid downstream supports [1], honeycomb structures with rigid supports upstream of the sphere [15, 23], and wires arranged azimuthally [34, 17].
+The drag on the spheres has been measured using hot-wire probes [1, 44], with flow visualisations performed using aluminium dust [38], smoke [39, 17], and dyes [18, 34, 2] to provide a qualitative understanding of the recirculation region, flow separation, and vortex shedding in the wake.
+Various forms of particle image velocimetry (PIV), including planar [17], stereoscopic [23], and tomographic [12], have been used to quantitatively measure the flow around spheres.
+While mounting the sphere is not a concern in numerical studies, simulations are subject to distinct challenges.
+The problem geometry and the wide range of scales involved require careful consideration of the mesh and solver to accurately resolve the flow at high Reynolds numbers.
+Because of this complexity, large eddy simulation (LES[41, 42, 9, 33] and detached eddy simulation (DES[9, 10] methods have been preferred over the more computationally expensive direct numerical simulation (DNS[41, 33, 31, 32] for resolving the primary features of the flow at high Reynolds numbers.
+In conjunction with the advancement of computational capabilities, these models have expanded the scope of numerical investigations of the flow over spheres to large Reynolds numbers.
+Experimental and numerical studies have examined the development of the flow over spheres over a wide range of Reynolds number regimes.
+At low Reynolds number, $Re_{D}<20$, the flow in the wake is laminar and axisymmetric, and recirculation is absent [38, 42].
 A small vortex ring forms near the rear stagnation point around $Re_{D}\approx 25$ [38], and elongates in the streamwise direction with increasing Reynolds number.
 This vortex ring begins to oscillate around $Re_{D}\approx 130$ with a long period, which becomes more prominent at higher Reynolds numbers.
-The recirculation region formed behind the sphere, which exhibits reverse flow in the mean, extends further downstream of the sphere as Reynolds number increases [9].
-The axisymmetry in the wake at low Reynolds number is lost due to the formation of a double-threaded wake around $Re_{D}\approx 210$ [9, 13], which consists of two streamwise vortices of opposing sign that can be seen as dye threads in the wake of a free-falling sphere [19].
+The recirculation region formed behind the sphere, which exhibits reverse flow in the mean flow, extends further downstream of the sphere as Reynolds number increases [18].
+The axisymmetry in the wake at low Reynolds number is lost due to the formation of a double-threaded wake around $Re_{D}\approx 210$ [18, 42], which consists of two streamwise vortices of opposite sign that can be seen as dye threads in the wake of a free-falling sphere [22].
 Single-frequency vortex shedding begins around $Re_{D}\approx 300$ [38, 34], and hairpin-shaped vortices form at $Re_{D}\approx 400$ [2].
-The generation of an unsteady random side force, which has been observed in both experimental [8] and numerical [20] studies results in the loss of planar symmetry in the wake in the range of $400<Re_{D}<500$ [7].
-From $Re_{D}=800$, the small-scale Kelvin-Helmholtz instability in the shear layer at the edge of the recirculation region manifests as axisymmetric vortex shedding [16, 21].
-This is the higher frequency of two instability modes identified in the wake of the sphere by [22], with the lower frequency mode arising from large-scale vortex shedding.
+The generation of an unsteady random side force, which has been observed in both experimental [39] and numerical [49] studies, leads to the loss of planar symmetry in the wake within the range of $400<Re_{D}<500$ [38].
+For $Re_{D}\gtrapprox 800$, the small-scale Kelvin-Helmholtz instability in the shear layer at the edge of the recirculation region appears as axisymmetric vortex shedding [10, 8].
+This is the higher-frequency of two instability modes identified in the wake of the sphere by Kim & Durbin [19], with the lower-frequency mode arising from large-scale vortex shedding.
 Small-scale vortical structures are shed from the main vortical structure around $Re_{D}\approx 1000$ [49], and the recirculation region behind the sphere begins to shrink with increasing Reynolds number.
-Periodic fluctuations are present in the wake for much higher Reynolds numbers, $Re_{D}=O(10^{6})$ [10].
-Since its first application to turbulent flows by [23], proper orthogonal decomposition (POD) has been widely used for identifying and analysing structures in turbulent flows [24, 25].
-As POD does not require temporal resolution of the flow and can thus be applied to non-time-resolved data, it is a robust and simple method for identifying the structures within a turbulent flow.
-In the case of the flow over a cylinder, the turbulent fluctuations due to the vortices shed from the sphere can be identified in the leading POD modes [26, 27].
+Periodic fluctuations are present in the wake for significantly higher Reynolds numbers, $Re_{D}=O(10^{6})$ [2].
+Since its first application to turbulent flows by Lumley [21], proper orthogonal decomposition (POD) has been widely used for identifying and analysing structures in turbulence [35, 4].
+As POD does not require temporal resolution of the flow and can thus be applied to non-time-resolved data, it is a robust yet straightforward method for identifying the structures within a turbulent flow.
+In the case of the flow over a cylinder, the turbulent fluctuations due to the vortices shed from the cylinder can be identified in the leading POD modes [24, 26].
 The energy of these modes is significantly larger than the higher-order modes, and their phase angle is clearly related to that of the pressure fluctuations resulting from the shed vortices [25].
-Similarly, periodic structures in screeching, impinging and free jets [29, 30, 31], are represented by the leading POD modes.
-Identifying pairs of POD modes which represent periodic structures allows for the a posteriori phase-averaging of experimental data which is collected asynchronously or at low frequency relative to these structures.
-However, the complexity of the flow over a sphere in this regime makes identification of periodic structures in the turbulent fluctuations challenging, as POD modes representing different phases of a structure are not linked by the decomposition.
-The analytic signal of the turbulent fluctuations, which is defined using the Hilbert transform, can be decomposed into complex modes, which include phase information, using POD [32].
-While the Hilbert transform is performed in time for time-resolved data, it can also performed in an analogous spatial direction when the data is not time-resolved [33, 34, 35].
-This results in modes emphasising structures which are periodic in the chosen direction, with the streamwise direction being the most sensible choice to identify structures which propagate downstream from the sphere.
-Although a useful method for identifying propagating structures, Hilbert POD (HPOD) introduces artefacts to the modes as a result of the Hilbert transform.
-The current work applies both POD and HPOD, with the Hilbert transform performed in the streamwise direction, to experimentally-measured velocity fluctuations in the wake of a sphere at $Re_{D}=7780$.
-The modes resulting from each of these decompositions are compared, and the correlation between the POD modes and the HPOD modes is used to determine pairs of POD modes representing propagating structures.
-The data is then phase-averaged using the elucidated POD mode pairs to observe the periodic behaviour of the propagating structures.
-By pairing POD modes using the HPOD modes, the dynamics of the turbulent structures within the wake can be determined while avoiding the filtering that is introduced by the Hilbert transform.
-Additionally, the same mode pairing can be achieved using the streamwise Hilbert transform of the individual POD modes.
-This method significantly reduces the computational cost of identifying paired POD modes, as it does not require POD to be performed on the complex-valued analytic signal.
-It is shown that the leading modes of POD and HPOD are similar in shape, and have comparable phase angles throughout the experimental data.
-However, the HPOD modes accentuate the propagating parts of the structures and introduce artefacts which are not present in the POD modes.
+Similarly, periodic structures in screeching, impinging, and free jets [13, 45, 46] are represented by the leading POD modes.
+Identifying pairs of POD modes which represent periodic structures enables the a posteriori phase-averaging of experimental data collected asynchronously or at a low sampling frequency relative to these structures.
+However, the complexity of the flow over a sphere in this regime makes identification of periodic structures in the turbulent fluctuations challenging, as POD modes representing different phases of a structure are not inherently linked by the decomposition.
+The analytic signal of the turbulent fluctuations, which is defined using the Hilbert transform, can be decomposed into complex modes that include phase information using POD [16].
+While the Hilbert transform is performed in time for time-resolved data, it can also be performed in an analogous spatial direction when the data is not time-resolved [27, 28, 20].
+This results in modes emphasising structures which are periodic in the chosen direction, with the streamwise direction being the most appropriate choice to identify structures which propagate downstream from the sphere.
+Although a useful method for identifying propagating structures, Hilbert POD (HPOD) introduces spectral leakage to the modes due to the Hilbert transform being derived from the Fourier transform, which inherently assumes periodicity.
+Raiola & Kriegseis [29] demonstrated the ability of spatial HPOD to extract spatiotemporally coherent wavepackets from flows with significant advecting structures.
+Their application of HPOD to DNS of a laminar vortex street in the wake of a cylinder at $Re_{D}=100$ showed that the spatial HPOD provides a more compact decomposition than POD.
+In this case, the first two POD modes representing the vortex street were captured by a single HPOD mode.
+In this example, the leading mode of the spatial HPOD was able to capture the first two POD modes even when the data were shuffled in time and the dataset reduced in size, whereas the modes of the conventional time-based HPOD converged towards the POD modes, and no longer captured paired POD modes in a single HPOD mode.
+They further applied HPOD to LES of a turbulent jet at $Re_{D}=10^{6}$ [43], showing that the spatial HPOD could still identify advecting wavepackets in a flow with a broader range of scales.
+Finally, HPOD was applied to PIV data from a turbulent subsonic jet [30] at $Re=33,000$, which lacked the temporal resolution afforded by the LES data.
+Despite this limitation, the spatial HPOD was able to extract spatial modes similar to those identified in the LES data.
+The current work applies both POD and HPOD, with the Hilbert transform performed in the streamwise direction following the methodology of [29], to experimentally measured velocity fluctuations in the wake of a sphere at $Re_{D}=7780$.
+Modes resulting from each decomposition are compared, and correlations between the POD and HPOD modes are used to identify POD mode pairs representing propagating structures.
+A novel approach is introduced in which the Hilbert transform is applied directly to the POD modes, allowing mode pairs associated with periodic or propagating structures to be identified without performing POD on the complex analytic signal of the turbulent fluctuations.
+This method significantly reduces the computational cost of identifying propagating structures and preserves the energy ranking of the modes relative to the original data.
+The identified POD mode pairs are then used to phase-average the velocity fluctuations in the wake of the sphere to examine the propagating structures, while phase-averaging based on HPOD modes is also performed to demonstrate the influence of HPOD on the identified structures due to its emphasis on propagating structures and inherent assumption of periodicity.
+Section 2 outlines the theoretical aspects of POD and HPOD and the rationale for applying the Hilbert transform directly to the POD modes.
+Section 3 provides details of the experimental facility and measurement methodology.
+Section 4 presents the analysis and results of the POD and HPOD, the pairing of POD modes using HPOD modes, the application of the Hilbert transform directly to the POD modes, and the resulting phase averages of the turbulent fluctuations.
+Section 5 summarises the work, highlights the key findings, and provides closing remarks.
 2 Theoretical Considerations
 2.1 Proper Orthogonal Decomposition
-POD decomposes statistically independent realisations of the random field $\boldsymbol{f}(\boldsymbol{x},t)$ into orthonormal spatial modes $\psi_{k}(\boldsymbol{x})$ with temporal coefficients $a_{k}(t)$, which are ranked by their energy $\lambda_{k}$, from largest to smallest, such that
+POD decomposes statistically independent snapshots of the random field $\boldsymbol{f}(\boldsymbol{x},t)$ into orthonormal spatial modes $\psi_{k}(\boldsymbol{x})$ with temporal coefficients $a_{k}(t)$, which are ranked by their associated energy $\lambda_{k}$ in descending order, such that
 $\boldsymbol{f}(\boldsymbol{x},t)=\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x}),$ (1)
-and is analogous to singular value decomposition (SVD) and principle component analysis (PCA).
+and is mathematically analogous to singular value decomposition (SVD) and principal component analysis (PCA).
 In the application of POD to turbulence, $\boldsymbol{f}(\boldsymbol{x},t)$ represents the turbulent fluctuations, and $\lambda_{k}$ represents the contribution of the $k^{th}$ mode to the total TKE of $\boldsymbol{f}(\boldsymbol{x},t)$.
@@
 In order to perform POD, the two-point correlation of $\boldsymbol{f}(\boldsymbol{x},t)$ in space
-[…]
-is calculated, where $T$ is the temporal domain. The spatial modes and the corresponding energy are given by the eigendecomposition of $C(\boldsymbol{x},\boldsymbol{x}^{\prime})$
-[…]
-where $S$ is the spatial domain. The temporal coefficients are the projection of $\boldsymbol{f}(\boldsymbol{x},t)$ onto $\psi(\boldsymbol{x})$
-[…]
-Alternatively, POD can be performed using the method of snapshots, which first determines the temporal coefficients, and then the spatial modes.
-This method is more computationally efficient when the number of spatial coordinates included in $\boldsymbol{f}(\boldsymbol{x},t)$ is larger than the number of realisations of the field.
-The two-point correlation in time
-[…]
-is used instead of that in space, given in equation (4). The eigendecomposition of $C(t,t^{\prime})$
-[…]
-yields the eigenvalues and temporal coefficients. The spatial modes are then determined by the projection of $\boldsymbol{f}(\boldsymbol{x},t)$ onto $a(t)$
-[…]
-The temporal coefficients of the $i^{th}$ and $j^{th}$ POD modes can be expressed in terms of their phase angle $\phi_{i,j}$ as
-$a_{i}(t)=\sqrt{2\lambda_{i}}\cos{\big{(}\phi_{i,j}(t)\big{)}},$ (10)
+$C(\boldsymbol{x},\boldsymbol{x}^{\prime})=\int_{T}\boldsymbol{f}(\boldsymbol{x},t)\boldsymbol{f}(\boldsymbol{x}^{\prime},t)\,dt,$ (4)
+is calculated, where $T$ denotes the temporal domain. The spatial modes and the corresponding energy are given by the eigendecomposition of $C(\boldsymbol{x},\boldsymbol{x}^{\prime})$
+$\int_{S}C(\boldsymbol{x},\boldsymbol{x}^{\prime})\psi(\boldsymbol{x}^{\prime})\,d\boldsymbol{x}^{\prime}=\lambda\psi(\boldsymbol{x}),$ (5)
+where $S$ denotes the spatial domain. The temporal coefficients are computed by projecting $\boldsymbol{f}(\boldsymbol{x},t)$ onto $\psi(\boldsymbol{x})$
+$a(t)=\int_{S}\boldsymbol{f}(\boldsymbol{x},t)\psi(\boldsymbol{x})\,d\boldsymbol{x}.$ (6)
+Alternatively, POD can be implemented using the method of snapshots, which first computes the temporal coefficients, and subsequently determines the spatial modes.
+This approach is more computationally efficient when the number of spatial coordinates in $\boldsymbol{f}(\boldsymbol{x},t)$ exceeds the number of realisations of the field.
+The two-instant temporal correlation
+$C(t,t^{\prime})=\int_{S}\boldsymbol{f}(\boldsymbol{x},t)\boldsymbol{f}(\boldsymbol{x},t^{\prime})\,d\boldsymbol{x},$ (7)
+is used instead of the spatial correlation defined in equation (4). The eigendecomposition of $C(t,t^{\prime})$
+$\int_{T}C(t,t^{\prime})a(t^{\prime})\,dt^{\prime}=\lambda a(t),$ (8)
+yields the eigenvalues and corresponding temporal coefficients. The spatial modes are subsequently obtained by projecting $\boldsymbol{f}(\boldsymbol{x},t)$ onto $a(t)$
+$\psi(\boldsymbol{x})=\int_{T}\boldsymbol{f}(\boldsymbol{x},t)a(t)\,dt.$ (9)
+When the $i^{th}$ and $j^{th}$ POD modes represent the same periodic structure separated in phase by $\pm\pi/2$, their temporal coefficients can be expressed as functions of the phase angle $\phi_{i,j}$ of the structure as
+$a_{i}(t)=\sqrt{2\lambda_{i}}\cos(\phi_{i,j}(t)),$ (10)
 and
-$a_{j}(t)=\sqrt{2\lambda_{j}}\sin{\big{(}\phi_{i,j}(t)\big{)}},$ (11)
-respectively. Thus, the phase angle $\phi_{i,j}(t)$ is given by
+$a_{j}(t)=\sqrt{2\lambda_{j}}\sin(\phi_{i,j}(t)),$ (11)
+respectively. Consequently, the phase angle $\phi_{i,j}(t)$ is given by
 $\phi_{i,j}(t)=\tan^{-1}\left(\frac{a_{j}(t)/\sqrt{2\lambda_{j}}}{a_{i}(t)/\sqrt{2\lambda_{i}}}\right).$ (12)
-The phase average of the data using the $i^{th}$ and $j^{th}$ POD modes at the phase angle $\phi$ is given by
+The phase-averaged representation of the data using the $i^{th}$ and $j^{th}$ POD modes at the phase angle $\phi$ is defined as
 $\widetilde{\boldsymbol{f}}(\boldsymbol{x},\phi)=\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})\,|\,\phi_{i,j}(t)=\phi$ (13)
 where $\phi_{i,j}(t)$ is defined by equation (12).
-While any two POD modes can be used for phase-averaging $\boldsymbol{f}(\boldsymbol{x},t)$, not all selections of modes result in physically meaningful phase averages.
-In the case of strongly dominant periodic fluctuations, the temporal coefficients of the corresponding modes form a circle [26].
-Large periodic structures can also be identified by comparing the phase angle of the modes to an independent measurement of the phase of the flow, such as pressure signals [28].
-However, the less prominent the periodic structures, the more difficult it is to relate the POD modes which represent them.
+Although any two POD modes can be employed for phase-averaging $\boldsymbol{f}(\boldsymbol{x},t)$, not all mode selections yield physically meaningful phase averages.
+For strongly dominant periodic fluctuations, the temporal coefficients of the corresponding modes exhibit a circular distribution in the coefficient space [24].
+Large-scale periodic structures can also be identified by comparing the phase angle of the modes to an independent phase measurement of the flow, such as pressure signals [25].
+However, as periodic structures become less prominent, establishing a clear relationship between the corresponding POD modes becomes increasingly challenging.
 2.2 Hilbert Proper Orthogonal Decomposition
-HPOD embeds the phase information in the modes by replacing $\boldsymbol{f}(\boldsymbol{x},t)$ with its analytic signal [32]
+HPOD incorporates phase information into the modes by replacing $\boldsymbol{f}(\boldsymbol{x},t)$ with its analytic signal [16]
 $\boldsymbol{f}^{a}(\boldsymbol{x},t)=\boldsymbol{f}(\boldsymbol{x},t)+i\mathcal{H}[\boldsymbol{f}(\boldsymbol{x},t)],$ (14)
-where $\mathcal{H}$ denotes the Hilbert transform of $\boldsymbol{f}(\boldsymbol{x},t)$ in time
-[…]
-The Hilbert transform is equivalent to a $\mp 90^{\circ}$ phase shift [36] of the fundamental frequencies of $\boldsymbol{f}(\boldsymbol{x},t)$ in Fourier space
+where $\mathcal{H}$ denotes the Hilbert transform of $\boldsymbol{f}(\boldsymbol{x},t)$ with respect to time
+$\mathcal{H}\big[\boldsymbol{f}(\boldsymbol{x},t)\big]=\frac{1}{\pi}\int_{-\infty}^{\infty}\frac{\boldsymbol{f}(\boldsymbol{x},\tau)}{t-\tau}\,d\tau.$ (15)
+The Hilbert transform is equivalent to a $\mp\pi/2$ phase shift of the fundamental frequency components of $\boldsymbol{f}(\boldsymbol{x},t)$ in Fourier space [40]
 $\mathcal{H}\big[\boldsymbol{f}(\boldsymbol{x},t)\big]=\mathcal{F}^{-1}\Big[-i\operatorname{sgn}(f)\mathcal{F}\big[\boldsymbol{f}(\boldsymbol{x},t)\big]\Big],$ (16)
@@
 is the sign of the instantaneous frequency $f$ of each term in the Fourier series given by $\mathcal{F}\big[\boldsymbol{f}(\boldsymbol{x},t)\big]$.
-For time-resolved data, the Hilbert transform can be performed in time to determine the analytic signal $\boldsymbol{f}^{a}(\boldsymbol{x},t)$.
-HPOD is performed on time-resolved data by replacing $\boldsymbol{f}(\boldsymbol{x},t)$ with its analytic signal in equations (4) to (6).
-However, where the coherent structures propagate downstream, the Hilbert transform can be performed in the streamwise direction, as propagation in this direction serves as an analogy for a change in time, to apply this method to data that is not time-resolved [33, 34, 35]
-[…]
-In this case, the analytic signal of $\boldsymbol{f}(\boldsymbol{x},t)$ is defined using $\mathcal{H}_{x}$
+For time-resolved data, the Hilbert transform is applied with respect to time to compute the analytic signal $\boldsymbol{f}^{a}(\boldsymbol{x},t)$.
+HPOD is implemented on time-resolved data by substituting $\boldsymbol{f}(\boldsymbol{x},t)$ in equations (4) to (6) with its analytic signal.
+For flows where coherent structures propagate downstream, the Hilbert transform can alternatively be applied in the streamwise direction
+$\mathcal{H}_{x}\big[\boldsymbol{f}(\boldsymbol{x},t)\big]=\frac{1}{\pi}\int_{-\infty}^{\infty}\frac{\boldsymbol{f}\big((\xi,\ldots),t\big)}{x-\xi}\,d\xi.$ (18)
+This approach uses streamwise propagation as an analogue for temporal evolution, allowing the method to be applied to data that are not time-resolved [27, 28, 20].
+In this case, the analytic signal of $\boldsymbol{f}(\boldsymbol{x},t)$ is defined using the streamwise Hilbert transform $\mathcal{H}_{x}$
 $\boldsymbol{f}^{a}_{x}(\boldsymbol{x},t)=\boldsymbol{f}(\boldsymbol{x},t)+i\mathcal{H}_{x}\big[\boldsymbol{f}(\boldsymbol{x},t)\big],$ (19)
-and substituted for $\boldsymbol{f}(\boldsymbol{x},t)$ in equations (7) to (9).
-The resulting eigenvalues $\lambda^{a}_{k}$ are the contribution of the $k^{th}$ mode to the total TKE of the analytic signal, and are real-valued.
-The spatial modes $\psi^{a}_{k}(\boldsymbol{x})$ are complex, with the imaginary part of $\psi^{a}_{k}(\boldsymbol{x})$ representing a phase shift of its real part by $\mp 90^{\circ}$.
-The instantaneous amplitude and phase angle of the $k^{th}$ mode are defined by the real $\Re$ and imaginary $\Im$ parts of $a^{a}_{k}(t)$ as
+and is substituted for $\boldsymbol{f}(\boldsymbol{x},t)$ in equations (7) to (9).
+The resulting eigenvalues $\lambda^{a}_{k}$ represent the contribution of the $k^{th}$ mode to the total TKE of the analytic signal and are real-valued.
+The spatial modes $\psi^{a}_{k}(\boldsymbol{x})$ are complex, with the imaginary component of $\psi^{a}_{k}(\boldsymbol{x})$ corresponding to a $\mp\pi/2$ phase shift of the real component.
+The instantaneous amplitude and phase angle of the $k^{th}$ mode are defined by the real $\Re$ and imaginary $\Im$ components of $a^{a}_{k}(t)$ as
 $|a^{a}_{k}(t)|=\sqrt{\Re{\big(a^{a}_{k}(t)\big)}^{2}+\Im{\big(a^{a}_{k}(t)\big)}^{2}},$ (20)
@@
 respectively.
-The instantaneous amplitude denotes the strength of the mode at each instant in time, and the instantaneous phase angle denotes the mode shape, which oscillates between the real and imaginary parts of $\psi_{k}^{a}(\boldsymbol{x})$ at $\phi_{k}^{a}(t)$ equal to $0^{\circ}$ and $\mp 90^{\circ}$, respectively.
-The analytic signal can be reconstructed using $\hat{K}$ modes by
+The instantaneous amplitude quantifies the strength of the mode at each instant, while the instantaneous phase angle characterises the mode shape, which oscillates between the real and imaginary components of $\psi^{a}_{k}(\boldsymbol{x})$ at $\phi^{a}_{k}(t)=0^{\circ}$ and $\phi^{a}_{k}(t)=\mp\pi/2$, respectively.
+The analytic signal can be reconstructed using $\hat{K}$ modes as
 $\hat{\boldsymbol{f}^{a}}(\boldsymbol{x},t)=\sum_{k=1}^{\hat{K}}a^{a}_{k}(t)\psi^{a}_{k}(\boldsymbol{x}),$ (22)
-the real part of which is the reduced-order model of $\boldsymbol{f}(\boldsymbol{x},t)$ including the modes which contribute the most TKE to the analytic signal.
-This reconstruction favours the propagating components of $\boldsymbol{f}(\boldsymbol{x},t)$, but introduces artefacts due to the Hilbert transform.
+where the real part is the reduced-order model of $\boldsymbol{f}(\boldsymbol{x},t)$ including the modes which contribute most to the TKE of the analytic signal.
+This reconstruction favours the propagating components of $\boldsymbol{f}(\boldsymbol{x},t)$, but introduces spectral leakage into the reconstruction due to the assumption of periodicity inherent to Hilbert transform.
 2.3 Hilbert Transform of POD Modes
-Although the POD modes do not contain temporal information, the streamwise Hilbert transform can be performed directly on each of the modes.
-Thus, the analytic signal of the POD mode $\psi_{k}(\boldsymbol{x})$ can be defined as
+Although POD modes do not inherently contain temporal information, the streamwise Hilbert transform can be applied directly to each mode.
+Thus, the analytic signal of the POD mode $\psi_{k}(\boldsymbol{x})$ is defined as
 ${(\psi_{k})}^{a}(\boldsymbol{x})=\psi_{k}(\boldsymbol{x})+i\mathcal{H}_{x}\big[\psi_{k}(\boldsymbol{x})\big].$ (23)
@@
 which can be expanded to
-$\sum_{k=1}^{K}a_{k}(t)\Big{(}(\psi_{k})^{a}(\boldsymbol{x})\Big{)}=\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})+a_{k}(t)i\mathcal{H}_{x}\big{[}\psi_{k}(\boldsymbol{x})\big{]}.$ (25)
-As the Hilbert transform is a linear operator, this can be rewritten as
-[…]
+$\sum_{k=1}^{K}a_{k}(t)\Big({(\psi_{k})}^{a}(\boldsymbol{x})\Big)=\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})+\sum_{k=1}^{K}a_{k}(t)i\mathcal{H}_{x}\big[\psi_{k}(\boldsymbol{x})\big].$ (25)
+Since the Hilbert transform is a linear operator, the expression can be rewritten as
+$\sum_{k=1}^{K}a_{k}(t)\Big({(\psi_{k})}^{a}(\boldsymbol{x})\Big)$ $=\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})+i\mathcal{H}_{x}\Big[\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})\Big]$ (26)
+$=\boldsymbol{f}(\boldsymbol{x},t)+i\mathcal{H}_{x}[\boldsymbol{f}(\boldsymbol{x},t)]$
+$=\boldsymbol{f}^{a}_{x}(\boldsymbol{x},t),$
 which is the analytic signal of $\boldsymbol{f}(\boldsymbol{x},t)$.
-Therefore, applying the transform to each of the POD modes of $\boldsymbol{f}(\boldsymbol{x},t)$ individually and using the resulting analytic signals in place of the original modes allows for the reconstruction of $\boldsymbol{f}^{a}(\boldsymbol{x},t)$.
-However, the analytic signal of the POD modes are not equal to the HPOD modes due to the filtering of the Hilbert transform being introduced after the decomposition of $\boldsymbol{f}(\boldsymbol{x},t)$.
-While this method is not equivalent to HPOD, where the propagation of structures is embedded in the modes, it is a computationally efficient alternative for matching pairs of POD modes which represent the same propagating structure.
+Therefore, applying the Hilbert transform to each POD mode of $\boldsymbol{f}(\boldsymbol{x},t)$ individually and substituting the resulting analytic signals for the original modes enables the reconstruction of $\boldsymbol{f}^{a}(\boldsymbol{x},t)$.
+However, the analytic signals of the POD modes are not equal to the HPOD modes, as the decomposition of the analytic signal of $\boldsymbol{f}(\boldsymbol{x},t)$ ranks HPOD modes by their energy contribution to $\boldsymbol{f}^{a}(\boldsymbol{x},t)$ rather than to $\boldsymbol{f}(\boldsymbol{x},t)$.
+Consequently, leading POD modes may contain non-propagating features that are relegated to higher-order HPOD modes.
+While this approach is not equivalent to HPOD, the tendency for propagating structures to be decomposed into paired POD modes—or real and imaginary parts of HPOD modes [29]—suggests that the Hilbert transform of a POD mode representing a propagating structure should closely match the POD mode corresponding to the $\pi/2$ phase-shift of the same structure.
 3 Methodology
@@
 Experiments were performed in a vertical water tunnel, the test section of which is 1.5 m long with a cross-section of 0.25 m $\times$ 0.25 m, shown in figure 1[7, 6, 11].
-The settling chamber is connected to the top of the test section by a 16:1 contraction, which includes a honeycomb and screens to reduce the turbulence level and scale in the tunnel test section.
+The settling chamber is connected to the top of the test section by a 16:1 contraction, which incorporates a honeycomb and screens to reduce the turbulence intensity and scale in the tunnel test section.
 Water is pumped from the plenum chamber below the test section up to the settling chamber such that the flow through the test section is downwards.
@@
 Seeding and filtration of the tunnel is performed via an auxiliary circuit, which is isolated during experiments.
-A sting of diameter 9.3 mm is mounted to a crossbeam of thin aerofoil cross-sections above the test section, and tapers in diameter over the last 100 mm down to 4.65 mm in diameter and ends in a threaded rod which is used to mount a sphere in the test section.
+A sting of diameter 9.3 mm is mounted to a crossbeam of thin aerofoil cross-sections above the test section.
+The sting tapers over the last 100 mm to a diameter of 4.65 mm and terminates in a threaded rod used to mount the sphere.
 A close-up view of the sphere mounting structure is shown in figure 1b.
-A sphere with a diameter of $D=40$ mm was manufactured using a Phrozen Sonic Mighty 8k 3-D printer, which has a 35 µm layer height and a planar resolution of 22 µm $\times$ 22 µm, as shown in figure 1c.
-The sphere has a shell thickness of 2.5 mm and an external boss at the top for mounting to a threaded rod on the end of the sting.
+A sphere with a diameter of $D=40$ mm was fabricated using a Phrozen Sonic Mighty 8k 3D printer, featuring a layer height of 35 µm and a planar resolution of 22 µm $\times$ 22 µm, as shown in figure 1c.
+The sphere has a shell thickness of 2.5 mm and an external boss at the top for mounting to the threaded rod on the end of the sting.
 The boss begins parallel to the sting and smoothly transitions to the sphere surface.
 The sphere is supported internally by a conical structure, which features slits to prevent uncured resin from being trapped in the cavities during printing.
-Figure 1: (a) Vertical water tunnel facility, and (b) close-up of the mounting structure and sphere, and (c) 3-D printed sphere design [39].
+Figure 1: (a) Vertical water tunnel facility, (b) close-up of the mounting structure and sphere, and (c) 3-D printed sphere design [11].
 3.2 Experimental Method
-Measurements of the flow in the wake of the sphere were taken with a freestream velocity of $U_{\infty}=200$ m/s, resulting in a Reynolds number of $Re_{D}=7780$.
+Flow measurements in the wake of the sphere were conducted at a freestream velocity of $U_{\infty}=200$ mm/s, corresponding to a Reynolds number of $Re_{D}=7780$.
 Under these conditions, the background turbulence is 2%, and the mean convective time of the flow
 $t_{c}=\frac{D}{U_{\infty}},$ (27)
-is 0.2 seconds [39].
-The flow was seeded with 11 µm glass spheres, and the single-exposed particle image pairs were recorded using a PCO Panda 26 DS camera with a 5120 $\times$ 5120 px array of 2.5 µm $\times$ 2.5 µm pixels.
-A Zeiss lens with a fixed 50 mm focal length was used to capture a 190 mm $\times$ 190 mm field of view, resulting in a spatial resolution is 37 µm/px.
+was 0.2 seconds [11].
+The flow was seeded with 11 µm glass spheres, and single-exposed particle image pairs were recorded using a PCO Panda 26 DS camera with a 5120 $\times$ 5120 px array of 2.5 µm $\times$ 2.5 µm pixels.
+A Zeiss lens with a fixed 50 mm focal length was used to capture a 190 mm $\times$ 190 mm field of view, resulting in a spatial resolution of 37 µm/px.
 The field of view began just behind the trailing edge of the sphere, to avoid reflections off the surface, and extended to 5.2 diameters downstream of the sphere.
 The field of view was aligned with the centreline of the sphere, and extended to 2.3 diameters either side of the centreline in the transverse direction.
-The f-stop was set to 2.8, resulting in a 5.0 mm depth of field.
-The diffraction limited minimum image diameter and expected particle image diameter for an 11 µm particle at these setting are 3.9 µm and 1.6 px, respectively.
-The particle images were illuminated using 7 ns pulses from a pair of 400mJ Spectra Physics Nd:YAG lasers.
-The beams of these lasers were aligned co-linearly before being shaped into a sheet aligned to the centre of the tunnel and the corresponding meridian of the sphere.
-The thickness of the laser sheet throughout the field of view was 300 µm.
-Pairs of single-exposed images where captured with a separation of $\Delta t=1$ ms at an acquisition rate of 1 Hz ($\Delta T=1$ s), using a BeagleBone Black pulse generator [40] to control the timing and synchronisation of the camera and laser pulses.
+The f-stop was set to 2.8, yielding a depth of field of 5.0 mm.
+The diffraction-limited minimum image diameter and expected particle image diameter for an 11 µm particle at these settings were 3.9 µm and 1.6 px, respectively.
+Illumination was provided by 7 ns pulses from a pair of 400 mJ Spectra Physics Nd:YAG lasers.
+The beams were aligned collinearly and shaped into a sheet aligned to the tunnel centreline and the corresponding meridian of the sphere.
+The thickness of the laser sheet throughout the field of view was approximately 300 µm.
+Image pairs were captured with a separation of $\Delta t=1$ ms at an acquisition rate of 1 Hz ($\Delta T=1$ s), using a BeagleBone Black pulse generator [14] to synchronise the camera and laser pulses.
 3.3 Data Processing
-The mean background images of the first and second exposures were subtracted from the corresponding single-exposed particle image pairs.
+Mean background images of the first and second exposures were subtracted from the corresponding particle images.
 The instantaneous streamwise and transverse velocities, $u(x,y,t)$ and $v(x,y,t)$, were determined from the particle image pairs using two-component-two-dimensional (2C-2D) multigrid/multipass cross-correlation digital (MCCD) PIV [48, 36].
-A final interrogation window of 64 px $\times$ 32 px on a grid with spacing of 32 px $\times$ 16 px, in $x$ and $y$, respectively, was used.
-This aspect ratio was chosen to capture the higher streamwise velocity component and the higher velocity gradients in the transverse direction, while still being sufficiently large for the interrogation window to contain approximately 10 particles common to both images of each pair.
+The final interrogation window size was 64 px $\times$ 32 px, with a grid spacing of 32 px $\times$ 16 px in the streamwise $x$ and transverse $y$ directions, respectively.
+This aspect ratio was selected to capture the higher streamwise velocity component and the higher velocity gradients in the transverse direction, while ensuring each interrogation window contained approximately 10 particles common to both images of each pair.
 The quality of the cross-correlation was increased through the application of an 8-point Hart filter.
-A maximum displacement limit of 20 px and normalised local median filter [43], with a threshold of 2 standard errors, was applied to the raw displacement vectors.
-The validation criteria resulted in a rejection of 2.3% of vectors, which were replaced by an interpolation of a 2nd-order polynomial fit of their 13 nearest neighbours [42].
+A maximum displacement limit of 20 px and normalised local median filter [47], with a threshold of 2 standard deviations, were applied to the raw displacement vectors.
+The validation criteria resulted in a rejection of 2.3% of vectors, which were replaced by an interpolation using a second-order polynomial fit of their 13 nearest neighbours [36].
 A total of $N_{T}=12,600$ two-component velocity vector fields, each containing 50,000+ vectors, were determined.
 The instantaneous velocity fluctuations $u^{\prime}(x,y,t)$ and $v^{\prime}(x,y,t)$ were calculated by subtracting the mean velocity field, $\overline{u}(x,y)$ and $\overline{v}(x,y)$, from the instantaneous velocity field, $u(x,y,t)$ and $v(x,y,t)$.
-The 95% confidence interval of the 2C-2D MCCDPIV algorithm used is 0.06 pixels [42], which corresponds to a standard error of $\sigma_{\varepsilon_{u}}=0.03$ pixels.
-The standard error of the instantaneous velocities is thus $1.1\times 10^{-3}$ m/s.
+The 95% confidence interval of the 2C-2D MCCDPIV algorithm used is 0.06 pixels [36], corresponding to a standard error of $\sigma_{\varepsilon_{u}}$ = 0.03 pixels.
+This translates to a standard error of $1.1\times 10^{-3}$ m/s in the instantaneous velocities.
 The uncertainty of the mean velocity arising from the random nature of the turbulent fluctuations can be estimated by including the mean streamwise Reynolds stress [3, 37]
@@
 where $\sigma_{\varepsilon_{u}}$ is the standard error of the instantaneous streamwise velocity measurements.
-As the maximum of both $\overline{u^{\prime}u^{\prime}}$ and $\overline{v^{\prime}v^{\prime}}$ are of the order of $0.05U_{\infty}^{2}$ [39], a conservative estimate of the uncertainty in the mean velocity components is $4.0\times 10^{-4}$ m/s.
+As the maxima of both $\overline{u^{\prime}u^{\prime}}$ and $\overline{v^{\prime}v^{\prime}}$ are of the order of $0.05U_{\infty}^{2}$ [11], a conservative estimate of the uncertainty in the mean velocity components is $4.0\times 10^{-4}$ m/s.
 The uncertainty in the fluctuations is thus
@@
 Parameter Symbol Units Relative Units
-Freestream velocity $U_{\infty}$ 200 mm/s --
-Sphere diameter $D$ 40 mm --
-Reynolds number $Re_{D}$ 7780 --
-Convective time $t_{c}$ 0.2 s --
+Freestream velocity $U_{\infty}$ 200 mm/s -
+Sphere diameter $D$ 40 mm -
+Reynolds number $Re_{D}$ 7780 -
+Convective time $t_{c}$ 0.2 s -
 Time between laser pulses $\Delta t$ 1 ms $0.005t_{c}$
 Time between velocity fields $\Delta T$ 1 s $5t_{c}$
-Depth of field -- 5.0 mm $0.125D$
-Laser sheet thickness -- 0.3 mm $0.0075D$
-Diffraction limited particle diameter -- 3.9 µm
-Spatial resolution $SR$ 37 µm --
-Grid spacing -- 32 px $\times$ 16 px $0.03D\times 0.015D$
-Interrogation window -- 64 px $\times$ 32 px $0.06D\times 0.06D$
-Field of view -- 190 mm $\times$ 190 mm $4.7D\times 4.7D$
-Number of 2C-2D velocity fields $N_{T}$ 12,600 --
-Velocity vectors per field $N_{S}$  50,000+ --
-Velocity vector validation rate -- 97.7%
+Depth of field - 5.0 mm $0.125D$
+Laser sheet thickness - 0.3 mm $0.0075D$
+Diffraction-limited particle diameter - 3.9 µm
+Spatial resolution SRSR 37 µm -
+Grid spacing - 32 px $\times$ 16 px $0.03D\times 0.015D$
+Interrogation window - 64 px $\times$ 32 px $0.06D\times 0.06D$
+Field of view - 190 mm $\times$ 190 mm $4.7D\times 4.7D$
+Number of 2C-2D velocity fields $N_{T}$ 12,600 -
+Velocity vectors per field $N_{S}$  50,000+ -
+Velocity vector validation rate - 97.7%
 Instantaneous velocity fluctuation uncertainty $\sigma_{u^{\prime}}$ $1.2\times 10^{-3}$ m/s $0.006U_{\infty}$
@@
 4.1 POD and HPOD Modes
-POD was performed by arranging the turbulent fluctuations into the matrix
+POD was implemented by arranging the turbulent fluctuations into the matrix
 $\boldsymbol{X}=\begin{bmatrix}u^{\prime}(\boldsymbol{x}_{1},t_{1})&\cdots&u^{\prime}(\boldsymbol{x}_{N_{S}},t_{1})&v^{\prime}(\boldsymbol{x}_{1},t_{1})&\cdots&v^{\prime}(\boldsymbol{x}_{N_{S}},t_{1})\\ \vdots&u^{\prime}(\boldsymbol{x}_{n},t_{m})&\vdots&\vdots&v^{\prime}(\boldsymbol{x}_{n},t_{m})&\vdots\\ u^{\prime}(\boldsymbol{x}_{1},t_{N_{T}})&\cdots&u^{\prime}(\boldsymbol{x}_{N_{S}},t_{N_{T}})&v^{\prime}(\boldsymbol{x}_{1},t_{N_{T}})&\cdots&v^{\prime}(\boldsymbol{x}_{N_{S}},t_{N_{T}})\\ \end{bmatrix}$ (30)
-where $N_{S}$ is the number of velocity vectors per velocity field and $\boldsymbol{x}_{n}={(x,y)}_{n}$ is the spatial coordinates of the $n^{th}$ vector, which has fluctuating components denoted by $u^{\prime}$ and $v^{\prime}$ in the streamwise and transverse directions, respectively.
-As $N_{S}>N_{T}$, the method of snapshots was used to perform the POD, with the two-point correlation of $\boldsymbol{X}$ in time given by the matrix form of equation (7)
+where $N_{S}$ denotes the number of velocity vectors per velocity field and $\boldsymbol{x}_{n}={(x,y)}_{n}$ represents the spatial coordinates of the $n^{th}$ vector, with fluctuating components $u^{\prime}$ and $v^{\prime}$ corresponding to the streamwise and transverse directions, respectively.
+Since $N_{S}>N_{T}$, the method of snapshots was employed to perform the POD.
+The two-instant correlation of $\boldsymbol{X}$ is expressed in matrix form as
 $\boldsymbol{C}=\boldsymbol{X}\boldsymbol{X}^{T},$ (31)
@@
 $\boldsymbol{C}\boldsymbol{\Phi}=\lambda\boldsymbol{\Phi},$ (32)
-yielded the eigenvectors $\boldsymbol{\Phi}$ and eigenvalues $\lambda$. The temporal coefficients were determined from $\boldsymbol{\Phi}$
+yields the eigenvectors $\boldsymbol{\Phi}$ and eigenvalues $\lambda$. The temporal coefficients are obtained from $\boldsymbol{\Phi}$
 $a_{k}(t_{m})=\boldsymbol{\Phi}_{m,k},$ (33)
-and the spatial modes were given by the projection of $\boldsymbol{X}$ onto $\boldsymbol{\Phi}$
+and the spatial modes are given by projecting $\boldsymbol{X}$ onto $\boldsymbol{\Phi}$
 $\psi_{k}(\boldsymbol{x}_{n})={(\boldsymbol{\Phi}^{T}\boldsymbol{X})}_{k,n}.$ (34)
-The analytic signal of $\boldsymbol{X}$
-$\boldsymbol{X}^{a}=\boldsymbol{X}+i\mathcal{H}_{x}[\boldsymbol{X}],$ (35)
+As the Fourier transform—and thus the Hilbert transform—assumes periodic signals, the matrix $\boldsymbol{X}$ was weighted in the streamwise direction using a Hann window [5] prior to performing HPOD to reduce spectral leakage in the modes.
+The weight at each streamwise index is defined as
+$w_{\textrm{Hann}}(x)=\tanh\!\left(\alpha\sin^{2}(\pi n/N)\right),$ (35)
+where $\alpha$ is a scaling parameter that controls the steepness of the weighting taper towards the domain edges, $n$ is the index of the streamwise location and $N$ is the total number of streamwise locations.
+A value of $\alpha=10$ was used to provide a relatively steep taper to zero at the domain edges while minimising the effect on the interior.
+The weighted analytic signal of $\boldsymbol{X}$
+$\boldsymbol{X}^{a}=w_{\textrm{Hann}}(x)\boldsymbol{X}+i\mathcal{H}_{x}[w_{\textrm{Hann}}(x)\boldsymbol{X}],$ (36)
 was determined using equations (18) and (19), and HPOD was performed by substituting $\boldsymbol{X}^{a}$ for $\boldsymbol{X}$ in equations (31) to (34).
-The resulting eigenvalues, temporal coefficients and spatial modes are denoted by $\lambda^{a}_{k}$, $a_{k}^{a}(t)$ and $\psi_{k}^{a}(\boldsymbol{x})$, respectively.
-The contribution of the leading HPOD modes to the TKE of $\boldsymbol{X}^{a}$ is significantly larger than the contribution of the leading POD modes to the TKE of $\boldsymbol{X}$.
-The first, second and third modes of each decomposition contribute 7.70%, 4.53% and 3.70%, and 4.8%, 3.39% and 2.55%, respectively.
-However, the difference in the contribution of the higher order modes to the TKE is less significant, as shown in figure 2a.
-The result is that the cumulative contribution to TKE of a given number of the HPOD modes is larger than that of the equivalent number of POD modes, as shown in figure 2b and table 2.
-However, this is not a one-to-one comparison, as the total TKE of the analytic signal $\boldsymbol{X}^{a}$ is larger than that of the data $\boldsymbol{X}$, at 571,871 m2/s2 and 300,542 m2/s2, respectively.
-This is due to the imaginary component of $\boldsymbol{X}^{a}$, which is not present in $\boldsymbol{X}$.
+The resulting eigenvalues, temporal coefficients, and spatial modes are denoted by $\lambda^{a}_{k}$, $a^{a}_{k}(t)$, and $\psi^{a}_{k}(\boldsymbol{x})$, respectively.
+The contribution of the leading HPOD modes to the TKE of $\boldsymbol{X}^{a}$ is significantly larger than that of the leading POD modes to the TKE of $\boldsymbol{X}$.
+The first, second, and third modes of the HPOD contribute 8.93%, 5.36%, and 4.05%, respectively, while those of the POD contribute 4.85%, 3.39%, and 2.55%, respectively.
+However, the difference in the contribution of the higher-order modes to the TKE is less pronounced, as shown in figure 2a.
+Consequently, the cumulative contribution to TKE of a given number of the HPOD modes is more than that of the equivalent number of POD modes, as shown in figure 2b and table 2.
+It should be noted that this is not a one-to-one comparison, as the total TKE of the analytic signal $\boldsymbol{X}^{a}$, 453,071 m2/s2, is greater than that of the data $\boldsymbol{X}$, 300,542 m2/s2.
+This difference arises from the imaginary component of $\boldsymbol{X}^{a}$, which is absent from $\boldsymbol{X}$.
 Figure 2: (a) Individual and (b) cumulative TKE contribution of the POD and HPOD modes.
 TKE [%] POD Modes HPOD Modes Ratio
-25 12 7 0.58
-50 65 38 0.58
-75 311 180 0.58
-90 1027 593 0.58
-95 1973 1172 0.59
-99 5904 4148 0.70
+25 12 6 0.50
+50 65 32 0.49
+75 311 153 0.49
+90 1027 513 0.50
+95 1973 1020 0.52
+99 5904 3753 0.64
 Table 2: Number of POD modes and HPOD modes required to meet 25% to 99% of the TKE of the data $\boldsymbol{X}$ and its analytic signal $\boldsymbol{X}^{a}$, respectively, and the ratio of HPOD modes to POD modes required to meet each proportion of TKE.
 Figure 3: (i) Streamwise and (ii) transverse components of the first ten POD modes.
-[…]
-The sixth and eighth POD modes, shown in figures 3f and 3h, respectively, are similar in structure to the first and second POD modes but with a shorter wavelength than those modes.
+The first POD mode consists of strong fluctuations directly behind the sphere at $x/D<2.5$, with substantially weaker fluctuations present around $x/D=3.5$.
+The streamwise fluctuations, shown in figure 3a(i), are opposite in sign across the centreline, where they vanish, while the transverse fluctuations, shown in figure 3a(ii), alternate in sign along the streamwise direction and are maximal at the centreline.
+The second POD mode, the streamwise and transverse components of which are shown in figures 3b(i) and 3b(ii), respectively, appears similar to the first mode, with the structures shifted downstream.
+The structures present at $x/D<2.5$ in the first mode are now located at $1.5<x/D<3.5$, and new structures with opposite sign are present at $x/D<1.5$, suggesting that these modes represent a propagating structure of fluctuations with opposing sign.
+Although the third POD mode, shown in figure 3c, is asymmetric in the measured plane, it is similar in structure to the first and second modes, suggesting that the structure is not always planar-symmetric.
+The fourth and fifth POD modes, shown in figures 3d and 3e, respectively, both exhibit symmetric V-shaped structures in the streamwise fluctuations and antisymmetric structures in the transverse fluctuations.
+Although the streamwise and transverse components of the fifth POD mode appear more symmetric and antisymmetric, respectively, than those of the fourth mode, it still resembles the fourth mode shifted downstream.
+The sixth and eighth POD modes, shown in figures 3f and 3h, respectively, are similar in structure to the first and second POD modes but with a shorter wavelength.
 The seventh POD mode superficially resembles the fourth mode in its transverse component, shown in figure 3g(ii).
 However, the streamwise component, shown in figure 3g(i), does not exhibit the streamwise periodicity of the other leading modes.
-While this mode may represent an asymmetric aspect of the structure represented by the forth and fifth modes, it is less clearly related to these modes than the third mode appears to be to the first and second modes.
-The ninth and tenth POD modes respectively exhibit V-shaped structures in their streamwise components, shown in figures 3i(i) and 3j(i), respectively, similar to the fourth and fifth modes for $x/D>2.5$, which increase in wavelength further downstream.
+While this mode may represent an asymmetric aspect of the structure represented by the fourth and fifth modes, it is less clearly related to these modes than the third mode appears to be to the first and second modes.
+The ninth and tenth POD modes exhibit V-shaped structures in their streamwise components, shown in figures 3i(i) and 3j(i), respectively, similar to the fourth and fifth modes for $x/D>2.5$, which increase in wavelength further downstream.
 These modes differ in shape from the lower-order modes upstream of $x/D=2$, where the structures have an aspect ratio closer to unity than those in the higher-order modes.
 The transverse components of the ninth and tenth modes, shown in figures 3i(ii) and 3j(ii), respectively, resemble those of the fourth and fifth modes with a short wavelength.
-The wavelength of these modes also increase in the downstream direction.
+The wavelength of these modes also increases in the downstream direction.
 Figure 4: (i) Streamwise and (ii) transverse components of the first four HPOD modes.
+Real parts are shown in (a), (c), (e), and (g) and corresponding imaginary parts are shown in (b), (d), (f), and (h)
 The first HPOD mode, the real and imaginary parts of which are shown in figure 4a and 4b, respectively, is qualitatively similar in structure to the first two POD modes, shown in figure 3a and 3b, respectively.
-The streamwise components of the first HPOD mode, shown in figures 4a(i) and 4b(i) for the real and imaginary parts, respectively, consist of alternating positive and negative regions in the streamwise direction and is antisymmetric about the streamwise axis.
-The transverse components, shown in figures 4a(ii) and 4b(ii) for the real and imaginary parts, respectively, consist of symmetric regions of alternating sign.
-This structure is very similar to that of the first two POD modes, with the exception of the ends of the measurement domain.
-This difference is due to the filtering of the Hilbert transform, which places more emphasis on the components of the structure which propagate in the streamwise direction.
-The effect of this filtering is seen most clearly comparing figures 4b(i) and 3b(i) around $x/D=5$.
-The second HPOD mode features the same V-shaped structure in the streamwise components, the real and imaginary parts of which are shown in figures 4c(i) and 4d(i), respectively, present in the fourth and fifth POD modes, shown in figures 3d(i) and 3e(i), respectively.
-The transverse components of the real and imaginary parts of the second HPOD mode are antisymmetric with regions of alternating sign in the streamwise direction, as shown in figures 4c(ii) and 4d(ii), respectively.
-While similar to transverse components of the fourth and fifth POD modes, shown in figures 3d(ii) and 3e(ii), respectively, the HPOD modes are significantly more symmetrical/anti-symmetrical than the POD modes, due to the filtering effect of the Hilbert transform.
-The third HPOD mode, the real and imaginary components of which are shown in figures 4e and 4f, have a similar structure to the sixth and eight POD modes, shown in figures 3f and 3h, respectively.
-Namely, antisymmetric streamwise velocity regions and symmetric transverse velocity regions of alternating sign in the streamwise direction, with a shorter wavelength than the first HPOD mode and the first two POD modes.
-The fourth HPOD mode, the real and imaginary components of which are shown in figures 4g and 4h, have a similar structure to the ninth and tenth POD modes, shown in figures 3i and 3j, respectively.
-The streamwise components of these modes feature symmetrical ovular regions of alternating sign in the streamwise direction in the near wake becoming V-shaped structures around $x/D=2$, while the transverse components feature antisymmetric regions of alternating sign in the streamwise direction.
+The streamwise component of the first HPOD mode, the real and imaginary components of which are shown in figures 4a(i) and 4b(i), respectively, consists of alternating positive and negative regions in the streamwise direction and is antisymmetric across the centreline.
+The transverse component, the real and imaginary parts of which are shown in figures 4a(ii) and 4b(ii), respectively, consists of symmetric regions of alternating sign.
+This structure is very similar to that of the first two POD modes, except for the ends of the measurement domain, where the intensity is reduced by the Hann window.
+The small structures near the sphere at $x/D=1$, in the second POD mode, shown in figure 3b(i), are extended towards the centreline in the real component of the HPOD mode, similar to the downstream structures, due to the assumption of periodicity implicit in the Hilbert transform.
+In the imaginary part of the HPOD mode, shown in figure 4b(i), these structures are less pronounced as their intensity is reduced by the Hann window.
+The favouring of propagating structures by HPOD also results in an increase in the relative intensity of the downstream structures in figures 4a(i) and 4b(i) compared to those in figures 3a(i) and 3b(i).
+The streamwise component of the second HPOD mode, the real and imaginary parts of which are shown in figures 4c(i) and 4d(i), respectively, features the same V-shaped structure present in the fourth and fifth POD modes, shown in figures 3d(i) and 3e(i), respectively.
+The transverse components of the real and imaginary parts of the second HPOD mode are antisymmetric with regions of alternating sign along the streamwise direction, as shown in figures 4c(ii) and 4d(ii), respectively.
+While similar to the transverse components of the fourth and fifth POD modes, shown in figures 3d(ii) and 3e(ii), respectively, the HPOD modes are significantly more symmetric and antisymmetric across the centreline than the POD modes, due to the accentuation of propagating structures in the HPOD modes.
+The third HPOD mode, the real and imaginary components of which are shown in figures 4e and 4f, respectively, has a similar structure to the sixth and eighth POD modes, shown in figures 3f and 3h, respectively.
+Specifically, it exhibits antisymmetric streamwise velocity regions and symmetric transverse velocity regions of alternating sign along the streamwise direction, with a shorter wavelength than the first HPOD mode and the first two POD modes.
+The fourth HPOD mode, the real and imaginary components of which are shown in figures 4g and 4h, has a similar structure to the ninth and tenth POD modes, shown in figures 3i and 3j, respectively.
+The streamwise component of this mode features symmetric ovular regions of alternating sign in the streamwise direction in the near wake, becoming V-shaped structures around $x/D=2$, while the transverse component features antisymmetric regions of alternating sign in the streamwise direction.
 4.2 Pairing POD Modes using HPOD Modes
-As $\boldsymbol{f}(\boldsymbol{x},t)$ is given by the real part of $\boldsymbol{f}^{a}(\boldsymbol{x},t)$, it can be expressed in terms of the HPOD modes as
-$\boldsymbol{f}(\boldsymbol{x},t)=\Re\big{(}\boldsymbol{f}^{a}(\boldsymbol{x},t)\big{)}=\sum_{k=1}^{K}\Re\big{(}a_{k}^{a}(t)\psi^{a}_{k}(\boldsymbol{x})\big{)},$ (36)
+As $\boldsymbol{f}(\boldsymbol{x},t)$ is given by the real part of $\boldsymbol{f}^{a}(\boldsymbol{x},t)$, it can be expressed using the HPOD modes as
+$\boldsymbol{f}(\boldsymbol{x},t)=\Re\big(\boldsymbol{f}^{a}(\boldsymbol{x},t)\big)=\sum_{k=1}^{K}\Re\big(a^{a}_{k}(t)\psi^{a}_{k}(\boldsymbol{x})\big),$ (37)
 which can be expanded to
-$\boldsymbol{f}(\boldsymbol{x},t)=\Re\big{(}\boldsymbol{f}^{a}(\boldsymbol{x},t)\big{)}=\sum_{k=1}^{K}\Re\big{(}a_{k}^{a}(t)\big{)}\Re\big{(}\psi^{a}_{k}(\boldsymbol{x})\big{)}-\Im\big{(}a_{k}^{a}(t)\big{)}\Im\big{(}\psi_{k}^{a}(\boldsymbol{x})\big{)}.$ (37)
-As the real and imaginary parts of $a_{k}^{a}(t)$ can be expressed as $|a^{a}_{k}(t)|\cos(\phi^{a}_{k}(t))$ and $-|a^{a}_{k}(t)|\sin(\phi^{a}_{k}(t))$, respectively, the shape of the $k^{th}$ HPOD mode at the phase angle $\phi_{k}^{a}(t)$ is given by
-$\frac{\psi_{k}^{a}(\boldsymbol{x},\phi_{k}^{a}(t))}{|a_{k}^{a}(t)|}=\cos\big{(}\phi_{k}^{a}(t)\big{)}\Re\big{(}\psi_{k}^{a}(\boldsymbol{x})\big{)}+\sin\big{(}\phi_{k}^{a}(t)\big{)}\Im\big{(}\psi_{k}^{a}(\boldsymbol{x})\big{)}.$ (38)
+$\boldsymbol{f}(\boldsymbol{x},t)=\Re\big(\boldsymbol{f}^{a}(\boldsymbol{x},t)\big)=\sum_{k=1}^{K}\Re\big(a^{a}_{k}(t)\big)\Re\big(\psi^{a}_{k}(\boldsymbol{x})\big)-\Im\big(a^{a}_{k}(t)\big)\Im\big(\psi^{a}_{k}(\boldsymbol{x})\big).$ (38)
+As the real and imaginary parts of $a^{a}_{k}(t)$ can be expressed as $|a^{a}_{k}(t)|\cos(\phi^{a}_{k}(t))$ and $|a^{a}_{k}(t)|\sin(\phi^{a}_{k}(t))$, respectively, the shape of the $k^{th}$ HPOD mode for the phase angle $\phi^{a}_{k}(t)$ is given by
+$\frac{\psi^{a}_{k}(\boldsymbol{x},\phi^{a}_{k}(t))}{|a^{a}_{k}(t)|}=\cos\big(\phi^{a}_{k}(t)\big)\Re\big(\psi^{a}_{k}(\boldsymbol{x})\big)+\sin\big(\phi^{a}_{k}(t)\big)\Im\big(\psi^{a}_{k}(\boldsymbol{x})\big).$ (39)
 The mode shape of the $k^{th}$ HPOD mode matches the $j^{th}$ POD mode when
 $\psi_{j}(\boldsymbol{x})=\begin{bmatrix}\cos(\Delta\phi^{a}_{k,j})&\sin(\Delta\phi^{a}_{k,j})\\ \end{bmatrix}\begin{bmatrix}\Re\big(\psi^{a}_{k}(\boldsymbol{x})\big)\\ -\Im\big(\psi^{a}_{k}(\boldsymbol{x})\big)\\ \end{bmatrix}$, (40)
-where $\Delta\phi_{k,j}^{a}$ is the phase offset of the $k^{th}$ HPOD mode which results in the best match of the $j^{th}$ POD mode.
-The negative of the imaginary part is used to account for the negative introduced by the Hilbert transform for positive instantaneous frequencies, given in equation (16), which results in the phase of the HPOD modes being reversed relative to the motion of the structure.
-The phase angles at which the $k^{th}$ HPOD mode matches each of the POD modes can be determined simultaneously by
+where $\Delta\phi_{k,j}^{a}$ is the phase offset of the $k^{th}$ HPOD mode which provides the best match to the $j^{th}$ POD mode.
+The negative sign on the imaginary part accounts for the negative introduced by the Hilbert transform associated with positive instantaneous frequencies, given in equation (16), which accounts for the phase of the HPOD modes being reversed relative to the motion of the structure.
+The phase angles at which the $k^{th}$ HPOD mode matches each of the POD modes can be determined by solving
 $\begin{bmatrix}\cos(\Delta\phi^{a}_{k,1})&\sin(\Delta\phi^{a}_{k,1})\\ \vdots&\vdots\\ \cos(\Delta\phi^{a}_{k,j})&\sin(\Delta\phi^{a}_{k,j})\\ \vdots&\vdots\\ \cos(\Delta\phi^{a}_{k,K})&\sin(\Delta\phi^{a}_{k,K})\\ \end{bmatrix}^{T}={\left(\begin{bmatrix}\Re\big(\psi^{a}_{k}(\boldsymbol{x})\big)\\ -\Im\big(\psi^{a}_{k}(\boldsymbol{x})\big)\\ \end{bmatrix}\begin{bmatrix}\Re\big(\psi^{a}_{k}(\boldsymbol{x})\big)\\ -\Im\big(\psi^{a}_{k}(\boldsymbol{x})\big)\\ \end{bmatrix}^{T}\right)}^{-1}\begin{bmatrix}\Re\big(\psi^{a}_{k}(\boldsymbol{x})\big)\\ -\Im\big(\psi^{a}_{k}(\boldsymbol{x})\big)\\ \end{bmatrix}\begin{bmatrix}\psi_{1}(\boldsymbol{x})\\ \vdots\\ \psi_{j}(\boldsymbol{x})\\ \vdots\\ \psi_{K}(\boldsymbol{x})\\ \end{bmatrix}^{T},$ (41)
@@
 $\Delta\phi^{a}_{k,j}=\tan^{-1}\left(\frac{\sin(\Delta\phi^{a}_{k,j})}{\cos(\Delta\phi^{a}_{k,j})}\right).$ (42)
-The correlation coefficient of the $k^{th}$ HPOD mode at $\Delta\phi^{a}_{k,j}$ and the $j^{th}$ POD mode is
-$\mathcal{R}_{k,j}=\frac{\operatorname{cov}\big{(}\psi_{k}^{a}(\boldsymbol{x},\Delta\phi^{a}_{k,j}),\psi_{j}(\boldsymbol{x})\big{)}}{\sqrt{\operatorname{var}\big{(}\psi_{k}^{a}(\boldsymbol{x},\Delta\phi^{a}_{k,j})\big{)}\operatorname{var}\big{(}\psi_{j}(\boldsymbol{x},)\big{)}}},$ (42)
-and the corresponding $R^{2}$ value is
-$R^{2}_{k,j}=1-\frac{\sum\big{(}\psi_{k}^{a}(\boldsymbol{x},\Delta\phi^{a}_{k,j})-\psi_{j}(\boldsymbol{x})\big{)}^{2}}{\sum\big{(}\psi_{j}(\boldsymbol{x})\big{)}^{2}}.$ (43)
-The HPOD modes which best match each of the first ten POD modes, and the corresponding phase offset, correlation coefficient, and $R^{2}$ value, are shown in table 3.
-POD Mode, $j$ HPOD Mode, $k$ $\Delta\phi_{k,j}^{a}$ $\mathcal{R}_{k,j}$ $R^{2}_{k,j}$
-1 1 $1.71\pi$ 0.89 0.78
-2 1 $1.18\pi$ 0.97 0.93
-3 3 $0.18\pi$ 0.62 0.26
-4 2 $1.70\pi$ 0.84 0.65
-5 2 $1.21\pi$ 0.96 0.91
-6 3 $1.52\pi$ 0.82 0.65
-7 4 $1.70\pi$ 0.63 0.29
-8 3 $0.88\pi$ 0.76 0.50
-9 4 $1.10\pi$ 0.85 0.70
-10 4 $0.50\pi$ 0.80 0.49
-Table 3: HPOD modes and phase shifts for best match of the first ten POD modes, and the corresponding $\mathcal{R}_{k,j}$ values.
-The first and second POD modes are the best matches to the first HPOD mode, with correlation coefficients of 0.89 and 0.97, and $R^{2}$ values of 0.78 and 0.93, respectively.
-The corresponding phase angles are $1.71\pi$ and $1.18\pi$, which are approximately $90^{\circ}$ apart.
-This is consistent with the separation of the real and imaginary parts of the HPOD modes, suggesting that the relationship between these POD modes is similar to that of the real and imaginary parts of the first HPOD mode.
-Similarly, the fourth and fifth POD modes are the best matches to the second HPOD mode, with correlation coefficients of 0.84 and 0.96, and $R^{2}$ values of 0.65 and 0.91, respectively.
-The corresponding phase angles are $1.7\pi$ and $1.21\pi$, which are also separated by approximately $90^{\circ}$.
-The sixth and eighth POD modes are the best match to the third HPOD mode, with correlation coefficients of 0.82 and 0.76, and $R^{2}$ values of 0.65 and 0.50, respectively.
-The separation of the corresponding phase angles, $1.52\pi$ and $0.88\pi$, respectively, is greater than $90^{\circ}$.
-This is also the case for the ninth and tenth POD modes, which are the best matches to the fourth HPOD mode at phase angles of $1.10\pi$ and $0.50\pi$, respectively.
-The corresponding correlation coefficients and $R^{2}$ values of these POD modes are 0.85 and 0.80, and 0.70 and 0.49, respectively.
-The phase average of the paired POD modes were compared to the corresponding HPOD modes at phase angles from 0 to $2\pi$ in increments of $0.01\pi$, accounting for the phase angle shift in the first POD mode in each pair, $\Delta\phi_{k,i}$.
-The $R^{2}$ values of each phase average
-$R_{k,(i,j)}^{2}=1-\frac{\sum{\Big{(}\psi^{a}_{k}(\boldsymbol{x},\phi+\Delta\phi^{a}_{k,i})-\widetilde{\psi_{i,j}}(\boldsymbol{x},\phi)\Big{)}}^{2}}{\sum{\big{(}\widetilde{\psi_{i,j}}(\boldsymbol{x},\phi)\big{)}}^{2}},$ (44)
-was calculated for each case.
-The joint probability density functions of the phase angle of the HPOD modes and the phase angle of the corresponding pairs of POD modes, adjusted for the phase angle shift between the HPOD mode and the first POD mode of each pair.
-HPOD Mode $k$ POD Modes $i,j$ $R_{k,(i,j)}^{2}$
-1 1,2 0.81
-2 4,5 0.76
-3 6,8 0.57
-4 9,10 0.61
-Table 4: $R_{k,(i,j)}^{2}$ values of the phase-averaged HPOD modes and the corresponding POD modes.
+The correlation coefficient between the $k^{th}$ HPOD mode at $\Delta\phi^{a}_{k,j}$ and the $j^{th}$ POD mode is given by
+$\mathcal{R}_{k,j}(\Delta\phi^{a}_{k,j})=\frac{\operatorname{cov}\big(\psi^{a}_{k}(\boldsymbol{x},\Delta\phi^{a}_{k,j}),\psi_{j}(\boldsymbol{x})\big)}{\sqrt{\operatorname{var}\big(\psi^{a}_{k}(\boldsymbol{x},\Delta\phi^{a}_{k,j})\big)\operatorname{var}\big(\psi_{j}(\boldsymbol{x})\big)}},$ (43)
+and the corresponding $R^{2}_{k,j}$ value is
+$R^{2}_{k,j}(\Delta\phi^{a}_{k,j})=1-\frac{\sum_{\boldsymbol{x}}{\Big(\psi^{a}_{k}(\boldsymbol{x},\Delta\phi_{k,j}^{a})-\psi_{j}(\boldsymbol{x})\Big)}^{2}}{\sum_{\boldsymbol{x}}{\Big(\psi_{j}(\boldsymbol{x})\Big)}^{2}},$ (44)
+where $\boldsymbol{x}$ represents the spatial coordinates of the modes.
+The $\mathcal{R}_{k,j}$ and $R_{k,j}^{2}$ values were calculated for the combined streamwise and transverse components of the modes, as well as individually for each component, denoted by superscripts u and v for the streamwise and transverse components, respectively.
+The HPOD modes which best match each of the first ten POD modes, and the corresponding phase offset $\Delta\phi^{a}_{k,j}$, correlation coefficient $\mathcal{R}_{k,j}$, and $R_{k,j}^{2}$ value, are shown in table 3.
+POD Mode, $j$ HPOD Mode, $k$ $\Delta\phi_{k,j}^{a}$ $\mathcal{R}_{k,j}$ $\mathcal{R}_{k,j}^{u}$ $\mathcal{R}_{k,j}^{v}$ $R^{2}_{k,j}$ $(R^{2}_{k,j})^{u}$ $(R^{2}_{k,j})^{v}$
+1 1 $1.64\pi$ 0.88 0.86 0.90 0.74 0.71 0.75
+2 1 $1.15\pi$ 0.98 0.96 0.99 0.95 0.92 0.96
+3 3 $0.65\pi$ 0.57 0.46 0.66 0.28 0.10 0.38
+4 2 $1.71\pi$ 0.86 0.85 0.88 0.68 0.70 0.74
+5 2 $1.21\pi$ 0.96 0.95 0.97 0.84 0.87 0.83
+6 3 $1.49\pi$ 0.87 0.83 0.92 0.75 0.68 0.80
+7 4 $0.21\pi$ 0.56 0.59 0.56 0.15 0.20 0.27
+8 3 $0.92\pi$ 0.80 0.65 0.90 0.60 0.35 0.78
+9 4 $1.67\pi$ 0.86 0.84 0.91 0.73 0.63 0.83
+10 4 $1.16\pi$ 0.86 0.86 0.92 0.68 0.57 0.85
+Table 3: HPOD modes and phase shifts for the best match of the first ten POD modes, and the corresponding $\mathcal{R}_{k,j}$ values.
+Superscripts u and v denote the streamwise and transverse components, respectively.
+The correlation coefficient between the $j^{th}$ POD mode and the $k^{th}$ HPOD mode as a function of streamwise position is given by
+$\mathcal{R}_{k,j}(x,\Delta\phi^{a}_{k,j})=\frac{\operatorname{cov}\big(\psi^{a}_{k}(x,\Delta\phi^{a}_{k,j}),\psi_{j}(x)\big)}{\sqrt{\operatorname{var}\big(\psi^{a}_{k}(x,\Delta\phi^{a}_{k,j})\big)\operatorname{var}\big(\psi_{j}(x)\big)}},$ (45)
+and are shown for the matched HPOD modes and POD modes in figure 5.
+Figure 5: Correlation coefficient between each paired POD mode and the corresponding HPOD mode at the best match phase angle as a function of streamwise position.
+Superscripts u and v denote the streamwise and transverse components, respectively.
+The first and second POD modes exhibit the strongest correspondence to the first HPOD mode, with correlation coefficients of 0.88 and 0.98, and $R_{k,j}^{2}$ values of 0.74 and 0.95, respectively.
+The corresponding phase angles are $1.64\pi$ and $1.15\pi$, which are approximately $\pi/2$ apart and consistent with the separation of the real and imaginary parts of the HPOD modes.
+This suggests that the relationship between these POD modes mirrors that of the real and imaginary components of the first HPOD mode.
+The correlation coefficient between these POD and HPOD modes is consistently close to unity, with troughs in the correlation occurring due to the values of the modes approaching zero at $x/D=2.6$ and $x/D=4.4$, as shown in figure 5a.
+Additional reductions occur near the domain boundaries, with the trough at the upstream end being significantly deeper.
+While the correlation of the combined and streamwise components decreases to approximately 0.5, the correlation of the transverse component approaches -1.
+This indicates that the transverse components of the first POD mode and the first HPOD mode have opposite signs close to the sphere.
+This discrepancy likely results from a combination of spectral leakage at the domain edge and the difference in length of the structures in the HPOD mode and POD modes, exemplified by figures 6c(i) and 6d(i).
+The reduction in correlation at the end of the domain is less severe, and most likely reflects the attenuation of the Hann window.
+Similarly, the correlation coefficient of the second POD mode with the first HPOD mode is consistently high throughout the domain, with small dips in the correlation coefficient occurring at $x/D=1.55$ and $x/D=3.5$, as shown in figure 5b.
+The spacing between drops in the correlation coefficient between the first HPOD mode and both the first and second POD modes is approximately $1.9D$, suggesting that these modes correspond to the same propagating structure, with a wavelength of approximately $3.8D$.
+The fourth and fifth POD modes are the best matches to the second HPOD mode, with correlation coefficients of 0.86 and 0.96, and $R_{k,j}^{2}$ values of 0.68 and 0.84, respectively.
+The corresponding phase angles are $1.71\pi$ and $1.21\pi$, which are consistent with the expected $\pi/2$ offset.
+The correlation of the fourth POD mode with the second HPOD mode at the corresponding phase offset, shown in figure 5c, is close to unity between $x/D=1.25$ and $x/D=2.4$, after which it drops before increasing again from approximately $x/D=2.7$.
+The downstream increase never reaches unity because the fourth POD mode is asymmetric about the centreline, as shown in figure 3d, whereas the second HPOD mode is symmetric, as shown in figures 4c–d.
+The correlation of the fifth POD mode with the second HPOD mode at the corresponding phase offset, shown in figure 5d, is consistently high across the measurement domain, except towards the edges of the domain, where the Hann window attenuates the intensity of the HPOD modes.
+This behaviour reflects the greater symmetry of the fifth POD mode, shown in figure 3e, compared to the fourth POD mode, shown in figure 3d.
+The sixth and eighth POD modes are the best matches to the third HPOD mode, with correlation coefficients of 0.87 and 0.80, and $R_{k,j}^{2}$ values of 0.75 and 0.60, respectively.
+The separation of the corresponding phase angles, $1.49\pi$ and $0.92\pi$, is slightly greater than $\pi/2$.
+The correlation between the sixth POD mode and the third HPOD mode at the corresponding phase offset, shown in figure 5e, remains close to unity across the measurement domain with regularly spaced troughs where the mode values approach zero up to $x/D=3$.
+Downstream of this point, the correlation remains at unity for a longer streamwise distance before decreasing towards the domain edge.
+This extended region of high correlation corresponds to the increased size of the structures in the sixth POD mode at this streamwise location, shown in figure 3f.
+The substantial decrease in correlation towards the downstream end reflects the reduced size of the structures in the HPOD mode, and the subsequent regions of opposing sign which are visible at the end of the domain in figure 4f(i).
+The correlation between the eighth POD mode and the third HPOD mode at the corresponding phase offset, shown in figure 5f, is consistently high across the measurement domain, with periodic troughs which increase in wavelength downstream, consistent with the structures in the eighth POD mode, shown in figure 3h(i).
+The ninth and tenth POD modes are the best matches to the fourth HPOD mode at phase angles of $1.67\pi$ and $1.16\pi$, respectively.
+The corresponding correlation coefficients are both 0.86, while the $R_{k,j}^{2}$ values are 0.73 and 0.68, respectively.
+The correlation between the ninth POD mode and the fourth HPOD mode at the corresponding phase offset, shown in figure 5g, is consistently high across the measurement domain, with troughs where the mode values approach zero.
+Additionally, the correlation drops towards the end of the measurement domain, with the streamwise component approaching zero at the end of the domain and the transverse component changing sign, as shown in figures 3i(i) and 3i(ii), respectively.
+The correlation between the tenth POD mode and the fourth HPOD mode at the corresponding phase offset, shown in figure 5h, is also consistently high across the measurement domain.
+The troughs in the correlation are more widely spaced further downstream, consistent with the increasing wavelength of the structures in the tenth POD mode, shown in figure 3j.
+The correlation also drops towards the ends of the measurement domain, with the transverse component changing sign at the beginning of the domain.
+The correlation between the streamwise components becomes negative at the end of the domain, reflecting structures of opposite sign appearing in the HPOD modes, due to the spectral leakage visible in figure 4g.
+The $R^{2}_{k,\widetilde{(i,j)}}$ values of the $k^{th}$ HPOD mode and the phase average of the $i^{th}$ and $j^{th}$ POD modes are defined as
+$R^{2}_{k,\widetilde{(i,j)}}=1-\frac{\sum_{\boldsymbol{x},\phi}{\Big(\psi^{a}_{k}(\boldsymbol{x},\phi+\Delta\phi_{k,i}^{a})-\widetilde{\psi_{i,j}}(\boldsymbol{x},\phi)\Big)}^{2}}{\sum_{\boldsymbol{x},\phi}{\Big(\widetilde{\psi_{i,j}}(\boldsymbol{x},\phi)\Big)}^{2}},$ (46)
+[…]
+HPOD Mode $k$ POD Modes $i,j$ $R^{2}_{k,\widetilde{(i,j)}}$ $(R^{2}_{k,\widetilde{(i,j)}})^{u}$ $(R^{2}_{k,\widetilde{(i,j)}})^{v}$
+1 1,2 0.83 0.79 0.85
+2 4,5 0.80 0.78 0.82
+3 6,8 0.67 0.53 0.76
+4 9,10 0.70 0.61 0.83
+Table 4: $R_{k,(i,j)}^{2}$ values of the phase-averaged HPOD modes and the corresponding POD modes. Superscripts u and v refer to the streamwise and transverse components, respectively.
 Figure 6: Phase-averaged (a) streamwise and (c) transverse components of the first and second POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the first HPOD mode, adjusted for the phase angle offset $\Delta\phi_{1,1}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
-The streamwise component of the phase-averaged first and second POD modes, shown on figure 5a, is good agreement with that of the phase average of the first HPOD mode at the adjusted phase angle, shown in figure 5b, with the main differences in the structures being a result of the filtering from the Hilbert transform.
-Namely, the intensity of the structures in the POD mode is higher directly behind the sphere than the HPOD modes, while the intensity in the structures further downstream is higher in the HPOD modes.
-This difference is also seen in the transverse comments, shown in figures 5c and 5d for the POD modes and HPOD modes, respectively.
-Additionally, the structures in the near wake of the sphere in the POD modes, shown in figure 5a(iv), are filtered out of the HPOD mode, shown in figure 5b(iv), as they do not propagate downstream.
-Despite these differences, it is clear that the first and second POD modes and the first HPOD mode represent the same propagating structure consisting of planar antisymmetric streamwise fluctuations and alternating positive and negative transverse fluctuations.
-This structure is consistent with flapping of the wake, and the similarity of the third POD mode, shown in figure 3c, to the first and second suggests that the flapping also has an out-of-plane component.
+See supplementary movie 1 for an animated version of this figure.
+The streamwise component of the phase-averaged first and second POD modes, shown in figure 6a, agrees well with that of the first HPOD mode at the adjusted phase angle, shown in figure 6b.
+The main differences in these structures result from the accentuation of propagating structures by HPOD.
+Specifically, the structures in the POD mode exhibit greater intensity directly behind the sphere, while those further downstream are more pronounced in the HPOD modes.
+This difference is also evident in the transverse components, shown in figures 6c and 6d for the POD modes and HPOD modes, respectively.
+Additionally, the shape of the structures in the near wake of the sphere in the POD modes, shown in figure 6a(iv), differs from that in the HPOD mode, shown in figure 6b(iv).
+This results from the assumed periodicity of the Hilbert transform, which couples the near-wake structures with those further downstream.
+Despite these differences, it is evident that the first and second POD modes and the first HPOD mode represent the same propagating structure.
+This structure consists of planar-antisymmetric streamwise fluctuations and alternating positive and negative transverse fluctuations, consistent with wake flapping.
+The similarity of the third POD mode, shown in figure 3c, to the first and second suggests that the flapping includes an out-of-plane component.
 Figure 7: Phase-averaged (a) streamwise and (c) transverse components of the fourth and fifth POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the second HPOD mode, adjusted for the phase angle offset $\Delta\phi_{2,4}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
-While the streamwise component of the phase-averaged fourth and fifth POD modes, shown in figure 6a, is similar to that of the second HPOD mode, shown in figure 6b, the latter demonstrates significantly stronger symmetry, due to the inherent filtering of propagating features by HPOD.
-This is most clearly seen when comparing figures 6a(ii) and 6a(vi) to figures 6b(ii) and 6b(vi), respectively, where the V-shaped structures in the HPOD case remain symmetric downstream, while they become asymmetric in the case of the POD modes.
-The transverse component of the second HPOD mode, shown in figure 6d, shows a stronger antisymmetry than the transverse components of the phase average of the fourth and fifth POD modes, shown in figure 6c.
-The variation between the phase average of the fourth and fifth POD modes and the second HPOD mode highlights the filtering effect of HPOD, which prioritises the structures that are periodic in the direction of the Hilbert transform.
-While the fourth and fifth POD modes are highly asymmetric, particularly after $x/D>3$, the second HPOD mode is symmetric in the streamwise component and antisymmetric in the transverse direction.
-The asymmetry in the POD modes is likely due to the three-dimensionality in the flow being more significant in the structure represented by the fourth and fifth POD modes than that represented by the first and second POD modes.
-In this case, the use of HPOD is beneficial for isolating the propagating structures in the flow in the presence of asymmetry in the POD modes.
-The alternating positive and negative streamwise fluctuations in these modes suggests that they represent a pulsating motion in the wake of the sphere.
+See supplementary movie 2 for an animated version of this figure.
+While the streamwise component of the phase-averaged fourth and fifth POD modes, shown in figure 7a, resembles the second HPOD mode, shown in figure 7b, the latter exhibits markedly greater symmetry, attributed to HPOD’s inherent emphasis on propagating features.
+This is most evident when comparing figures 7a(ii) and 7a(vi) to figures 7b(ii) and 7b(vi).
+In the HPOD case, the V-shaped structures remain symmetric downstream, whereas they become asymmetric in the POD modes.
+Additionally, the V-shaped structures extend further upstream structures close to the sphere ($x/D<1.5$) in the POD modes, but remain distinct in the HPOD modes.
+The transverse component of the second HPOD mode, shown in figure 7d, exhibits greater antisymmetry than the transverse components of the phase average of the fourth and fifth POD modes, shown in figure 7c, particularly upstream of $x/D=3$.
+The asymmetry in the POD modes likely arises because three-dimensionality in the flow is more pronounced in the structures represented by the fourth and fifth POD modes than that represented by the first and second POD modes.
+In this case, the use of HPOD proves advantageous for isolating the propagating structures in the flow in the presence of asymmetry in the POD modes.
+The alternating positive and negative streamwise fluctuations in these modes indicates that they correspond to a pulsating motion in the sphere’s wake.
 Figure 8: Phase-averaged (a) streamwise and (c) transverse components of the sixth and eighth POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the third HPOD mode, adjusted for the phase angle offset $\Delta\phi_{3,6}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
-The streamwise component of the phase average of the sixth and eight POD modes, shown in figure 7a, is qualitatively similar to that of the third HPOD mode, shown in figure 7b.
-However, the alternating positive and negative structures are better defined in the HPOD mode than the phase average of the POD modes.
-This is most evident when comparing figures 7a(i) and 7a(ii) to figures 7b(i) and 7b(ii), in which the positive region in the HPOD modes around $x/D=3$ and $y/D>0$ is not present in the phase-averaged POD modes.
-This is most likely due to the structure moving in the out-of-plane direction as it propagates, which is filtered out by the Hilbert transform.
-Additionally, the intensity of the structures is generally lower in the HPOD modes than those in the POD modes due to the filtering of the non-propagating components.
-The transverse component of the phase average of the sixth and eighth POD modes, shown in figure 7c, shows a transverse shifting of the structures of alternating sign, particularly in figures 7c(i)-(iii).
-This shifting is not present in the transverse component of the third HPOD mode, shown in figure 7d.
-The shifting of the structures in the transverse velocity of the POD modes is consistent with the the three-dimensional motion of the wake, and coupled with the behaviour of the streamwise components of the POD modes.
-The structure represented by these modes appears to be a flapping in the wake with a shorter wavelength than that represented by the first and second POD modes and the first HPOD mode.
+See supplementary movie 3 for an animated version of this figure.
+The streamwise component of the phase average of the sixth and eighth POD modes, shown in figure 8a, is qualitatively similar to that of the third HPOD mode, shown in figure 8b.
+However, the alternating positive and negative structures are more distinctly defined in the HPOD mode than the phase average of the POD modes.
+This is particularly evident when comparing figures 8a(i) and 8a(ii) to figures 8b(i) and 8b(ii), in which the positive region in the HPOD modes around $x/D=3$ and $y/D>0$ is not present in the phase-averaged POD modes.
+This likely attributable to the out-of-plane motion of the structure as it propagates, while the Hilbert transform enforces periodicity.
+Additionally, the intensity of the structures tends to be lower in the HPOD modes than those in the POD modes because the non-propagating components are relegated from the leading HPOD modes.
+The transverse component of the phase average of the sixth and eighth POD modes, shown in figure 8c, exhibits transverse displacement of alternating-sign structures, most evident in figures 8c(i)–(iii).
+This shifting is absent from the transverse component of the third HPOD mode, shown in figure 8d.
+The shifting of the structures in the transverse velocity of the POD modes aligns with the three-dimensional motion of the wake and corresponds to the observed behaviour of their streamwise components.
+The structure represented by these modes suggests wake flapping with a shorter wavelength than that associated with the first and second POD modes and the first HPOD mode.
 Figure 9: Phase-averaged (a) streamwise and (c) transverse components of the ninth and tenth POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the fourth HPOD mode, adjusted for the phase angle offset $\Delta\phi_{4,9}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
-The streamwise component of the phase-averaged ninth and tenth POD modes, shown in figure 8a, is qualitatively similar to the phase average of the fourth HPOD mode, shown in figure 8b, for $x/D>2.5$.
-In this domain, the streamwise components of both the POD modes and HPOD mode exhibit V-shaped structures similar to those in the fourth and fifth POD modes and the second HPOD mode.
-For $x/D<2.5$, there is more detail in the structure of the POD modes, which is most noticeable in figures 8a(iii) and 8a(iv) compared to figures 8b(iii) and 8b(iv).
-As the V-shaped structures in the streamwise components are more prominent in the phase-averaged ninth and tenth POD modes than the fourth HPOD mode, it can be inferred that they do not propagate in the streamwise direction.
-The transverse components of phase average of the ninth and tenth POD modes and the phase average of the fourth HPOD modes, shown in figures 8c and 8d, respectively, are also qualitatively similar.
-As is the case for the streamwise components, the transverse component of the phase-averaged POD modes contains features not present in the HPOD mode in the very near wake, due to the filtering effect of the Hilbert transform.
-This can be seen clearly in figures 8c(ii) and 8c(iii) compared to figures 8d(ii) and 8d(iii).
+See supplementary movie 4 for an animated version of this figure.
+The streamwise component of the phase-averaged ninth and tenth POD modes, shown in figure 9a, is qualitatively comparable to the phase average of the fourth HPOD mode, shown in figure 9b, for $x/D>2.5$.
+In this domain, the streamwise components of the POD and HPOD modes exhibit V-shaped structures similar to those in the fourth and fifth POD modes and the second HPOD mode.
+For $x/D<2.5$, the structure of the POD modes exhibits greater detail, most evident in figures 9a(iii) and 9a(iv) compared with figures 9b(iii) and 9b(iv).
+The transverse components of the phase average of the ninth and tenth POD modes and the phase average of the fourth HPOD mode, shown in figures 9c and 9d, respectively, also exhibit qualitative similarity.
+Similar to the streamwise components, the transverse component of the phase-averaged POD modes includes features absent from the HPOD mode in the immediate near wake, attributable to the assumed periodicity of the Hilbert transform.
+This is clearly evident in figures 9c(ii) and 9c(iii) compared to figures 9d(ii) and 9d(iii) around $x/D=1$.
 Figure 10: Joint probability density function of the phase angles of (a) the first HPOD mode and the first and second POD modes, (b) the second HPOD mode and the fourth and fifth POD modes, (c) the third HPOD mode and the sixth and eighth POD modes, and (d) the fourth HPOD mode and the ninth and tenth POD modes.
-Phase angles are relative to the phase angle of each mode in the first instantaneous velocity field.
-The joint probability density functions of the HPOD modes and the corresponding POD mode pairs, shown in figure 9, are all the greatest around $\phi^{a}_{k}=\phi_{i,j}$.
-The phase angles of both have been adjusted by the phase angle of the first instantaneous velocity field to account for the phase shift between the POD modes and the HPOD modes and to ensure all phase angles are in the domain $\phi\in[0,2\pi]$.
+Phase angles of the HPOD modes are relative to the phase offset to the first POD mode of the corresponding pair.
+The joint probability density functions of the HPOD modes and the corresponding POD mode pairs, shown in figure 10, are highest around $\phi^{a}_{k}=\phi_{i,j}$ when adjusted for the appropriate phase offset.
+This indicates that the phase relationship between each HPOD mode and its corresponding POD mode pair is consistent throughout the measurement domain.
+The spread of the distributions varies between the different mode pairs, with the tightest clustering occurring for the first HPOD mode and the first and second POD modes, shown in figure 10a, and the widest spread being that of the third HPOD mode and the sixth and eighth POD modes, shown in figure 10c.
+This is consistent with the $R^{2}_{k,\widetilde{(i,j)}}$ values of the mode shapes summarised in table 4, with the highest value of 0.83 occurring for the first HPOD mode and the first and second POD modes, and the lowest value of 0.67 occurring for the third HPOD mode and the sixth and eighth POD modes.
 4.3 Pairing POD Modes using the Hilbert Transform
-While HPOD modes can be used to determine matched pairs of POD modes, this process requires performing both decompositions.
-As the Hilbert transform can be applied in the streamwise direction, it can be applied directly to the POD modes despite their lack of temporal information.
-The result is a shift of the fundamental wavenumbers of each POD mode by $\mp 90^{\circ}$, which is analogous to a propagation of the mode downstream.
-Thus, the streamwise Hilbert transform of a POD mode is analogous to a POD mode which represents a $\mp 90^{\circ}$ phase shift of the same streamwise periodic structure.
-In order to determine matched pairs of POD modes using the Hilbert transform, the Hilbert transform of each POD mode $\psi_{k}(\boldsymbol{x})$ in the streamwise direction $\mathcal{H}_{x}[\psi_{k}(\boldsymbol{x})]$ is used in place of the imaginary part of the HPOD modes to define the analytic signal of $\psi_{k}(\boldsymbol{x})$
+While HPOD modes can be used to identify matched pairs of POD modes, this process requires performing both decompositions.
+Because the Hilbert transform can be applied in the streamwise direction, it can be applied directly to the POD modes, despite their lack of temporal information.
+The result is a $\mp\pi/2$ phase shift of the fundamental Fourier components of each POD mode, analogous to the downstream propagation of the mode.
+Thus, the streamwise Hilbert transform of a POD mode is analogous to a POD mode that represents a $\mp\pi/2$ phase shift of the same streamwise periodic structure.
+In order to determine matched pairs of POD modes using the Hilbert transform, the Hilbert transform of the $k^{th}$ POD mode in the streamwise direction $\mathcal{H}_{x}[\psi_{k}(\boldsymbol{x})]$ is used to define the analytic signal of $\psi_{k}(\boldsymbol{x})$
 ${(\psi_{k})}^{a}(\boldsymbol{x})=\psi_{k}(\boldsymbol{x})+i\mathcal{H}_{x}\big[\psi_{k}(\boldsymbol{x})\big].$ (47)
@@
 The corresponding correlation coefficient is
-$\mathcal{R}_{k^{a},j}=\frac{\operatorname{cov}\big{(}(\psi_{k})^{a}(\boldsymbol{x},\Delta\phi_{k,j}),\psi_{j}(\boldsymbol{x})\big{)}}{\sqrt{\operatorname{var}\big{(}(\psi_{k})^{a}(\boldsymbol{x},\Delta\phi_{k,j})\big{)}\operatorname{var}\big{(}\psi_{j}(\boldsymbol{x})\big{)}}}.$ (49)
-The $R^{2}$ value of the phase average of the $k^{th}$ and $j^{th}$ POD modes compared to the phase average of the analytic signal of the $k^{th}$ POD mode is given by
-$R^{2}_{k^{a},\widetilde{(k,j)}}=1-\frac{\sum\big{(}(\psi_{k})^{a}(\boldsymbol{x},\phi)-\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)\big{)}^{2}}{\sum\big{(}\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)\big{)}^{2}},$ (50)
-where $(\psi_{k})^{a}(\boldsymbol{x},\phi)$ is the analytic signal of the $k^{th}$ POD mode and $\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)$ is the phase average of the $k^{th}$ and $j^{th}$ POD modes at phase angle $\phi$, which was determined using phase angles $\phi=[0,2\pi]$ in increments of $0.01\pi$.
-The POD mode which best matches the analytic signal of the each of the first ten POD modes, the corresponding correlation coefficient, and the $R^{2}$ of the phase averages are given in table 5.
-POD Mode $k$ Paired POD Mode $j$ $\Delta\phi_{k,j}$ $\mathcal{R}_{k^{a},j}$ $R^{2}_{k^{a},\widetilde{(k,j)}}$
-1 2 $-0.5\pi$ 0.88 0.89
-2 1 $0.5\pi$ 0.76 0.76
-3 2 $0.5\pi$ 0.70 0.74
-4 5 $-0.5\pi$ 0.79 0.81
-5 4 $0.5\pi$ 0.72 0.72
-6 8 $-0.5\pi$ 0.80 0.82
-7 9 $-0.5\pi$ 0.60 0.67
-8 6 $0.5\pi$ 0.76 0.77
-9 10 $-0.5\pi$ 0.64 0.65
-10 9 $0.5\pi$ 0.63 0.65
-Table 5: Paired POD modes using the analytic signal of the $k^{th}$ POD mode, and the corresponding phase angle, correlation coefficient and $R^{2}$ of the phase averages.
-For each of the first ten POD modes, the paired POD mode corresponds to the analytic signal at a phase angle of $\mp 0.5\pi$.
-Thus, the $j^{th}$ POD mode best matches the Hilbert transform of the corresponding $k^{th}$ POD mode.
-A pair of POD modes representing a structure propagating in the streamwise direction therefore requires that the POD modes are a mutual best match with phase angle shifts of $0.5\pi$ with opposite signs.
-Thus, the POD mode pairs identified using the Hilbert transform are the first and second POD modes, the fourth and fifth POD modes, the sixth and eighth POD modes, and the ninth and tenth POD modes.
+$\mathcal{R}_{k^{a},j}(\Delta\phi_{k,j})=\frac{\operatorname{cov}\big((\psi_{k})^{a}(\boldsymbol{x},\Delta\phi_{k,j}),\psi_{j}(\boldsymbol{x})\big)}{\sqrt{\operatorname{var}\big((\psi_{k})^{a}(\boldsymbol{x},\Delta\phi_{k,j})\big)\operatorname{var}\big(\psi_{j}(\boldsymbol{x})\big)}}.$ (51)
+The $R^{2}_{k^{a},\widetilde{(k,j)}}$ value of the phase average of the $k^{th}$ and $j^{th}$ POD modes compared to the phase average of the analytic signal of the $k^{th}$ POD mode is given by
+$R^{2}_{k^{a},\widetilde{(k,j)}}=1-\frac{\sum_{\boldsymbol{x},\phi}\big((\psi_{k})^{a}(\boldsymbol{x},\phi)-\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)\big)^{2}}{\sum_{\boldsymbol{x},\phi}\big(\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)\big)^{2}},$ (52)
+where $\boldsymbol{x}$ denotes the spatial coordinates, $(\psi_{k})^{a}(\boldsymbol{x},\phi)$ is the analytic signal of the $k^{th}$ POD mode and $\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)$ is the phase average of the $k^{th}$ and $j^{th}$ POD modes at phase angle $\phi$, which was determined using phase angles $\phi=[0,2\pi]$ in increments of $0.01\pi$.
+The POD mode which best matches the analytic signal of each of the first ten POD modes, is given in table 5, along with the corresponding correlation coefficient, and the $R^{2}_{k^{a},\widetilde{(k,j)}}$ of the phase averages.
+Both the correlation coefficients and $R^{2}_{k^{a},\widetilde{(k,j)}}$ values were calculated for the combined streamwise and transverse components of the modes, as well as the individual components.
+POD Mode $k$ POD Mode $j$ $\Delta\phi_{k,j}$ $\mathcal{R}_{k^{a},j}$ $\mathcal{R}_{k^{a},j}^{u}$ $\mathcal{R}_{k^{a},j}^{v}$ $R^{2}_{k^{a},\widetilde{(k,j)}}$ $(R^{2}_{k^{a},\widetilde{(k,j)}})^{u}$ $(R^{2}_{k^{a},\widetilde{(k,j)}})^{v}$
+1 2 $0.47\pi$ 0.85 0.82 0.89 0.86 0.82 0.88
+2 1 $0.46\pi$ 0.78 0.75 0.82 0.78 0.74 0.81
+3 2 $0.47\pi$ 0.77 0.75 0.79 0.79 0.77 0.80
+4 5 $0.51\pi$ 0.86 0.84 0.89 0.87 0.84 0.89
+5 4 $0.51\pi$ 0.79 0.76 0.82 0.79 0.75 0.82
+6 8 $0.49\pi$ 0.81 0.73 0.86 0.83 0.77 0.87
+7 9 $0.47\pi$ 0.59 0.61 0.58 0.67 0.67 0.67
+8 6 $0.48\pi$ 0.73 0.63 0.82 0.73 0.61 0.80
+9 10 $0.46\pi$ 0.75 0.71 0.82 0.75 0.70 0.83
+10 9 $0.45\pi$ 0.71 0.65 0.80 0.71 0.65 0.80
+Table 5: Paired POD modes using the analytic signal of the $k^{th}$ POD mode, and the corresponding phase angle, correlation coefficient and $R^{2}_{k^{a},\widetilde{(k,j)}}$ of the phase averages.
+Superscripts u and v refer to the streamwise and transverse components, respectively.
+For each of the first ten POD modes, the paired POD mode corresponds to the analytic signal at a phase angle of $\mp\pi/2$.
+Consequently, the $j^{th}$ POD mode best matches the Hilbert transform of the corresponding $k^{th}$ POD mode.
+A pair of POD modes representing a structure propagating in the streamwise direction therefore requires that the POD modes be a mutual best match with phase angle shifts of $\pm 0.5\pi$.
+Based on this criterion, the POD mode pairs identified using the Hilbert transform are the first and second POD modes, the fourth and fifth POD modes, the sixth and eighth POD modes, and the ninth and tenth POD modes.
 These are the same pairings as those identified using the HPOD modes.
 Figure 11: Phase-averaged (a) streamwise and (c) transverse components of the first and second POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the first POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
-The streamwise component of the analytic signal of the first POD mode, shown in figure 10b, is a close match to the phase average of the first two POD modes, shown in figure 10a.
-Both show the same long regions of alternating sign propagating downstream from the sphere, with the main variation between them being for $x/D<2$, where the non-propagating features are located, as shown in figure 5a(iv).
-This is an artefact of applying the Hilbert transform to the POD mode, due to these structures being present in the second POD mode but not the first POD mode, on which the Hilbert transform was performed.
-This also explains the variation in the $R^{2}$ of the analytic signal of the first POD mode and the second POD mode and that of the analytic signal of the second POD mode and the first POD mode, being 0.89 and 0.76, respectively.
-This difference is due to the non-propagating components of these modes being absent in the other mode, but propagated by the Hilbert transform.
-The variation between the transverse components of the phase-averaged first and second POD modes, shown in figure 10c, and the analytic signal of the first POD mode, shown in figure 10d, is significantly less than in the streamwise components.
-Both exhibit the same symmetric regions of alternating sign, which increase in length and decrease in intensity as they travel downstream.
-This is consistent with the vertical fluctuations from the flow coming around the sphere decaying once they are sufficiently far into the wake.
+See supplementary movie 5 for an animated version of this figure.
+The streamwise component of the analytic signal of the first POD mode, shown in figure 11b, closely matches the phase average of the first two POD modes, shown in figure 11a.
+Both exhibit extended regions of alternating sign propagating downstream from the sphere.
+The primary difference occurs for $x/D<2$, where the structures in the POD modes are seen coming around the sphere, as illustrated in figure 6a(iv).
+Conversely, the assumed periodicity of the Hilbert transform couples these structures with those further downstream in the HPOD mode, as shown in figure 11b(iv).
+This discrepancy arises from applying the Hilbert transform to the first POD mode, as these structures exist in the second POD mode but not in the first.
+Consequently, the $R^{2}_{k^{a},\widetilde{(k,j)}}$ values differ: 0.89 for the analytic signal of the first mode when compared to the second mode, and 0.76 for the reverse comparison.
+This variation arises because non-propagating components in one mode are absent from the other but are propagated by the Hilbert transform.
+The variation between the transverse components of the phase-averaged first and second POD modes, shown in figure 11c, and the analytic signal of the first POD mode, shown in figure 11d, is considerably smaller than that observed in the streamwise components.
+Both exhibit symmetric regions of alternating sign that elongate and weaken in intensity as they propagate downstream.
+This behaviour is consistent with the transverse fluctuations generated by the flow around the sphere decaying as they travel downstream.
 Figure 12: Phase-averaged (a) streamwise and (c) transverse components of the fourth and fifth POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the fourth POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
-The streamwise component of the analytic signal of the fourth POD mode, shown in figure 11b, retains more of the asymmetry of the phase-averaged fourth and fifth POD modes, shown in figure 11a, than the second HPOD mode, shown in figure 6b.
-This is also true for the transverse component of the analytic signal of the fourth POD mode, shown in figure 11d.
-The transverse component of the analytic signal of the fourth POD mode, shown in figure 11d, also retains the asymmetry seen in the phase average of the fourth and fifth POD modes, shown in figure 11c.
-Although the filtering of the Hilbert transform is less prominent in the analytic signal of the fourth POD mode than in the second HPOD mode, shown in figures 6b and 6d, it is still present, most noticeably figures 11b(iii) and 11d(iv).
+See supplementary movie 6 for an animated version of this figure.
+The streamwise component of the analytic signal of the fourth POD mode, shown in figure 12b, retains more of the asymmetry of the phase-averaged fourth and fifth POD modes, shown in figure 12a, than is observed in the second HPOD mode, shown in figure 7b.
+The transverse component of the analytic signal of the fourth POD mode, shown in figure 12d, also retains the asymmetry present in the phase average of the fourth and fifth POD modes, shown in figure 12c.
+Although the bias of the Hilbert transform towards propagating structures is less pronounced in the analytic signal of the fourth POD mode than in the second HPOD mode, shown in figures 7b and 7d, it remains evident, particularly in figures 12b(iii) and 12d(iv).
 Figure 13: Phase-averaged (a) streamwise and (c) transverse components of the sixth and eighth POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the sixth POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
-While there is some filtering in the analytic signal of the sixth POD mode, the asymmetry seen in the phase average of the sixth and eighth POD modes is preserved where the HPOD modes were planar symmetric.
-This can be seen in the positive region introduced to the streamwise component of the third HPOD mode at $x/D=2.5$, shown in figure 7b(i), which is not present in the phase average of the sixth and eight POD modes, shown in figure 12a(i), or the analytic signal of the sixth POD mode, shown in figure 12b(i).
-It is also shown in the asymmetric region at $x/D=3.5$ in the transverse component of the sixth and eighth POD modes, shown in figure 12c(i), which is also off-centre in the analytic signal of the sixth POD mode, shown in figure 12d(i), but located at the centreline in the third HPOD mode, shown in figure 7d(i).
-Additionally, the intensity of the structures further downstream, such as in figure 12a(i), is clearly preserved in the analytic signal of the sixth POD mode, as shown in figure 12b(i), where the intensity of the same region is significantly reduced in the third HPOD mode, shown in figure 7b(i).
-This intensity in the phase average of the sixth and eighth POD modes around $x/D=4$, shown in figure 12a(v), is propagated downstream in the analytic signal of the sixth mode, as shown in figure 12b(v), whereas the intensity is lower in the same region in the third HPOD mode, as shown in figure 7b(v).
-This shows that the filtering which occurs from the Hilbert transform also affects the analytic signal of the POD modes.
+See supplementary movie 7 for an animated version of this figure.
+While the analytic signal of the sixth POD mode exhibits more symmetry than the phase average of the sixth and eighth POD modes, its influence on the turbulent structures is less pronounced than that observed in the third HPOD.
+The analytic signal of the sixth POD mode introduces the negative region at $x/D=2.5$ in the streamwise component shown in figure 13a(v), similar to that observed in the third HPOD mode shown in figure 8b(v), which is not present in the phase average of the sixth and eighth POD modes, shown in figure 13b(v).
+The analytic signal of the sixth POD mode also retains more of the transverse shifting in the structures in the transverse velocity, as seen in figure 13d, than the third HPOD mode, shown in figure 8d.
+Additionally, the intensity of the structures further downstream is preserved in the analytic signal of the sixth POD mode, as illustrated in figure 13b(i), where the same region exhibits significantly reduced intensity in the third HPOD mode, shown in figure 8b(i).
 Figure 14: Phase-averaged (a) streamwise and (c) transverse components of the ninth and tenth POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the ninth POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
-While there is a significant difference between the streamwise components of the analytic signal of the ninth POD mode, shown in figure 13b, and that of the phase average of the ninth and tenth POD modes, shown in figure 13a, for $x/D<2$, they are more similar further downstream of the sphere.
-The structures in the phase average of the ninth and tenth modes are smaller scale for $x/D<2$ than they are for $x/D>2$, which is consistent with smaller scale fluctuations in the near wake breaking down as they travel downstream while the larger scales remain.
-As a result, these scales are not propagated by the Hilbert transform and are thus filtered in the analytic signal.
-The differences between the transverse components of the analytic signal of the ninth POD mode, shown in figure 13d, and the phase average of the ninth and tenth POD modes, shown in figure 13c, are significantly less prominent.
-However, there is still variation in the structures upstream of $x/D=1.5$, as shown in figure 13c(ii) and 13d(ii), and figures 13c(iii) and 13d(iii).
+See supplementary movie 8 for an animated version of this figure.
+While there is a significant difference between the streamwise components of the analytic signal of the ninth POD mode, shown in figure 14b, and that of the phase average of the ninth and tenth POD modes, shown in figure 14a, for $x/D<2$, they become more similar further downstream of the sphere.
+The structures in the phase average of the ninth and tenth modes are smaller in scale for $x/D<2$ than they are for $x/D>2$, which is consistent with the breakdown of smaller-scale fluctuations in the near wake as they travel downstream while the larger-scale structures persist.
+Consequently, these structures are less significant to the Hilbert transform and therefore attenuated in the analytic signal.
+The differences between the transverse components of the analytic signal of the ninth POD mode, shown in figure 14d, and the phase average of the ninth and tenth POD modes, shown in figure 14c, are less pronounced.
+However, some variation remains in the upstream structures at $x/D>1.5$, as shown in figure 14c(ii) and 14d(ii), and figures 14c(iii) and 14d(iii), particularly at the end of the domain, which results from spectral leakage in the analytic signal caused by the non-periodicity of the POD modes.
 4.4 Propagating Structures in the Wake of a Sphere
-The contribution of the pairs of POD modes identified to the total TKE of $\boldsymbol{X}$ is given in table 6.
+Table 6 compares the contributions of the identified POD mode pairs and their corresponding HPOD modes to the total turbulent kinetic energy (TKE).
+The first pair of POD modes (modes 1 and 2) represents the largest contribution to TKE, at 8.24%, which slightly exceeds the 7.70% contribution of the first HPOD mode.
+Subsequent pairs contribute progressively less energy, with the fourth pair (modes 9 and 10) accounting for 2.60% of the total TKE.
+The corresponding HPOD modes exhibit similar trends, although their contributions differ from those of the POD mode pairs.
+The first two HPOD modes contribute more proportionally to the analytic signal than the corresponding POD mode pairs contribute to the original data, whereas the third and fourth HPOD modes contribute less than their POD counterparts.
+This is consistent with the higher-order POD modes containing a greater proportion of non-propagating structures, which are excluded from the leading HPOD modes.
 HPOD Mode TKE [%] POD Modes TKE [%]
@@
 Table 6: Turbulent kinetic energy contributions of the HPOD modes to $\boldsymbol{X}^{a}$ and of the POD mode pairs to $\boldsymbol{X}$.
-In order to determine the nature and effects of the propagating structures identified from the POD modes, the paired POD modes were used to phase average the instantaneous velocity fluctuations
+In order to examine the nature and effects of the propagating structures identified from the HPOD and POD modes, the HPOD modes and corresponding POD mode pairs were employed to phase-average the instantaneous velocity fluctuations.
+The phase-averaged velocity fluctuations based on the phase angle of the $i^{th}$ and $j^{th}$ POD modes are defined as
 $\begin{bmatrix}\widetilde{u^{\prime}}(\boldsymbol{x},\phi)\\ \widetilde{v^{\prime}}(\boldsymbol{x},\phi)\\ \end{bmatrix}=\frac{1}{N_{\phi}}\sum_{k=1}^{K}a_{k}(t)\psi_{k}(\boldsymbol{x})\,|\,\phi_{i,j}=\phi,$ (53)
-where $i$ and $j$ are paired POD modes, and $N_{\phi}$ is the number of instantaneous velocity fields included for the phase angle $\phi$.
-The mean velocity was added to the phase-averaged velocity fluctuations to determine the phase-averaged velocity
+where $i$ and $j$ are paired POD modes, and $N_{\phi}$ is the number of instantaneous velocity fields used in the phase average at corresponding phase angle $\phi$.
+The mean velocity is combined with the phase-averaged velocity fluctuations to obtain the phase-averaged velocity
 $\begin{bmatrix}\widetilde{u}(\boldsymbol{x},\phi)\\ \widetilde{v}(\boldsymbol{x},\phi)\\ \end{bmatrix}=\begin{bmatrix}\widetilde{u^{\prime}}(\boldsymbol{x},\phi)\\ \widetilde{v^{\prime}}(\boldsymbol{x},\phi)\\ \end{bmatrix}+\begin{bmatrix}\overline{u}(\boldsymbol{x})\\ \overline{v}(\boldsymbol{x})\\ \end{bmatrix}$, (54)
-from which the phase-averaged vorticity is defined as
+from which the phase-averaged out-of-plane vorticity is computed as
 $\widetilde{\omega}(\boldsymbol{x},\phi)=\frac{\partial\widetilde{v}(\boldsymbol{x},\phi)}{\partial x}-\frac{\partial\widetilde{u}(\boldsymbol{x},\phi)}{\partial y}.$ (55)
-The phase-averaged planar Reynolds stress is given by the phase average of the product of the streamwise and transverse components of the instantaneous velocity fluctuations.
+The phase-averaged planar Reynolds stress is obtained by averaging the product of the streamwise and transverse components of the instantaneous velocity fluctuations at a given phase angle
+$\widetilde{u^{\prime}v^{\prime}}(\boldsymbol{x},\phi)=\frac{1}{N_{\phi}}\sum_{k=1}^{K}\big(a_{k}(t)\psi^{u}_{k}(\boldsymbol{x})\big)\big(a_{k}(t)\psi^{v}_{k}(\boldsymbol{x})\big)|\,\phi_{i,j}=\phi,$ (56)
+where $\psi^{u}_{k}(\boldsymbol{x})$ and $\psi^{v}_{k}(\boldsymbol{x})$ are the streamwise and transverse components of the $k^{th}$ POD mode, respectively.
 The TKE of the phase-averaged velocity is given by
@@
 Figure 15: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the first and second POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
+See supplementary movie 9 for an animated version of this figure.
 Figure 16: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the first HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
-The motion represented by the first and second POD modes, shown in figure 14, and the first HPOD mode, shown in figure 15, is a flapping of the wake.
-The phase-averaged velocity from using the first and second POD modes and the first HPOD mode, shown in figure 14a and 15a, respectively, are qualitatively indistinguishable, as is the phase-averaged vorticity, shown in figure 14b and 15b, respectively.
-The structures in the planar Reynolds stresses of the phase-average of POD modes, shown in figure 14c, are longer and more intense for $x/D<3$ than those in the first HPOD mode, shown in figure 15c, and disappear after $x/D=3$.
-The structures in the planar Reynolds stress of the HPOD mode are shorter and more consistent in size across the measured domain.
-The intensity of the structures is lower than those in the POD modes, but decays more slowly, with structure visible past $x/D=3$.
-The phase-averaged TKE of the POD modes, shown in figure 14d, is concentrated in periodically shed packets which initially increase in intensity until $x/D=2$, and then decay by $x/D=3$.
-The TKE of the HPOD mode, shown in figure 15d, reaches exhibits the same periodic structures as the POD modes, however, the intensity is increased.
-This is consistent with the TKE of the analytic signal of the turbulent fluctuations being significantly higher than that of the actual turbulent fluctuations.
+See supplementary movie 10 for an animated version of this figure.
+The first and second POD modes, shown in figure 15, and the first HPOD mode, shown in figure 16, represent a flapping motion in the wake.
+The phase-averaged velocity obtained using the first and second POD modes and the first HPOD mode, shown in figure 15a and 16a, respectively, are qualitatively indistinguishable, as is the phase-averaged vorticity, shown in figures 15b and 16b, respectively.
+The structures in the planar Reynolds stresses from the phase-averaged POD modes, shown in figure 15c, are longer and more intense for $x/D<3$ than those in the first HPOD mode, shown in figure 16c, and disappear after $x/D=3$.
+The structures in the planar Reynolds stress associated with the HPOD mode are shorter and more consistent in size across the measured domain.
+The intensity of the structures is lower than that of the POD modes, but decays more slowly, with structure remaining visible beyond $x/D=3$.
+The phase-averaged TKE of the POD modes, shown in figure 15d, appears in periodically shed packets that increase in intensity until $x/D\approx 3$, and then decay after $x/D\approx 3$.
+The TKE of the HPOD mode, shown in figure 16d, exhibits the same periodic structures as the POD modes but with greater intensity.
+This is consistent with the TKE of the analytic signal of the turbulent fluctuations being significantly greater than the actual turbulent fluctuations.
 Figure 17: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the fourth and fifth POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
+See supplementary movie 11 for an animated version of this figure.
 Figure 18: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the second HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
-The structure represented by the fourth and fifth POD modes and the second HPOD mode is a pulsating in the streamwise fluctuations, as can be seen from the deformation of the low velocity region behind the sphere, shown in figure 16a and 17a, respectively.
-The streamwise pulsation can also be seen in the phase-averaged out-of-plane vorticity, with the high-vorticity regions either side of the sphere extend further toward the centreline close to the sphere in figures 16b(ii) and 17b(ii).
-This extension toward the centreline occurs around $x/D=2$ in figures 16b(v) and 17b(v).
-While the structures in the planar Reynolds stress are highly variable in shape and size in the phase-averaged POD modes, shown in figure 16c, they are consistent in size and shape in the HPOD mode, shown in figure 17c.
-The size and shape of the phase-averaged TKE of the POD modes, shown in figure 16d, is more consistent than that of the planar Reynolds stresses.
-These structures are similar to, but less planar symmetric, than the structures in the phase-averaged TKE of the HPOD mode, which also exhibit a higher intensity with a more consistent pattern.
-The intensity of the TKE increases from directly behind the sphere until $x/D=2$, as seen in figure 17d(ii), before decreasing further downstream, as seen in figure 17d(iv).
+See supplementary movie 12 for an animated version of this figure.
+The structure represented by the fourth and fifth POD modes and the second HPOD mode corresponds to a pulsation in the streamwise fluctuations, as evidenced by the deformation of the low-velocity region behind the sphere, shown in figure 17a and 18a, respectively.
+The streamwise pulsation is also evident in the phase-averaged out-of-plane vorticity, where the high-vorticity regions on either side of the sphere extend further toward the centreline near the sphere, as shown in figures 17b(ii) and 18b(ii).
+This extension toward the centreline occurs at approximately $x/D=2$, as shown in figures 17b(v) and 18b(v).
+While the structures in the planar Reynolds stress vary considerably in shape and size in the phase-averaged POD modes, shown in figure 17c, they remain relatively uniform in size and shape in the HPOD mode, as shown in figure 18c.
+The size and shape of the phase-averaged TKE associated with the POD modes, shown in figure 17d, are more uniform than those of the planar Reynolds stresses.
+These structures are similar to, but less planar-symmetric than, those in the phase-averaged TKE of the HPOD mode, which also exhibit greater intensity and a more uniform pattern.
+The intensity of the TKE increases immediately behind the sphere up to $x/D=2$, as seen in figure 18d(ii), before decreasing further downstream, as seen in figure 18d(iv).
 Figure 19: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the sixth and eighth POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
+See supplementary movie 13 for an animated version of this figure.
 Figure 20: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the third HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
-The sixth and eighth POD modes represent a flapping motion in the wake, similar to that of the first and second POD modes but with a shorter wavelengths.
-Thus, the phase-averaged velocity, shown in figure 18a, and phase-averaged vorticity, shown in figure 18b, are similar to those of the first and second POD modes.
+See supplementary movie 14 for an animated version of this figure.
+The sixth and eighth POD modes correspond to a flapping motion in the wake, similar to that of the first and second POD modes, but with a shorter wavelength.
+Consequently, the phase-averaged velocity, shown in figure 19a, and phase-averaged vorticity, shown in figure 19b, closely resemble those of the first and second POD modes.
 The smaller structures are evident in the phase-averaged planar Reynolds stress, shown in figure 19c, particularly near the sphere.
-Further downstream of the sphere, the structures increase in streamwise extent, most notable in figure 18c(iv).
-The smaller structures in the phase-averaged TKE, shown in figure 18d, are more definitively separated across the centreline of the sphere than in the case of the first and second POD modes, where the larger structures extended to the centreline.
-The structures in the phase-averaged planar Reynolds stresses of the third HPOD mode, shown in figure 19c, are more consistent in size, shape and arrangement than those of the POD modes.
-There is also a region with low planar Reynolds stress around $x/D=2$ for all phase angles shown in figure 19, which reflects the smaller structures in the streamwise turbulent fluctuations in this regions, shown in figure 7c.
-The structures in the phase-averaged TKE are also more clearly defined for the HPOD modes, as shown in figure 19d.
+Further downstream of the sphere, the structures increase in streamwise extent, as illustrated in figure 19c(iv).
+The smaller structures in the phase-averaged TKE, shown in figure 19d, are more distinctly separated across the sphere’s centreline than in the first and second POD modes, where larger structures extend to the centreline.
+The structures in the phase-averaged planar Reynolds stresses of the third HPOD mode, shown in figure 20c, exhibit more uniformity in size, shape and arrangement compared to those of the POD modes.
+Additionally, a region with low planar Reynolds stress occurs around $x/D=2$ for all phase angles shown in figure 20b, corresponding to smaller structures in the streamwise turbulent fluctuations in this region, shown in figure 8b.
+The structures in the phase-averaged TKE are more distinctly defined for the HPOD modes, as shown in figure 20d.
 Figure 21: Phase-averaged (a) velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE using the ninth and tenth POD modes at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
+See supplementary movie 15 for an animated version of this figure.
 Figure 22: (a) Velocity, (b) vorticity, (c) planar Reynolds stress, and (d) TKE of the fourth HPOD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
-The structure represented by the ninth and tenth POD modes, and the fourth HPOD mode, is similar to that represented by the fourth and fifth POD modes, and the second HPOD mode, namely periodic streamwise fluctuations in the wake.
-The phase-averaged velocity, shown in figures 20a and 21a for the POD modes and HPOD mode, respectively, and the phase-averaged vorticity, shown in figures 20b and 21b for the POD modes and HPOD mode, respectively, are qualitatively similar to those of the lower order modes.
-The structures in the planar Reynolds stress phase-averaged using the POD modes, shown in figure 20c, are small and round in the near wake of the sphere and larger and more ovular further downstream, with the latter being angles away from the centreline in accordance with the V-shaped structures in the POD modes.
-While the arrangement of these structures is variable, particularly close to the sphere, in the phase-average using the POD modes, the planar Reynolds stress of the HPOD mode, shown in figure 21c, clearly shows the round structures for $x/D<2$ and more V-shaped structures for $x/D>2$.
-Both the POD modes and HPOD modes show a noticeable reduction in the planar Reynolds stress around $x/D=2$, which delineates the rounder structures from the V-shaped ones.
-However, the HPOD modes display this more prominently than the POD modes, which also exhibit additional structures in the very near wake which are filtered out in the HPOD modes.
-The change between round structures near the sphere to more V-shaped structures is reflected in the phase-averaged TKE, shown in figure 20d and 21d, for the POD modes and HPOD mode, respectively.
-This also shows that the structures close to the sphere are more intense and decay in intensity toward $x/D=2$, with the V-shaped structures increasing in intensity around $x/D=3$ before decaying in intensity downstream.
-As in the case of the planar Reynolds stress, the structures in the TKE are more clearly defined in the HPOD mode due to the filtering of the Hilbert transform.
-The structures elucidated by the POD mode pairs identified from the HPOD modes differ from those in the corresponding HPOD modes.
-The primary difference between these methods results from the filtering inherent in the Hilbert transform, as the non-propagating structures are given less importance in the HPOD, while propagating structures are amplified.
-The phase-averaging of the resulting modes thus results in more consistently shaped structures making it easier to identify and interpret the propagating structures.
-However, the filtering out of non-propagating structures may result in a poorer representation of the flow structure where the flow is three-dimensional.
-While applying the Hilbert transform directly to the POD modes is more computational efficient than performing the HPOD for pairing POD modes, it also forcibly propagates structures in the flow which may not propagate in reality.
-Thus, both methods of pairing POD modes introduce non-physical artefacts into the analysis of the flow and this should be considered when interpreting results.
+See supplementary movie 16 for an animated version of this figure.
+The structure represented by the ninth and tenth POD modes and the fourth HPOD mode resembles that of the fourth and fifth POD modes and the second HPOD mode, namely a periodic streamwise pulsation in the wake.
+The phase-averaged velocity and vorticity for the POD modes and HPOD mode, shown in figures 21a–b and 22a–b, respectively, closely resemble those of the lower-order modes.
+The structures in the planar Reynolds stress phase-averaged using the POD modes, shown in figure 21c, are small and rounded in the near wake of the sphere but become larger and more ovular further downstream, with the latter angled away from the centreline in alignment with the V-shaped structures in the POD modes.
+While the arrangement of these structures varies considerably near the sphere in the POD mode phase average, the planar Reynolds stress of the HPOD mode, shown in figure 22c, clearly exhibits rounded structures for $x/D<2$ and more V-shaped structures for $x/D>2$.
+Both the POD modes and HPOD modes exhibit a pronounced reduction in the planar Reynolds stress near $x/D=2$, which separates the rounded structures from the V-shaped ones.
+However, the HPOD modes display this transition more prominently, whereas the POD modes exhibit additional structures in the very near wake which are absent from the HPOD modes.
+The transition from round structures near the sphere to V-shaped structures further downstream is evident in the phase-averaged TKE, shown in figure 21d and 22d, for the POD modes and the HPOD mode, respectively.
+This also indicates that the structures near the sphere are more intense and decay towards $x/D=2$, whereas the V-shaped structures increase in intensity around $x/D=3$ before gradually weakening downstream.
+Similar to the planar Reynolds stress, the structures in the TKE are more distinctly defined in the HPOD mode due to the reduced significance of non-propagating structures in the analytic signal.
+The structural details revealed by the POD mode pairs identified from the HPOD modes differ from those of the corresponding HPOD modes.
+The primary distinction between these methods stems from emphasis on propagating structures in the leading HPOD modes, whereas non-propagating structures are relegated to higher-order modes.
+Phase-averaging the resulting modes yields more uniformly shaped structures, thereby facilitating the identification and interpretation of propagating structures.
+However, the reduced emphasis on non-propagating structures may result in a less accurate representation of flow features that do not propagate in the direction of the Hilbert transform, such as those confined to the near wake of the sphere.
+Although applying the Hilbert transform directly to the POD modes is computationally more efficient than performing HPOD to pair POD modes, it also artificially imposes propagation on structures that may not physically propagate.
+Thus, both methods of pairing POD modes introduce non-physical assumptions into the flow analysis, which should be carefully considered when interpreting the results.
 5 Concluding Remarks
-The velocity in the wake of a sphere at $Re_{D}=7780$ was measured using 2C-2D MCCDPIV, and POD and HPOD were applied to the turbulent fluctuations.
-The resulting POD and HPOD modes were compared to determine if the HPOD modes, which accentuate propagating structures, could be used to match POD modes which represented the same structure at different phase angles.
-The Hilbert transform of the POD modes was then used as a substitute for the HPOD modes in order to determine paired POD modes without performing the computationally-intensive HPOD.
-The POD mode pairs identified using these methods were then used to phase average the turbulent fluctuations to examine the propagating structures in the wake of the sphere.
+The velocity in the wake of a sphere at $Re_{D}=7780$ was measured using 2C-2D MCCDPIV, with POD and HPOD applied to the turbulent fluctuations.
+The resulting POD and HPOD modes were compared to assess whether the HPOD modes, which emphasise propagating structures, could be used to identify POD mode pairs representing the same structure at different phase angles.
+The analytic signal of the POD modes was then employed as a substitute for the HPOD modes to identify paired POD modes without performing the computationally intensive HPOD.
+The POD mode pairs identified using these methods were then used to phase-average the turbulent fluctuations to enable the examination of propagating structures in the sphere’s wake.
 The leading HPOD modes are shown to represent the same structures as the leading POD modes, with a greater emphasis on the propagating structures in the former.
-While each HPOD mode consists of a real and imaginary part representing a phase-shift of 90 in the mode, this comes at the cost of filtering from the Hilbert transform.
-As a result, it can be difficult to differentiate between what is part of the flow structures and what is an artefact of this filtering, particularly in the case of two-dimensional measurements of a three-dimensional flow.
-Performing the Hilbert transform on the POD modes to create an analytic signal of the POD modes is shown to identify the same pairs of POD modes identified using the HPOD modes.
-This method allows the same structures identified using the HPOD modes to be identified, without introducing the filtering resulting from the HPOD.
-The phase average of the turbulent fluctuations using the identified POD mode pairs represented flapping and pulsating motions in wake of the sphere.
-As the velocity fields in the current work are not time-resolved, the Hilbert transform was performed in the streamwise direction, and thus the structures identified are those which propagate in that direction.
-As a result, the temporal information of the structures is not given by either decomposition.
-However, the paired POD modes gave insight into the dynamics of the structures in the turbulent fluctuations when the complexity of the flow made identifying related POD modes challenging.
+While each HPOD mode consists of a real and imaginary part corresponding to a phase-shift of $\pi/2$ in the mode, this comes at the cost of spectral leakage due to the assumption of periodicity imposed by the Hilbert transform.
+As a result, it can be difficult to differentiate between genuine flow structures and artefacts introduced by the transform, particularly when taking two-dimensional measurements of a three-dimensional flow.
+Performing the Hilbert transform on the POD modes to create an analytic signal of the POD modes is shown to identify the same POD mode pairs as the HPOD modes.
+This method allows the same structures to be identified, without introducing the non-physical elements introduced by the HPOD.
+The identified POD mode pairs represent flapping and pulsating motions in the wake of the sphere.
+As the velocity fields in the present study are not temporally resolved, the Hilbert transform was applied in the streamwise direction, meaning the identified structures are those propagating along this axis.
+As a result, the temporal information of the structures is not captured by either decomposition.
+However, the paired POD modes offer insight into the dynamics of turbulent structures, even when flow complexity makes it difficult to identify corresponding POD modes.
 Acknowledgements
 Shaun Davey gratefully acknowledges the support of the Australian Commonwealth Government through a Research Training Program (RTP) Scholarship.
-This research was undertaken with the support of resources provided via a Monash HPC-NCI Merit Allocation from the National Computational Infrastructure (NCI Australia), an NCRIS-enabled capability supported by the Australian Government.
-The authors would like to acknowledge the support of the Australian Research Council (ARC) in funding this research through Linkage Infrastructure, Equipment and Facilities (LIEF) grants LE100100222 and LE180100166.
+This research was undertaken using resources provided through a Monash HPC-NCI Merit Allocation from the National Computational Infrastructure (NCI Australia), an NCRIS-enabled capability supported by the Australian Government.
+The authors also acknowledge the support of the Australian Research Council (ARC) in funding this research through Linkage Infrastructure, Equipment and Facilities (LIEF) grants LE100100222 and LE180100166.
 Declaration of Interests